A Complete Bibliography of Publications in the Journal of Modern Applied Statistical Methods

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

18 January 2025 Version 1.00

Title word cross-reference

(Y < X) [AW05]. 0.05 [Kna15]. 20 [LFG19]. 2 × 2 [HL07, KRP02, MCL09]. 2 × C [Pos02a]. 4^{r}2^{n−p} [ACBM06]. ^{+} [Saw02a]. _{2} [Saw02a]. α [RNN13b]. X̄ [Cox13, Kho07]. br^{2} [KKZB16]. C(t, p, s) [Ree11b]. χ^{2} [Saw02a]. D [ATvBB17, FC04, Wal05a, Wal15, Wil18d]. ϵ2 [Wal15]. η2 [Wal15]. F [BB19, BR14, El 11, KSS10, PA07, PMZ08, ZA08]. g [Wal15]. h [SS16]. K [AM14, DNUH22, SG21, KDMAA18]. ω2 [Wal15]. p [And13a, And13b, Goo13, Man13, PB13, Wil19e]. P (π1 > π2) [KSM13]. P (X < Y) [Bak03, EB04]. P (Y < X) [Ng06]. P ∗ [BR03]. p > 0.05 [SY02]. P [Y < X] [MESSE16]. r [Smi09, Wal05b]. R^{2} [WS16a]. r × c [LA14]. S_{u} [PS02b]. S_{w} [BS02]. σ_{1}^{2} ≠ σ_{2}^{2} [Saw02b]. T [Lec07, RR24, ZRB23, ASS19, AB11, FKL^{+}03, GJ17, LL16, MOD07, NT25, Ozg16, SM02, Saw05, TGhC^{+}05, Web07, WS09]. U [BRT17]. φ [SSM18].

-based [ZA08]. -Curvature [RNN13b]. -Distribution [KSS10]. -Divergence [SSM18]. -maps [SG21]. -Nearest [DNUH22]. -Occasion [SS16]. -Optimal [ATvBB17]. -Repeated [AM14]. -Test [GJ17, LL16, BRT17, PA07, PMZ08]. -tests [RR24, El 11]. -th [KDMAA18]. -Values [Goo13, Man13]. -weighting [KKZB16]. -Year [LFG19].

1 [DiS02, KIR12]. 10 [Alo04]. 11 [Ree04]. 16 [Dem05]. 17 [BB05]. 18 [OO05]. 19 [Wal05b].

2 [FZ09, MADT03]. 2018 [HPSR23]. 21 [DP05]. 21st [ZK16, Bra03, RL16]. 22 [KMG06]. 23 [Alo06]. 24 [Hea06]. 25 [BB06]. 26 [NS07]. 27 [SH07]. 28 [SH09]. 29 [FC09]. 2PNO [SH07, SH09].

3 [Hea02]. 31 [Tay11]. 32 [CKP12, LOCMS13]. 33 [MS14]. 34 [Wal15]. 35 [KH16]. 36 [WS16a]. 37 [ASM^{+}16]. 38 [Wal16a]. 39 [ASR^{+}16].

4 [Fay02]. 4.0 [Pos02a]. 40 [RM16]. 407 [Whi07]. 41 [ANAS16]. 42 [OMN16]. 44 [Tak17a]. 45 [ATvBB17]. 46 [SAZ17]. 47 [LCW^{+}17]. 48 [Wal17b]. 49 [ÖK17].

5 [Whi07]. 50 [MBS17]. 51 [AK18a]. 52 [CDPT19]. 53 [Lan19]. 54 [AAH^{+}19]. 55 [BD20]. 57 [AK20].

6 [Hea03].

8 [RP03a].

9 [Wal03]. 90/95 [Fah02a].

Aalen [GVCM18]. AB-Type [LFG19]. AB/BA [Ree06]. ABE [Eni09]. Abilities [Weh03]. Absolute [Die06, Die09b, OMN16]. Academic [AK10]. Accelerated [Ebr06, LRT20]. Acceptance [AOCAN19, Rao11, SLC17]. Accommodate [MRKM^{+}17]. Accommodation [LB16, LLB17]. According [EYS^{+}07, OE12]. Account [MLS10]. Accounting [KRP02]. Accuracy [Kan05, Leh15, SF19, SS09, Wal04]. Accurate [BH06, Cen04, Leh15]. ACF [Saf11, SJAZ14]. Achievement [GCH08]. Across [CD09, LZ07]. Activity [LHO^{+}18, WBN07]. Adaptive [BI02, GRM20, KWAO04, LRT20, MAP08, NH06, Wu04, Zum02]. Adding [DS16]. Addition [Mey16]. Additive [GVCM18, Lip16, NAY^{+}18, Wil06b]. Address [ZCB^{+}17]. Addressing [Abe15]. Adhering [ZZ15]. Adjusted [AS10, EO13, GCP24, LMME09, MBWZ07, PA07, UAD16]. Adjustment [Fra19, Wal17b]. Admissibility [BRT17]. Adolescent [PN03, PC14]. Advanced [WBN07]. Advantages [CID^{+}14, NR05]. Advice [NKC18, RN18]. Affect [TSS20]. Affects [KBKBG07]. Affiliated [Wal05a]. After [Pér07a, Die09a, Rah14]. Against [Ots05, Saw03a]. Age [RRS07, RS17, WBN07, WCW08]. Aggregate [CD09, SSAH18].

Agreement [BS09, Heo08]. agricultural [SKSV25]. Agriculture [EO09]. Ahead [RR03]. AIDS [PK06, LLHT03, LX06]. Air [KKZB16]. Airport [LIT24]. Akaike [EAS13]. Alert [Ber05]. Algebraist [CS06]. Algorithm [ASR^{+}16, BB04, BB05, BS04, DNUH22, Hea03, LM17, LS13, MDS03, OO05, OMN16, SAZ17, SH07, SSMK22, Su11, SJ25, SP16, Tak17a, Tak17b, TK06]. Algorithms [AM17, BD20, ML24, ÖK17, PC15]. Aligned [BZ09, LHB04, NS07, Pet02]. Allocation [BC19, DR18, Pap16, VS08]. Allows [Wil17a]. Almost [Kna19a, SGM16, PK06]. Alpha [Ben10, EYS^{+}07, JAJ24, TNH17, ZGZ07]. Alphabet [Weh03]. Alter [LS13]. Alterations [NM17]. Alternative [ANAS16, BR14, BF14, CID^{+}14, GS06, Kho04, Lan13, MK08, OMN16, RL10, WA07, YII13]. Alternatives [ASS19, BS02, El 11, LHB04, XYJ03]. Alzheimer [XGY^{+}08]. Ambulatory [SLA16]. American [Wil04]. Among [WF09, Cha24, Koc16, Tor18]. Analog [Wil18d]. Analogs [Wil19a]. Analyses [BLEL02, CF03, KOW16, Wal03, WA07, WK02a, WZ17, XGY^{+}08, ZK05]. Analysing [BO13]. Analysis [AK20, AAK10, AKA11, AK18a, AK10, AK04b, AS05, Alo06, Bha14, BB11, Bos09, BLNA^{+}22, BO13, CD09, DS02, EO13, EO09, FA11, FBR^{+}08, FS15, FC09, FF08a, FF08b, FF16, Fuk07, GMS05, Gan14, GL06, Goo09, HMNW07, HKP17, HA14, HP03, HM10, HZ16, HML07, ICS07, Kan05, KS09, Kel05, KZ08, KCCD10, KJ06, LS16a, Les11, LF11, LC05, Lip13a, Lip17a, LC20, LW05, Liu09, LOK11, LK12, LKW^{+}19, LL07, MADT03, MN17, MN21, ML03, MK05, MBWZ07, Nan10, Nan05, NAA16, Neu09, NM17, NM19, ORSCMG03, PM14, PM10, Pra15, PNZM24, Raj21, RRB12, RRS23, RNN12, RNN13b, RP03a, RP03b, RMS17, SLLL11, Saw03c, SJ07, SE12, Spe04, Tho16, TSK04, VSS16, WY03, Wal04, WPN^{+}18, Wat06, WBN07, YHPR17, YII13, ZL04]. Analytic [FB10, Rus08, Wal03]. Analytical [Kor16, LM19]. Analyze [Lip17a, ZLB19, PW24]. Analyzing [AKO03, BB07, FR04, LEKG10, LZ07, Lu16, MCL09, Per06, PI08, SK09, SIA11, WA03]. Anchoring [CPT04]. Ancillary [PS18]. ANCOVA [LP06, NS07, Wil06a, Wil16a, Wil17b, Wil18c]. and/or [Fay02]. Andersen [BS04, Kra18]. Anderson [PB13]. Animal [EO09]. ANN [OAA10]. ANOM [RRB12]. Anomaly [NSJ^{+}24]. ANOVA [BB04, BB05, BJ11, CP13, Hea03, Lan13, LCW^{+}17, NKW^{+}19, Wil05, WH16, ZA08, ZA11]. ANOVA HOV [LCW^{+}17]. AOQL [SS12]. Application [ADSBH02, AK18a, AOBCP16, ASS04, AT09, BPZ12, Bha14, BW05, CMR02, DEH^{+}06, DGJ16, FLE04, HR16, HWS^{+}24, KRP02, LX06, LHO^{+}18, MBS17, MA24, MMS08, MPS11, MBWZ07, Pau05, RAJ18, RNN12, SS04, SMC16, SSMS10, SKSV25, WE05, Wil04, XGY^{+}08, YHPR17]. Applications [ANO^{+}25, AC15, AALF14, ALF16, AFL12, AK13, AEOD24, BK09a, BZBS03, ESAH20, Gau17, Goo02, GZ17, HA07, HB08b, IS24, JAJ24, Kib06, Kum13, LFO07, MSS20a, MK23, NAY^{+}18, OON24, RS20, RNN13b, Raz13, SMSN20, SS23, SFA16a, SZH^{+}18]. Applied [DAC02, Fer12, Hus07, LH09, PW04, SM04, SLA16]. Apply [LMRM^{+}20]. Applying [AL15, Les11, MLS10, WB02, WBSMI09]. Approach [ANAS16, AASIJ10, AS16, Bha14, BA11, Cam09b, CdlTE22, CPDP18, Cox13, GSP24, GZ07, Hur09, KI16, Kho05, KCA^{+}17, LS16b, MCL09, ML03, NXFN19, PPC18, PKRK25, RTS16, SAA19b, Saw13, SM04, SFA16b, SY19, SSS21, Su05, YII13, YOO14, YS14]. Approaches [AAH^{+}19, Kib08, LKW^{+}19, SSAH18, WSN15]. Appropriate [MC11, SSMS10]. Approximate [Cam03, Cam07, Cam09a, Cam09b, Cam12, Cox13, SF19, Wal17b, WA07]. Approximating [Fra17, RC17, Wal17a]. Approximation [JA18, PDC13]. Approximations [Fah02b, VM09]. Aptitude [MK05]. Aptitude-Treatment [MK05]. AQL [SS12]. Arbitrary [Bha08]. Archimedes [SJ25]. Area [PA14, SAA19b, WBSMI09]. Arguments [Saw03a]. Arising [Gan14]. Arithmetic [DS25]. ARL [Rah14]. ARMA [DAC02]. Arrays [HY03]. Art [Kle19]. Artery [Cen04]. Article [NKC18]. Artificial [MYY^{+}06, PNZM24, SSM18]. Asian [Wil04]. Assess [CPDP18, RRB08]. Assessing [AK04a, Cen04, KCCD10, MC08, NRSGRN02, PC15, SF19, Tho09]. Assessment [BKS13, PKRK25]. assigned [TTBG06]. Assignment [Saw04b]. associate [SG21]. Associated [AK21]. Association [KJI10, Kin03, Wil11, Wil15a, YT10]. Associations [LA14]. Assumption [Eni09, LH13, LH04, MK05]. Assumptions [AB11, BZBS03, BZ09, WWB12]. Asymmetric [Die09b, SR07, YS14]. Asymmetry [PS02b]. Asymptotic [CM02, Pos02a]. Attitude [AS16]. Attractiveness [TL15]. Attributable [FB04]. Attribute [KIR12, PS22]. Auditory [Zum02]. Augmentation [KCA^{+}17]. Authentic [Lov19b]. Author [And13b]. Auto [SW06, Wan03]. Auto-Correlated [SW06]. Autocorrelated [AO08, Bha10, Die11, Hol06]. Autocorrelation [Bha10, Die09a, Die11, RB09]. Autocovariance [Eni09]. Automated [Leh15]. Autonomous [NSJ^{+}24]. Autoregressive [MF11, OYO16, Saf11, YS16]. Auxiliary [ADAAM09, Bou10, DR18, Dan25, MKA23a, MKA23b, ÖÇ17, RSQ20, RTA22, SK18, SK13, SP14a, SP14b, TS11, TuAH19, VS08]. Average [IT18, KS06, MF11, RB11, RR24, ST07, ST08, SA11]. Averaged [LW05].

B [LH09]. B-Spline [LH09]. BA [Ree06]. Back [NR05]. Backward [LB19]. Balanced [ASG16a, ASG16b, GA17, GDG20, JVA10, VJ11]. Balancing [Lip10]. Ban [BD16]. Band [Wil17a]. Bandwidth [EME10]. Bandwidths [EM14]. Bank [CPDP18, BB11]. Bar [OS10]. Bar-Lev [OS10]. Base [Eni09]. Based [AS17, ADAR09, AL15, AA20, AS16, BE06, BAN06, BZBS03, Bea17, Bic11, BII11, CA08, DW19, DBH02, Dye16, EB04, kAEAE21, FEB13, Fuk05, Gao04, HK10, HH16, HNG19, Hur03, KJMA18, KWAO04, KM02, KH16, KDMAA18, LM17, LT07, LBC^{+}19, Lov19a, LAA17, MESSE16, MCL09, MJ20, NH12, Neu10, NW09, PPS14, RB11, RSQ20, RRB08, Rao11, RRB12, RK13, Rao14, RRS23, SMSN20, SSM18, SA10, SA07, SL16, SS09, SP14a, SA13, TK06, WW10, WK02b, Wil05, Wil06a, Wil15a, Wil16a, Wil18d, Zha05, AA21, AGKJ19, DK19, Dub04, HSA18, JR16, Koc18, MBS17, ML24, ÖÇ17, PW24, SJ25, ZA08]. Bathtub [MSS20a]. Bayes [BP16, FKP^{+}25, Lee15, Lov19a, LMME09]. Bayesian [ATvBB17, AAK22, AK20, AAK10, AKA11, AK18a, AHAO09, ACB06, AK18b, AM19, Bha14, BK22, BW05, Cam03, CT05, Cam07, Cam09a, Cam09b, Cam12, DR14, EN23, FA13, FF12, GK04, GG09, GS19, GRM20, GHSF^{+}15, HP03, HZ16, IS24, JR16, JA18, KKT13, KMA18, KSM13, KM14, KJ12, Kib08, KCA^{+}17, KK20a, KMA21, LC20, Lov19a, MESSE16, MPC^{+}16, Nan10, Nan05, NAA16, NXFN19, NH12, NE16, OY17, PZ11, PW04, Pra10a, Pra12, Pra15, Qu05, RD16, RTS16, SH14, SAA19b, SFA14, SFA16b, SA11, SBY07, SU16b, Wu04, ZC13]. Bayesian-based [JR16]. BBM [AK18b]. BBM-IRT [AK18b]. Be [Ber05, GT10, Sha06a]. Behavioral [Fra15, HKP17, PN03, Saw03b]. Behaviors [HXH06]. Behrens [HL03, Saw02b]. Being [CS08]. Believe [Saw04b]. Bell [Gil24]. Benefit [MEH08]. Benefits [BB07]. Bernoulli [MADT03]. Best [DP05, GD02, Lip13a, LOCMS13, MS03, SK09, SW08, Wil18b, ZRB23]. Beta [AK18b, FLE04, Im20, KLS07, LFO07, RJ14]. Beta-Binomial [Im20]. Beta-Mixture [AK18b]. Beta-Weibull [LFO07]. Better [Ree07]. Between [AS05, BC02, Lov19a, MMH19, RK13, RL02, Saw02b, SHA06b, Wal05b, AKQ10, And13a, And13b, KS09, Kha13, KJI10, LS06, PB13, Ree09b, Saw13, Sha08a, SRA17]. Beyond [BS09, Ben10]. Bi [Lip10, PA14]. Bi-Criteria [Lip10]. Bi-Weibull [PA14]. Bias [GW09, Har18, Har19, HS11, KCCD10, LH14, LEKG10, ML03, Pan18, SLLL11, Wal05a]. Biased [Kha13, Mir08, MK23, NLQ10, RNN12, RNN13a, SHA06b, TP16, RNN13b]. Biases [MEH08]. Biasing [KWOF02]. Bilevel [QT06]. Bimodal [FZ09]. Bimodality [FLE04, FZ02, FZ09, Kna07]. Binary [ATvBB17, AL15, AZ17, Cen04, Hea02, Im20, ICS07, KHC05, Mal02, MA06, RYZ16, Ree06, SK14, WS16a]. Binomial [AK18a, FA11, HB08b, HB12, Im20, Mir08, RNN13a, Raz02, Ree04, Ree07, SBY07, Wil19b]. Bioleaching [AOBCP16]. Biomarker [XGY^{+}08]. Birth [Saw03c]. Bivariate [ASS04, ACB06, BS11a, DR18, KKK19, MN21, SLD14, WS16b, Wil19a]. Blair [Saw04a, Sto04]. Blinder [LF11]. Block [JVA10, Kon13, MK05, TTBG06]. Blog [AS16]. Blood [SLA16]. BLUE [ADAR09, PH25]. BMDP [PS02a]. BMI [ANA12]. BOA [KS25b]. BOA-Enhanced [KS25b]. Bobovitch [OS10]. Book [ML24, Rao17]. boosting [FKP^{+}25]. Bootstrap [AM14, ANAS16, Alo04, ASR^{+}16, AA14, BAN06, CDPT19, CM02, DP21, Ebr06, HM08, OY17, PD13, Qum13, RWS09, Rus08, SM04, VO03, Wil10, ZA08]. Bootstrap-t [Wil10]. Bootstrapping [KWOF02, Kin03, Man05, PPC18]. Borda [EG11]. Borda-Winner [EG11]. Borel [TP16]. Both [DRTW17, DWT19]. Botswana [SSMS10]. Boukai [OS10]. Bound [Ben17, MA06]. Bounds [Ben10, Cam07, JR16]. Bracketed [BS02]. Bradley [Mey16]. Brahmaputra [BB11]. Brain [KKT13]. Breast [CT09, ORSCMG03, RTS16, TX11]. Brown [Wal06]. Building [LBC^{+}19]. Burr [KR16, PJ14]. Burr-Type [PJ14]. Buy [RR24].

C [KK20a]. C-AR [KK20a]. Calculate [Hur09]. Calculation [KSM13, KM14, LW05, ZRB23]. Calculations [Saw02a, Tay11]. Calculator [MS14]. Calculus [DSCS08]. Calibration [AM17, CERA24, GSP24, GCP24, IB24, KZ18, NSSO14, RSQ20, RTA22]. Calibration-Based [RSQ20]. Call [Kes15]. Can [GT10, NKC18]. Cancer [AAH^{+}19, BC05, Bos09, CT09, Jay17, KKT13, MMS08, ORSCMG03, RTS16, SZH^{+}18, TX11, WCW08]. Cannabis [WE05]. Canonical [Bos09, LH14]. Capability [SV08]. Capture [GG09, GS19]. Cardiac [MMM07]. Care [IS24]. caribaea [OCB13]. Carlo [AŞ15b, BA11, CT05, CC09, CCH19, Dye16, El 11, Ele17, FBR^{+}08, FC04, FF12, FF13, Har18, Har19, HS11, Kan05, LAA17, Mic13, MSK^{+}11, MŞ09, NKW^{+}19, Opd03, RM16, Tho09, YOO14, YS14, YS16]. Case [BH16, Bak03, BB11, BW05, BO13, HP03, KY06, KKST05, KAF05, Lee13, LH13, Lee15, LFG14, LFG19, Lip13a, LHH11, LOCMS13, MT24, NS07, PC15, PC18, Saw03a, Zha05, Zum02, Gan12]. Case-Control [Lee13, LH13, Lee15, Zha05, Gan12]. Cases [FB04]. Casual [AS05]. Categorical [AM17, BI02, LHO^{+}18, Neu09, PI08, Pos02a]. Categories [Wal05b]. Causal [Bar15, LM15, RL10]. Causality [Kna18, MSS07]. Cause [PM08]. Cause-of-Death [PM08]. Caused [NM17, RRS07]. Caution [LMS15]. Cautionary [RL10]. ’cbnet [BD20]. Cell [AZ17, Jay17]. Censored [AA14, BE06, CGH15, FEB13, HA14, KS14, KR16, Kit10, KJ06, LFO07, LAMB17, LRT20, MRKM^{+}17, PPS14, PTC18, Pra15, SFA16a, SFA16b]. Censoring [AA21, BP16, GRM20, KJK09, PP17, PS17, Pra20, SFA14]. Census [WBSMI09]. Central [CPDP18, NKC18, RN18, ZA08]. Century [ZK16, Bra03, RL16]. CEO [AGHL08]. Certain [EK03]. CFNN [ML24]. Chain [AK10, CCH19, KSB10, SIA11, XZ05]. Chain-Type [KSB10]. Challenges [Sen02]. Change [DMBR16, Fuk05, Hig04, Lip17b, LZ07, MC07, Pra14, QT06, WBN07]. Character [Bou10]. Characteristic [AASIJ10]. Characteristics [GZ07, ZK05]. Characterization [AGKJ19, AEOD24, AF23]. Characterizations [ASK16, BPZ12]. Chart [AS17, BS11b, Dye16, KY06, Kho04, KN05, KQC06, KS06, Kho07, MIN18, Rah14]. Charting [Dye11]. Charts [Alk13, Cox13, DAC02, Ela08, EC08, Kho05, RB11, Rah14, RRB12, SD17]. Chemical [SA02]. Chi [DW19, WY03, Web07]. Chi-Square [DW19, WY03, Web07]. Childhood [KKT13, RRB08]. Children [LBC^{+}19, SHA06b, Weh03]. China [PW24, ZT12]. Chinese [Pos02b]. Choice [DDM07]. cholerae [RRS07]. Choosing [Che14, Die06, EG11, MC11, Saw05]. Chronic [DS02]. Cigarettes [ANA12]. CIPP [Fin19]. Circle [LM17]. Circular [MRMH17, MMH19, OAC03]. Circumventing [KWOF02]. Cities [VA16]. Claim [IJ06]. Claims [LA18]. Class [KCCD10, KSB10, LFG14, NAK13, SS13, SPT12, SK18, BKS13, ZD10]. Classes [CEI24, OHQ^{+}20]. Classic [Ree11a]. Classical [Cam09b, Kib08, NH12, SH14, SA13]. Classification [DNUH22, IJ06, KS25b, KCCD10, Leh15, LAA17, MA17, PNZM24, SKB15]. Classification-Based [LAA17]. Classifications [BS09]. Classified [CKPJ18, YD17]. classifying [FKP^{+}25]. Cliff [Sto04]. Clifford [Saw04a]. Clinical [BLEL02, Bha14, BW05, HL07]. Closed [GG09, GS19, LM19, Ree09a]. Closed-Form [LM19]. Cluster [Alo06]. Clustered [ANA12, Dub04, Hus07, Im20]. Clustering [ÇE14, KRP02, LS13, Mar12, TK06]. Clusters [JPH23]. cN [KIR12]. CO [Saw02a]. Cochran [SAA12]. Cochrane [Die09a]. Code [BB04, BB05]. Codes [BD20, MB05, ÖK17]. Coding [SA07]. Coefficient [AKP08, BKS12, BF09, Bha10, BF13, Bha14, Cam12, Ele17, LH14, PD13, Per06, SK13, SP14b, TNH17, Wal17b, Wil15b]. Coefficients [Abe15, BF11, BF14, BG15, BKS13, EYS^{+}07, Gid07, Hus07, LS16a, Ola20, SLLL11, ZGZ07]. Cohen [HXH06, Wal05a, Wal15, Wil18d]. Coin [Man05]. Cold [BP16]. Collected [WA03]. Column [BS14]. Combination [CS04, Mec03]. Combinations [Rah14, Sta02]. Combined [KQC06, SK10, TK06]. Combines [Cli11]. Combining [ADARAM05, ANAS16, AM10, ASR^{+}16, BZ11, Bla02, Saw02a]. Comments [And13b, RL10]. Common [Bak03, OS10, Rah14]. Commonly [Dem04b]. Communicable [AA12]. Community [WB02, WBSMI09]. Comparative [ARM10, AK10, CID^{+}14, EME10, GG09, ML24, Mar12, MSS20b, RSQ20, WS09, WSN15]. Comparing [CD09, CFSF10, CCC19, FF08a, KCCD10, LAM16, NKC18, PNG08, Ree04, Rus08, RN18, ST02, Tur17, VM09, WK02a, Wil11, Wil12]. Comparison [AKQ10, AK04b, AAH^{+}19, AŞ15b, BB16, BB19, BE06, BM05, BZ11, BJ11, CE15, CC09, CC19, CGH15, DRTW17, DWT19, FKP^{+}25, Fin11, FF12, FF13, Fin14, FZ09, Geo14, GK04, GHSF^{+}15, HL16, HS18, Im20, IJ06, KKST05, KSM13, KM14, KWAO04, Kha13, KJI10, KKW^{+}06, LS16b, LLHT03, LL16, LHO^{+}18, LL07, LS17, MMH19, MA17, Mec03, MOD07, MW07, MB05, MBWZ07, MŞ09, NKW^{+}19, OCB13, ON14, OE12, PW04, PS02b, QS12, Rah14, RM07, SK14, SAZ17, SE12, SA13, SRA17, TW07, VSS16, Wal06, Wal07, Wan03, WF09, YHPR17, ZS17, ZL04]. Comparisons [ACB06, Day03, LB16, MS03, Min13, Opd03, PD05, WK02b, Wil16b]. Compensation [AGHL08]. Competing [KJ06, Lip13a, WF09]. Competition [Rod10]. Competitive [Tor18]. Compilation [ÖK17]. Complementary [KDMAA18]. Complex [LK13, Lor19, MC11, Sah05, ZLB19]. Complexity [Lan09]. Compliance [Lui06, Whi07]. Component [ANM15, AW16, FA13, KJ14a, Nan10, NLQ10, PNZM24, SFA14, SFA16a]. Components [BS14, Kan05, Kan06, Lip16, SS18, Wil06b]. Composite [BE17, LS16b]. Compound [LWZK16, MPS11, RJ14]. Compounded [HNWD05]. Comprehensive [RYZ16]. Compression [HR16]. Computer [ATvBB17, HKP17]. Computing [BB04, BB05, BB06, Doy09, Hea06, RR03, SV08, ZD10]. Concept [Goo02].

Concepts [Gau17, SKO09]. Conceptual [And13a, And13b, PB13]. Concerning [BD16]. Concomitant [KKK19]. Condition [JR16, LH14]. Conditional [AM10, NXFN19, RBAR18, SLD14, WPN^{+}18]. Conditioning [KI14]. Conditions [Mic13, Wal06, Wal07, WS19, WFW02]. Conducted [BM05]. Conducting [MBS17, PM14, PM10]. Confidence [AS10, ASS19, AKP06, AKP08, AA14, BD16, Bak03, BF09, Bha10, Cam03, Cam07, Cam09a, Cam12, CA08, CP13, CPT04, DP21, FC04, Fra17, GHSF^{+}15, HM08, JR16, KAF05, Kin03, Lec07, PS17, Ree07, Ree09a, Ree09b, Rus08, RL02, Sha06a, SF19, TNH17, Wal15, Wal16a, Wal17a, WC05, Wil17b, Wil17a, ZT06, ZA08]. Confirmative [Fin19]. Confirmatory [Bea11, FF16, KZ08, WZ17, ZK05]. Conflicts [Lov19a]. Confounders [LMRM^{+}20]. Confounding [Lee13]. Conjoint [GMS05]. Connectionist [GCH08]. Conover [BR14]. Consideration [SJ25]. Considerations [BJ11, LOCMS13]. Considering [Cox13, Wil19f]. Consistency [KS09, PZ11, SSAH18]. Consistent [HSA18]. Constant [DS16, PA14, Res08]. Constrained [LMME09]. Constraints [RTA22]. Constriction [ZC13]. Construct [CS09, NRSGRN02]. Constructing [CP13, Kon13, Rus08, SLC17]. Construction [ASG16a, ASG16b, IJ08, RB11, SG21, ZT06]. Constructive [Ser02]. Contamination [DW19]. Content [Spe04]. Context [BH16]. Context-Driven [BH16]. Contingency [IT18, JADB10, LA14, PD11, Pos02a, Sha08b, WY03, YT10]. Continuation [LOK11, LB19]. Continuity [Lan13]. Continuous [ASK16, ANA12, Eic05, KH16, Opd05, Rah17, SRW21]. Continuous-Data [Rah17, Opd05]. Continuously [HNWD05]. Contoured [BZBS03, Kib08]. Contrails [Bar15]. Contrast [LS16b, SH14]. Contrasts [LH04, TTBG06]. Control [AS17, Alk13, BS11b, DAC02, Dye11, Dye16, Ela08, EC08, Gan12, KY06, Kes15, KN05, Kho05, KQC06, KS06, Kho07, Lee13, LH13, Lee15, MIN18, RB11, SW08, SD17, VJ11, Zha05]. Controlling [HWS^{+}24]. Convenience [TD03]. Convenient [KMG06]. Conventional [FKL^{+}03, KS10, Opd03, SY19]. Convergence [DEH^{+}06]. Convergent [WFC15]. Conversation [Saw04a]. Converting [Wal03]. Cook [BPZ12]. Cookies [Han03]. Copula [BPZ12, PDC13, ZC13]. Coronary [Cen04]. Correct [Kan05]. Corrected [BB19, Ola20]. Correcting [KRP02, MEH08, ML03]. Correction [Die09a, Die11]. Corrections [LS06]. Correlated [Hea02, Lip17b, OA11, Oya11, Pap16, RJR14, SW06, SSK13, SD17, SAA12, Tig03]. Correlation [AL15, AKP08, BS11a, BF09, BF11, BF13, BF14, BG15, Bha14, BKS13, CC09, Con02, DP21, Ele17, Gid07, HW11, Hit09, HK10, Hus07, LeB16, LH14, Lip09, Mak17, NXFN19, Per06, Rus08, SHA06b, Sha08a, SP14b, Wal17b, WS16b, Wil15b, Wil19e]. Correlational [Wal03, WFW02]. Correlations [BH06, Lip18, Wil19e, Bea17]. Corresponding [WC09]. Cost [Cox13]. Count [Doy09, Gil24, MEH08, PM03]. Countries [Fer12]. Counts [CS08, KC06]. Covariance [Bha08, CFSF10, GZ07, HSA18, MK05, SLLL11]. Covariances [LHB04]. Covariate [AAI15, AE11, ICS07, LMME09, NS07, Tig03, Wil19c].

Covariate-Adjusted [LMME09]. Covariates [DDM07, GVCM18, MRKM^{+}17, SRW21, Wil17b, Wil19f]. Coverage [HM08, MC07, MC08, WC05, Wil17b, Wil17a, ZA08]. Cox [MRKM^{+}17, Tig03]. Cox-Type [Tig03]. Crashes [Kor16]. Created [RH03]. Credible [GHSF^{+}15, Lov19a]. Criteria [DP05, Fuk07, GD02, GG09, Lip10, PD05, Wat06, WF09]. Criterion [EAS13, Gha14, Kra18, Lip06, UAD16]. Critical [And13a, And13b, BB04, BB05, Fay02, Hea03, Per03, PB13, Raj21]. Criticism [Ser02]. Cronbach [EYS^{+}07]. Cross [AM17, CKPJ18, HK10, SH15, YD17]. Cross-Calibration [AM17]. Cross-Classified [CKPJ18, YD17]. Cross-Correlation [HK10]. Crossing [TZA08]. Crossover [Ree06, Ree11a, Ree11b, Ree12]. CSF [XGY^{+}08]. cT [KIR12]. cum [VZ20]. Cumulative [EAM09, HNG19, Wal16b]. Current [DS02, SSS16, SZH^{+}18]. Curriculum [LP06]. Curse [Wil18c]. Curvature [RNN13b, Wil08b, Wil16a, Wil17b, Wil18a, Wil18c]. Curve [Bak07, PA14]. Curves [MJ20, SR07]. Cycle [ZT12]. Cyclic [GDG20]. Cyclical [SS18].

D [MADT03]. D’Agostino [PS02b]. Daily [SIA11]. Dangers [Ber05]. Data [AE12, AM17, AGHL08, AK18a, AKA20, AKO03, AK04b, Alo06, ACP05, AE11, AA14, ANA12, BPWR10, BAN06, BK09a, BS04, BZBS03, BBB12, BI02, BS11a, Bha10, Bha14, BII11, BLNA^{+}22, BE17, BO13, Che14, CTDVD05, CGH15, CWBK07, DS02, Dem04a, DGJ16, Doy09, Dub04, EO13, Eic05, EB04, EAS13, FBR^{+}08, FC09, FEB13, Fin11, Fin14, FRJ02, FR04, GV03, Gil24, HPSR23, HMNW07, HL16, HKP17, HA14, Hea02, Hus07, IS24, IJ06, KOW16, Kit10, Kna19a, KZ08, KCA^{+}17, KJ06, LH14, LFO07, Lee11a, Lee11b, Lee15, LB10, LIT24, LX06, LC05, LEKG10, LK12, LK13, LOCMS13, LBC^{+}19, LAMB17, LRT20, Lor19, MRMH17, MF11, MEH08, MDS03, MOD07, MMS08, MPS11, MK10, Mos12, MLS10, Nan05, Nan02, Neu09, NM17, NIN16, OD07, OAC03, ODLR17, PMZ08, PPS14]. Data [PNG08, Pau05, PN10, PM10, PC18, Pér07b, PTC18, PI08, PC14, Pra10a, Pra15, QWL^{+}16, RYZ16, Rah17, RJR14, RNN12, RNN13b, Raz13, Rus08, SAA19b, SLLL11, SK14, SKB15, SK09, SSMS10, SS24, SFA16a, SFA16b, SSK13, SD17, SIA11, SSMK22, SZH^{+}18, SBY07, Su05, SLA16, TK06, TSK04, Tor18, WS19, WA03, WBSMI09, Xu05, YHPR17, YII13, YD17, ZS17, Zha05, ZCB^{+}17, ZLB19, Zum02, bZbWA13, ANO^{+}25, FKP^{+}25, Koc18, Opd05, SKSV25]. Datasets [CD09, MC11, RB09]. Daubechies [NSSO14]. Day [ANA12]. Days [RR24]. Deal [LHH11]. Dealing [HS18, Saw13, Wil15b]. Death [PM08]. Decision [Cam09b, WB06]. Decision-makers [WB06]. Decisions [Cen04]. Deck [Pér07b]. Decomposition [LF11, Lip16, YT10]. Deconstructing [Saw03a]. Default [Kel05]. Defects [RB11]. Defense [Kna15]. Definitions [ZZ15]. Degradation [DEE17]. Degree [PA07]. Degrees [Web07, WA07]. Delete [Ela08]. Deletion [SV12]. Delphi [SK09]. Demand [MEH08, PK06]. Demonstration [Ber17, Lor19]. Density [Aly07, Bic11, EME10, HNG19, MMM07, RC17]. Departing [LIT24].

Departure [IT18]. Dependence [ACB06, Fer17, SH15]. Dependency [ZC13]. Dependent [AE11, CFSF10, GJ17, ICS07, Lip12, MS14, MBS17, Wil19e]. Depth [Gao04, MA17]. Derivation [Par10]. Derived [NLQ10]. Design [AOCAN19, ASG16a, ASG16b, BW05, Cox13, Dye16, EK03, Gan12, Han03, HKP17, HBA20, JPH23, Les11, Lui10, PA04, PA07, PMZ08, RJR14, Ree11a, WY03, WA03, Wil06a]. Designing [Rah14, SU16b]. Designs [ATvBB17, ACBM06, AKO03, AK04a, ACP05, BZ09, CP13, CKP12, DBH02, EK03, GA17, GDG20, Goo09, HS18, JVA10, KY08, Kon13, LHB04, LFG14, LFG19, LK13, LH04, Lu16, MN05, MN17, MC11, MCL09, MK05, MBWZ07, OO18, QS12, Ree11b, Ree12, SL17, SM17, SB09, TTBG06, TW07, VJ11, WA07, SG21]. Destructive [SU16b]. Detecting [RB09]. Detection [AE12, CKCL18, EAS13, HMNW07, LM17, MRMH17, NSJ^{+}24, SL16, WSN15]. Determinants [CPDP18]. Determination [Rav17, SH14, SP16, TTBG06]. Determining [Kan05, Kan06, Wal05a, Wal05b, Wat06, WFW02, Wil18a, Wil19e, YTP24]. Developing [Fer12, JR16, MMM07]. Development [WA07, WZS11]. Developmental [CBP18, LH09, LBC^{+}19]. Deviates [Oya11]. Deviation [Aly07, KI10, OMN16, TK06]. Deviations [Lip07, Nav07]. DEWMA [Alk13]. Diabetes [KS25b, MMM07]. Diagnostic [HSA18]. Diagnostics [NAK13, OL11]. Diagonal [DS16]. Dial [Lee12]. Dialogue [Ber09]. Diameter [Oya11, OCB13]. Diarrhea [RRB08]. Dichotomized [NM17, NM19]. Dichotomous [Fin11, GVCM18, PM10, TS16, ZD10]. Diet [PH25]. DIF [GZ07]. Difference [BC02, LW05, LZ07, Lui06, Ree09b, RL02, Saw02b, SB09, WWB12, Wil18d, Whi07]. Differences [Gao04, LF11, MK10, Tur17, Wal05b, ZK05]. Differencing [MDS03]. Different [AAH^{+}19, FF16, KJ12, Lu16, OHQ^{+}20, PNG08, RSQ20, SKB15, Saw03b, Tho09, Web07, WZS11, YHPR17]. Differential [DSCS08, LKW^{+}19, LH13]. Differentiating [Bos09]. Dilemma [BB16]. Dimension [TZA08]. Dimensional [BBB12, Fin14, Ura20]. Dimensionality [Wal04, Wil18c, YII13]. Dimensioning [NM19]. Disconnect [LS06]. Discontinuity [Les11]. Discovery [OE12]. Discrepancy [HPSR23]. Discrete [AM19, BK09a, HZ16, Mos12, MBS02, PJ14, ST14]. Discretized [Su05]. Discriminant [Bos09, BLNA^{+}22, HM10, LC05, MN21, RRHA11, SLLL11, Wal04, WFC15]. Discriminating [RK13]. Disease [Cen04, DS02, Sie05, XGY^{+}08]. Diseases [AA12, RS17]. Disjoint [Hur03]. Dispersed [ANA12]. Dissimilarity [ÇE14]. Distance [Bha14, Bos09, HSA18, KJMA18, MRMH17, Mug12, PNZM24, RSQ20]. Distinction [And13a, And13b, PB13]. Distributed [AO08, JR16, Ola20]. Distribution [ADSBH02, ADARAM05, AS10, AA12, AC15, AAA15, Ahm20, AJ17, ANM15, AALF14, AHAO09, AM10, AOAH11, AOCAN19, ASS04, ASAIJ12, AW05, ALF16, AASIJ10, AFL12, AK13, AGS14, AGKJ19, AEOD24, AA20, AT09, AM19, AF23, AŞ15b, BZBS03, BJJ22, BJJ23, BK22, BR03, CCT09, Cam03, Cam07, Cam09a, Cam09b, Cam12, CdlTE22, DEE17, DGJ16, Ebr06, FLE04, FA11, Fay02, Fay06, FA13, FZ09, Geo14, Gil24, GRM20, GZ17, HA07, HB08b, HB12, HHN20, Hit09, HNG19, HZ16, Hur03, Hur09, IS24, JA18, JAJ24, JPH23, KSS10, KRR13, KRR14, KPR14, KR16, KMA18, KJ12, KJ14b, KK13, Kib08, KLS07, KDMAA18, KD18, KKK19, KMA21, KMG06, LSAA02, Lec07, LFO07, LL16, LAMB17, MJ17, MESSE16, MA24, MSS20a, MEP09, MZ17, Mir08, MA18, Mug12, MK23, MŞ09]. Distribution [NAA16, NH12, NE16, NAY^{+}18, OO05, OHQ^{+}20, ÖK17, Ots05, OGKH17, PPS14, PJ14, PDC13, Pas09, PJ07, PS22, Pau05, PN10, PTC18, PS17, Pra10a, Pra10b, Pra12, Pra20, RS20, RMA14, Rao11, RRB12, RK13, Rao14, RD16, RRS23, RJ14, RAJ18, RWS09, RNN12, RC17, SMSN20, SQ15, SLD14, SMC16, SS23, SHA06b, SF19, SA14, SFA14, SFA16b, SSMK22, SA12, SA13, TP16, VA16, Web07, ANO^{+}25, AA21, AK21, OON24, SKSV25]. Distribution-Free [Fay02, Fay06]. Distributional [BZ09, WFW02]. Distributions [ADAR09, AAK10, ASK16, AKA20, AFL19, BB06, BK09a, BC19, DP21, Dem04b, Dem05, Die09b, ESAH20, FS15, FZ09, Gan14, GR11, Hea04, Heo02, HBA20, KJ12, KS25a, KH16, LA18, LS05, MN21, Mos12, Ng06, NKdG^{+}16, Ozg16, RNN13b, SFA16a, Su05, Su11, TZA08, TGhC^{+}05, VA16, WS16b, Wil19b, YOO14, ZJ08]. Disturbances [SW06, Saf11]. Divergence [SSM18]. Division [Ozg16]. Do [Saw04b]. Doctoral [Hal04]. Does [BH16, BHH12, TSS20]. Domain [AKK24, Res08]. Domestic [LIT24]. Dominance [FC09]. Dose [LH09, XYJ03]. Dose-Specific [LH09]. Double [AS17, ADSBH02, AA14, CERA24, Hur09, KSB10, PS18, PP17, Rao11, SMSN20, ST02]. Doubly [LAMB17, SFA14, SFA16a]. Down [MSS20a]. Downside [KI10]. Drainage [SJ25]. Drawn [Kho07]. drill [NT25]. Driven [BH16, Saa19a]. Drivers [LC20]. Dual [KC20]. Ductal [TX11]. Due [HS11, Sha11, Pér07b]. Duration [MEP09]. Durbin [BR14]. Dwelling [MPC^{+}16]. Dyadic [BKS13]. Dynamic [MMS08, NXFN19].

E-Bayesian [NE16]. Easily [FZ02]. Eating [LHO^{+}18]. Ecological [BKS13]. Economic [AOCAN19, Cox13, Mar12, SSMS10]. economy [PW24]. EDA [KS25b]. EDF [ASAIJ12]. Editor [Con18]. Education [FS15, Ste16]. Educational [Boo19, Che14, LK12, LK13, Liu14, OD07, Xu05]. Eether [Kes15]. Effect [AKP06, CP13, EN23, Fay06, Hit09, KAF05, LR03, LT07, MSS07, MEH08, MA06, Mec03, OYO16, OL03, Saw09, Sie05, SK10, Ste16, Wal05b, Web07, Wil18d, YTP24, ZA08, ZA11, ZRB23, NT25]. Effective [NKC18, RN18, SKB15, SS24]. Effectively [Tur17]. Effectiveness [CID^{+}14, HM10]. Effects [AKO03, AK04a, AZN20, CKCL18, Ele17, GCAC07, HMNW07, HL16, Hit09, KWOF02, KKZB16, LLHT03, Mak17, PC15, RYZ16, RH03, SLLL11, Tho16, Tig03, WY03, WBN07, ZA08, ZA11]. Efficacy [WPN^{+}18]. Efficiency [ASG16a, Bos09, CEI24, Dye11, NW10, SW06, Saw02c]. Efficiency-Balanced [ASG16a]. Efficient [CDPT19, CERA24, Dan25, MBS16, OAC03, SK18]. Eggenberger [HA07]. Ehrenberg [Lip13a]. Eigenvalues [Heo02]. Eight [Wal07]. Einstein [Saw02b]. Elastic [AL15, HWS^{+}24]. Elastic-Net [HWS^{+}24]. Electorate [EG11]. Elemental [HM08]. Elementary [TTBG06]. Elicitation [CPT04]. Elimination [RR07]. Elliptically [BZBS03, Kib08]. EM-Algorithm [SSMK22]. Email [Die09a]. EMB [Tak17a, Tak17b]. Emergent [Weh03]. Empirical [AK10, AASIJ10, Bak07, Ber17, NR05, NKW^{+}19, Odi12, SKB15, SRA17, Wil10]. Endpoints [BE17]. Energy [Saw02a]. Enhanced [KS25b, SS24]. Enhancing [KN05]. Enough [BBW11]. Enterprises [Mar12]. Entropy [AO15, ÇE14, DW19, Lip06, RC17]. Envelopment [YII13]. Environmental [MLS10, RRB08, RJR14]. Enzyme [Bos09]. Epidemiologal [TSK04]. Epidemiological [Alo06]. Epileptic [SHA06b]. Equal [AKP06, CEI24, RH03, MC07]. Equality [BK09b, BF11, BF14, BG15, CWBK07, KOW16, PS02b]. Equating [CKCL18]. Equation [CE15, CB12, FB10, Oya11, RM16, RMS17, Koc18]. Equations [Im20, Saf11, SM04]. Equity [LF11]. Equivalence [BC02, GCAC07, Mec03, Neu10]. Era [Lov19b]. Erlang [KJ12, TZA08]. Error [AB11, AOY11, ANM15, AO15, And13a, And13b, AO08, BM05, CMR02, DK19, Die09b, Dye11, FA11, Fay06, FKL^{+}03, Fra15, Han03, Har19, Kes15, KM02, KK20a, LS06, LB16, LHH11, MŞ09, Nav07, OE12, PA04, PA07, PMZ08, Pas09, PB13, RBAR18, SLLL11, SS12, SLC17, SM17, SW08, SPT12, TMB15, WA03, WA07]. Errors [Bha10, Die06, Die11, Dor16, Fra17, IB24, Lip13a, OA11, RBAR18, SS13, SK10, SS15, SY19, SAA12, SJAZ14, YTP24, Wal17a]. Eruptions [BPZ12]. Esscher [DGJ16]. Establishment [Pér07a]. Estimate [BS04, DEE17, GVCM18, LF11, LK12, OR02, OAC03, PM03, ZRB23]. Estimates [AA14, Bea11, CB12, Ela06, HM08, JPH23, KJ14b, MEH08, Pér07a, SA12, SA13, TSS20, Wan03, YOO14]. Estimating [AZN20, BS09, BKS13, Bou10, CGH15, Dan25, EAN05, EAM09, Im20, KSS10, KS25a, KKZB16, Lan09, Leh15, Lor19, MEP09, PZ11, PS22, Qum13, SSS21, SSMK22, TS16, Wal08, Wil08a, Wil15a, KC20]. Estimation [ADSBH02, AC15, AM14, AJ17, AN04, ASS04, AA21, AAH^{+}19, AA20, AM19, AŞ15b, BP16, Bha08, BF09, Bic11, BK22, BE17, Cha22, CBP18, CC09, CS12, CCH19, CERA24, DSK18, DP21, Die09b, Die11, Dub04, Eic05, EB04, Eid04, EME10, kAEAE21, Eni09, FA13, FF12, FRJ02, GZ07, GR11, HM08, HL16, HS08, IS24, KRR13, KRR14, KS14, KMA18, KI16, KJ12, KJ14a, KV14, KMA21, LHM14, Lui06, MBS16, MN05, MC11, MPC^{+}16, MAP08, MMM07, NE16, NAK13, Ola20, OO18, PD13, PS18, Pan18, PPC18, Pap16, PJ07, PTC18, PI08, Pra10b, Pra14, Pra20, Qu05, QT06, RMA14, Rao14, RWS09, Raz02, RC17, SMSN20, SAA19b, Saf08, SQ15, SS16, SS24, SFA14, SKB11, SS15, SVS16, SSS16, SB09, SZH^{+}18, Su11, SK13, Sub13, SP14b, SRA17]. Estimation [TS11, TP16, WBSMI09, XZ05, YTP24, YAS14, ZT06, MESSE16, Whi07]. Estimations [FEB13, HNG19, PN10]. Estimator [AOBCP16, Aly07, AW16, AKK24, Ben17, Dor16, Dye11, EAS13, Eni09, Fer17, GSP24, Gil24, HSA18, Kad16, Kha12a, Kha12b, Kho04, KC20, MKS19, MMH19, MKA23a, MKA23b, Mug12, NW09, NLQ10, OMN16, ÖÇ17, PS18, SPT12, Sie05, SGM16, Sub13, TC12, VZ20, Wal17b, Wil07a, YTP24, ZT06]. Estimators [ADAAM09, ARM10, AOAH11, AO15, Asa16, AO08, BE06, CA08, CEI24, GCP24, Geo14, IB24, JK14, KWAO04, Kha13, Kib06, KB16, KSB10, Kum13, LAA17, MA06, MSS20b, MT24, MŞ09, OHQ^{+}20, RSQ20, RTA22, SK14, ST02, SS04, SA10, SS13, SS24, SK10, STJ11, SK18, SY19, SJAZ14, SR07, SP14a, SP14b, TL14, VS08, WK04, Wil11, WC13]. EU [TL15]. Euclidean [PNZM24]. Eurasian [DEH^{+}06]. Evaluate [GCH08, Pér07b]. Evaluating [Dye16, NM17, TX11, WPN^{+}18]. Evaluation [Boo19, FC04, Fin19, FB10, Kle19, Les11, Mey16, MK08, NR05, Nan05, Pér07a, PA14, Saa19a, TNH17, VA16]. Évariste [SC06]. Even [TuAH19]. Event [CGH15]. Events [RNN13a, SJ07]. Evidence [AS16, BB07, Fuk07, HPSR23, OL03]. Evidence-Based [AS16]. Evidences [YS16]. EWMA [AS17, Alk13, MSS19]. Exact [BB04, BB05, Ber17, Bha10, BF11, CM02, Gan12, Hea03, Lee11a, Mal02, OO05, Pos02a]. Examination [Per03, SV12]. Examinee [GCH08]. Examining [EYS^{+}07, Kor16, OE12, SE12]. Example [KHC05, PW04, WY03]. Examples [AKA11, Doy09]. Excel [Hea06, Hei06, MS14, SV08, Sta02]. Exchange [PDC13]. Exchangeability [Goo02]. Existent [KSS10]. Expected [CCT09, CC09]. Experiment [OO05, SM17, Con18]. Experiment-wise [SM17]. Experimental [CKP12, Goo09, Han03, HKP17]. Experiments [BO13, GMS05, Per06, Qum13, RB09]. Expert [Leh15]. Explanatory [MIN18, TZZD08, Wil08a]. Explicit [Saf11, SJ25]. Exploration [Heo02, RNN12, SWM19, Xu05]. Exploratory [FF08a, FF08b, Kel05, LA14, NM17, NM19, PM10]. Explore [LBC^{+}19]. Exploring [CCC19]. Exponential [AE12, AS10, AC15, AM10, ACB06, AGKJ19, Bak03, BJJ22, BJJ23, Cam07, CEI24, Die06, ESAH20, GSP24, GRM20, HNG19, Hur09, Kad16, KPR14, KS14, KDMAA18, KD18, KMG06, LIT24, MJ17, MKA23a, MN21, NAA16, Ng06, OCB13, PP17, PN10, Pra12, RMA14, RK13, RWS09, SQ15, SVS16, SY19]. Exponential-Geometric [KDMAA18]. Exponentially [KS06]. Exponentiated [HA14, HHN20, NAY^{+}18, RRS23, SQ15, SA14]. Exponents [Nav07]. export [PW24]. Exposure [EAM09, FB04, Gan12, KRP02, Lee13]. Exposure-Risk [KRP02]. Expression [DGJ16, MPS11]. Extendable [HH16]. Extended [ANO^{+}25, GVCM18, KD18, MJ17, MRKM^{+}17, MK23, RMA14, Rao11]. Extending [Ahm20, ZL04]. Extension [Day17, HHN20, KE12, NSSO14, Ree11a, SAA12]. Extensions [Goo02]. Extract [Kan05]. Extraction [LM17, Tho10]. Extreme [ADAR09, AOAH11, ASS04, CdlTE22, KS25a, RRB12, SS04, SJ07, Wal07, ZJ08]. Extremely [CDPT19]. Extremes [AHAO09]. Eyether [Kes15].

Face [PNZM24]. Factor [Bea11, BS14, Bea17, FF08a, FF08b, Fin11, FF16, HH16, KS09, KY08, Kel05, KZ08, Lee15, LCW^{+}17, Lip17a, LM19, NKW^{+}19, NM17, NM19, PA04, PM10, RWS09, SPT12, Wal04, WWB12, WZ17, YTP24, ZK05]. Factor-Type [SPT12, YTP24]. Factorial [ACBM06, ASG16a, FF16, KY08, QS12]. Factors [JVA10, Lip17a, Lov19a, PD11, RRB08, WW10]. Faculty [LF11]. Failure [AK18a, AE11, BS04, DEE17, FRJ02, HA14, KPR14, KR16, Pra10a, ANO^{+}25]. Failures [AE11]. Fallibility [KBKBG07]. False [OE12]. Families [Weh03]. Family [AK20, AAK10, AKA20, AFL19, BF09, BF11, BF13, BF14, ESAH20, Fra15, HNG19, KRP02, Kes15, LHO^{+}18, Pra10a, SHA06b, TMB15, VS08]. Familywise [Kes15]. Fano [SA07]. Fast [Opd03]. Fatal [HML07]. Favor [RL10]. Feasibility [LHM14]. Feature [LM17, Ura20]. Features [KHC05]. Fecundability [Pau05]. Feed [LSAA02, Lan09]. Feed-Forward [LSAA02, Lan09]. Feedforward [KK16]. Female [Lee12]. Fermat [Saw02b]. Ferrieri [Fer12]. Fertility [Pos02b, PM03]. Field [DS02, RNN12]. Filter [EO09]. Final [Saw03c]. Financial [EN23]. Finding [Bak07, LC20]. Finite [Kho07, Kum13, LHH11, SS13, SK18, Sub13, SP14b, TS16, VZ20, DSK18, Kna19b]. Firms [XZ05]. Firth [Mey16]. Fisher [Ser10, Cam07, HL03, Lee11a, Saw02b, Wal17b]. Fit [ASAIJ12, AASIJ10, BH16, CE15, DW19, KS09, LT07, MJ20, Nee04, NKC18, ÖK17, PD11, Pau05, PTC18, RMS17, Saw02a, SRW21, Su05, Zha05, ZD10]. Fits [OGKH17]. Fitted [HL03]. Fitting [GV03, Liu09, LK13, Liu14, Pan18, Su16a]. Fixed [BW05, ZA08, ZA11]. Fixed-Effects [ZA08, ZA11]. Fixed-Sample [BW05]. Flanagan [Wal06]. FLASHE [LHO^{+}18]. Flaws [HML07]. Fleishman [Smi09]. Flexible [JADB10, LA18, PM14, Su16a, Wil07b]. Flexibly [Su05]. Flights [LIT24]. Flood [BB11]. Florida [OGKH17]. FMDFB [LS16a]. Follow [MBS02]. Follow-up [MBS02]. Follows [SHA06b]. Food [DNUH22]. Forcasting [ST07]. Forecast [DAC02]. Forecasting [LIT24, OYO16, OAA10, SSMS10, ST08, WW10, PW24]. Forecasts [SS18, YOO14]. Forests [CPDP18]. Form [HH16, LM19, Ree09a, SLD14]. Format [KZ08]. Formative [WFC15]. Forms [Abe15, RC17]. Formulas [Wal06, Wal07]. Fortran [Fah02a, Alo04, Bra03, FC09, SH07, SH09]. Fortune [Han03]. Forward [Jay17, LSAA02, Lan09, LB19]. Four [AAH^{+}19, BM05, Fay02, Kan06, MSK^{+}11, Ree12, Tho09, WA03]. Fractional [ACBM06, KY08, Pér07b, QS12]. Fractionally [MDS03]. Frailty [BS04, FRJ02]. Framework [GK04, Kor16, Pan18]. Frane [Wal17a]. Frechat [AA21]. Fréchet [ANO^{+}25]. Free [Fay02, Fay06, LS13, Sha11]. Freedom [PA07, Web07, WA07]. Frequency [BB11, Cli11, Lip19, Res08]. Frequentist [GHSF^{+}15, PW04, SAA19b]. Friedman [BB05]. Full [GMS05]. Full-Profile [GMS05]. Function [AE12, ADSBH02, AASIJ10, BD20, Bos09, BW05, EAS13, Fer17, GK04, GVCM18, HNG19, Mug12, Pas09, RWS09, Raz13, SLLL11, Saw02a, Udo13, Udo20, WCW08]. Functional [AN04, RJR14]. Functioning [LKW^{+}19]. Functions [AGKJ19, HR16, KDMAA18, LL16, MA17, QT06, RRHA11, SWM19, Udo13, Wal05b]. Fundamental [Gau17]. Further [CCC19]. Future [GJ17, Saw03b, WB06]. Fuzzy [ANAS16, PPS14, SKO09, SPB19, SK11, Tho16, SKSV25].

GA [WW10]. GA-Based [WW10]. Gain [QWL^{+}16]. Gained [BB07]. Galois [SC06]. Gamma [AFL12, AK13, AGS14, BAN06, BS04, KS14, KLS07, MZ17, Ola20, RNN13b, SLD14]. Gamma-Half [AK13]. Gamma-Logistic [AGS14]. Gap [MBS02]. GAR [KC06]. GARCH [Eni09, MSS07, NXFN19, OYO16, YOO14, YS14]. Gaussian [Cam03, Cam09a, Cam09b, Cam12, SJAZ14]. GDP [SSMS10]. GEE [ZT06]. Geeta [RJ14]. Gender [HPSR23, LF11]. Gene [DGJ16, MPS11]. General [BP16, Che14, CAM08, Gao04, KSB10, LM19, Pap16, Tur09, WZS11]. General-To-Specific [Tur09]. Generalizability [CF03]. Generalization [ALF16, HA07, RS20, SS23, Wil12]. Generalized [AA12, AKA20, AFL19, AEOD24, AA20, AKK24, CFSF10, FA11, GV03, HSA18, HMNW07, HA14, Im20, KRR13, KRR14, KMA18, KOW16, KJ12, KJ14b, KMG06, Lip16, LK12, MA24, Mir08, MA18, NAA16, PJ14, PN10, PTC18, QS12, RS20, RK13, Rao14, RJ14, Raz13, SS23, SS13, SSS16, SAA12, Su05, Su11, Sub13, VZ20, Wil06b, Wil04]. Generalizing [Wal16b]. Generated [AGKJ19, CPT04, Pre04]. Generating [AGKJ19, CD09, DiS02, ESAH20, Hea03, KDMAA18, KMG06, MRKM^{+}17, OO05]. Generation [Dem04b, Dem05, GT10, MK10, Sta02]. Generator [HS11]. Generators [Fog15, Mar03]. Genetic [AM17, CMR02, SP16]. Geometric [CCT09, KJ14b, KDMAA18, PJ07, Pau05, RMA14]. Geometry [DDM07]. Geostatistical [MLS10]. Gibbs [SH07, SH09]. Gill [BS04]. Give [Kna18]. Given [Wil19c]. GLDreg [Su16a]. GLM [SK14]. Global [Aly07, Wil16a]. Glossary [LBC^{+}19]. Goals [Sen02]. Goldilocks [BB16]. Gompertz [FA13, KKK19]. Good [KOW13, Lip13a, OL03]. Goodness [ASAIJ12, AASIJ10, KS09, MJ20, Nee04, ÖK17, Pau05, PTC18, SRW21, Zha05]. Goodness-Of-Fit [PTC18, Nee04, SRW21]. Got [Saw03d]. gradient [FKP^{+}25]. Granger [MSS07]. Graph [SJ25]. Graphical [BBB12, Gau17, LA14, SV12]. Graphics [CAM08]. Graphing [Tho16]. Graphs [ACBM06]. Greatest [Ben17]. Grids [Wil19f]. Grizzle [Ree11a]. Group [AKO03, AK04a, CWBK07, Fin14, FF16, KAF05, KOW16, KZ08, LHB04, MC08, SL17, Wal15, ZK05]. Grouped [BAN06, PN10]. Groups [Bos09, MCL09, RRS07, WA03, WA07, WK02b, Wil12, WH16]. Groups-Versus-Individual [WA07]. Growth [CCH19, HB08a, Jay17, LHM14, Mar12, OCB13, PKRK25, SE12, WZS11]. Guidelines [ACP05, Che14, MRKM^{+}17]. Gumbel [AALF14, AŞ15b, SFA16b, YAS14]. Gupta [HS18].

Hadamard [DSCS08]. Haghighi [AEOD24]. Half [AK13, KRR13, KRR14, PS17, Wal06]. Half-Logistic [PS17]. Handling [Che14, PC18, bZbWA13]. Harris [ANO^{+}25]. Hasanuddin [LIT24]. Having [OS10]. Hawai‘i [Min13]. Hazard [AAI15]. Health [IS24, KKZB16, LEKG10, LHO^{+}18, Min13, YHPR17, bZbWA13]. Healthy [SHA06b]. Heavy [MSS20a]. Heavy-tailed [MSS20a]. Hedges [VO03, Wal15]. Hegemony [Kel05]. Height [Oya11, OCB13]. Height-Diameter [Oya11]. Height/Diameter [OCB13]. Hellinger [Mug12]. Helpful [Die11]. Heterogeneity [BZ09, GS19, MA06, NKW^{+}19, RP03a, RP03b, WPN^{+}18, XGY^{+}08]. Heterogeneous [BKS13, LHB04, MEH08, Udo13, Udo20]. Heteroscedastic [CWBK07, Dor16, MCL09, MAP08, RBAR18]. Heteroscedasticity [CCC19, GCAC07, HSA18, KWOF02, Saf11, Wil15b, Wil17a]. Hettmansperger [NS07]. Hidden [BA11, NAK13]. Hierarchical [BP16, ÇE14, FBR^{+}08, LP06, PM14, SM17, WF09, ZCB^{+}17]. High [BBB12, Fin14, LEKG10, LHO^{+}18, MW07]. Higher [BII11, CS12, Ree11b, SIA11]. Hinge [SR07]. Hoc [HM10]. Hoel [Wil19a]. Homeless [MPC^{+}16]. Homogeneity [KM02, LCW^{+}17, VO03, XYJ03]. Homogeneous [OR02, ORSCMG03, TX11]. Homoscedasticity [YSVS19, Wil07b]. Honor [Sto04]. Hospital [RRB08]. Hot [BPZ12, Pér07b]. Hotelling [Nan02]. Huberty [Hit09]. Huffman [SA07]. Human [Pau05]. Hurricane [OGKH17]. Hybrid [KJ06, KJK09, LRT20, Pra20]. Hyperbolic [OCB13, Udo13]. Hypercube [JPH23]. Hypernatreamia [KKST05]. Hypothesis [BH16, CDPT19, Die09a, Die11, GK04, GT10, Har10, Hei06, LS16b, Lui10, Mec03, Nav07, OY17, Ozg16, RL16, RL10, Saw03a, Saw16, ZK16].

Ian [Lui07]. Ideal [PK06]. identically [JR16]. Identification [CC19, LWZK16, MF11]. Identify [RTS16]. Identifying [LA14, SK11, Wil18b, Wil19b]. II [BP16, FEB13, GRM20, KRR14, KJ06, KJK09, Lui10, OON24, PPS14, PP17, PS17, Pra14, SFA16b, TX11]. Ill [KI14]. Ill-Conditioning [KI14]. Image [HR16, LM17]. imbalanced [FKP^{+}25]. Impact [AZ17, CKCL18, CKPJ18, DAC02, Fan10, Lan13, LeB16, LB16, LLB17, SS09, WY03, WWB12, YD17, PW24]. Impacts [BB16, LM15]. Implementation [BE17, DNUH22, OMN16]. Implemented [FZ02]. Implementing [SH07, Tak17a]. Importance [Bra03, Har19, LZW14, OL03, SL17, TZZD08, WFW02, WZ17]. Important [KY08, Wil18a]. Improve [Kha12b, OS10, PE08, SS18]. Improved [AKK24, Dor16, FC04, Kha12a, Kib06, Kum13, LFG14, LFG19, MS03, Ree09b, SSS16, TS11]. Improvement [MKA23a]. Improvements [Eid04]. Improving [EM14]. Imputation [CTDVD05, DMBR16, FB10, LB10, LMRM^{+}20, LHO^{+}18, LOCMS13, MLS10, Pér07a, Pér07b, SVS16, SSS21, Tak17a, Tak17b, ZCB^{+}17, Koc18]. In-Sample [OYO16]. Inappropriate [YD17]. Incidence [Kit10]. Include [DRTW17, DWT19]. Inclusion [WB02]. Income [MJ20, Weh03]. Incomplete [Dem04a, PI08]. Incorporating [CF03, Kho04, Kho05, SLC17, WW10]. Increasing [GS19]. Increment [OCB13]. Indefinite [RC17]. Independence [CPDP18, CS09, JADB10, PD11, Sha08b, Wil07b, Wil08b]. Independent [ADARAM05, AM10, BZ11, DRTW17, Fay02, FC09, FKL^{+}03, HXH06, HK10, HH16, KRP02, KHC05, LAM16, LFG19, LT07, MCL09, NT25, Par10, RYZ16, VM09, WS09, Wil12, Wil18a, Wil19b, Wil19e]. Independent-Samples [FKL^{+}03, LFG19]. Indeterminacy [Bea11]. Index [BR03, Fer12, OAA10, PD11, ZD10]. Indexed [KIR12, SS12]. Indexes [BP16]. India [BB11, PKRK25, RRS07, VA16]. Indian [PH25]. Indicators [HB08a, Koc16]. Indices [CE15, PDC13, RMS17, SV08, WS16a, Wal17b]. Individual [Lip19, Rah14, TSK04, WA07]. Individuals [AS17, KY06, KQC06, WA03]. Indonesia [DNUH22]. Inductive [Hur09]. Industrialization [EO09]. Industry [PE08, Wan03, ZT12]. Inequality [MJ20]. Infant [HPSR23]. Infection [RRS07]. Inference [AAK22, AHAO09, AW05, AK21, BY11, Bar15, BII11, BH06, DR14, Die09b, GRM20, KSM13, KM14, Kib08, KK20a, LRT20, Lov19a, Ng06, OA11, PPS14, RD16, SZH^{+}18, WK04, ZC13, Zum02, ANO^{+}25]. Inferences [HW11, NH12, WE05, Wil06b, WC09, Wil10, Wil15b, Wil18d, Wil19c, Wil19d, Kna19b]. Inferential [LAMB17, RRS07, Wil18a, Wil18b]. Inferiority [GT18]. Inflated [Doy09, PPC18, SAA19b, YHPR17, bZbWA13, PM03]. Inflation [AB11]. Influence [JPH23, Kan06, MK10]. Influencing [PD11]. Information [ADAAM09, BY11, BB07, Cli11, Day03, DP05, EAS13, Fuk07, GD02, Gha14, ÖÇ17, PS18, PD05, Per06, SK18, SPB19, UAD16, VS08]. Informative [QWL^{+}16, Seo19]. Inhibition [SA02]. Initial [HS11]. Initiatives [RB11]. Innovations [YS14]. Instability [PKRK25]. Instantaneous [GVCM18]. Insurance [IJ08, LA18, ZT12]. Integer [HB12]. Integrated [Dye11, MF11]. Integration [CT05]. Integrity [Fan10]. Inter [Bea17, BS09]. Inter-correlations [Bea17]. Inter-Rater [BS09]. Interaction [ACBM06, Pet02]. Interactions [BZ09, KY08, LHB04, MK05, Saw13, WE05, Wil19a]. Intercept [BY11, Zha05]. Intercepts [LMME09]. Interest [DMBR16]. Interface [CAM08]. Interference [AA12]. Intermediate [Smi09]. Intermittent [QWL^{+}16]. Internal [KS09, PZ11, SSAH18]. International [LIT24]. Interpretability [RRHA11]. interpretation [Con18]. Interpretations [GJ17]. Interpreting [LR03]. Interrater [Heo08]. Interrupted [RB09]. Interval [AS10, ASS19, AE11, AA14, BF09, Bha10, BS02, FC04, LAMB17, Lui06, MC07, MC08, PD13, Whi07]. Intervals [ASS19, AKP06, AKP08, Alo04, BD16, Bak03, BAN06, Bak07, Cam03, Cam07, Cam09a, Cam12, CA08, CP13, CPT04, DP21, Ebr06, Fra17, GHSF^{+}15, HM08, KAF05, Kin03, Lec07, Lov19a, PS17, Pre04, Ree07, Ree09a, Ree09b, Rus08, RL02, Sha06a, SF19, TNH17, Wal15, Wal16a, Wal17a, WC05, Wil17b, ZT06, ZA08, SKSV25]. Intervention [HML07, LFG14, LFG19, PC15, Sie05]. Intra [BKS13]. Intra-class [BKS13]. Intraclass [BF09, BF11, BF13, BF14, Per06]. Intrinsically [EO13]. Invalid [Eni09]. Invariance [FF08b, FF16, KZ08, ZK05]. Invariant [FF08b]. Inverse [ANM15, IS24, JAJ24, KMA18, MA18, OON24, Pra10b, SFA16a, SKSV25]. Inverted [kAEAE21, Pra12, RRS23, MA24]. Investigating [HNG19, LHM14, LKW^{+}19]. Investigation [DW19, El 11]. Investigations [LR03]. Involving [Udo13, VJ11]. IPS [Raj21]. Iran [ZC13]. IRT [AK18b, AK18b, CC19, SH07]. ISE [OAA10]. Ishita [AK21]. Issue [ZZ15]. Italian [CS06]. Item [DW19, FF12, KZ08, LKW^{+}19, MLS10, Nee04, Pra10a, SH09, ZK05]. Item-Fit [DW19]. Item-Level [ZK05]. Items [LB10, LHO^{+}18, NRSGRN02]. ITSACORR [HML07]. IV [SZH^{+}18].

J [Whi07, Wil19b]. Jackknife [AM14]. Jacques [DSCS08]. Japanese [MMS08]. Jeffreys [Lov19b]. JMASM [Whi07, ATvBB17, AK20, ANAS16, AK18a, Alo04, Alo06, ASR^{+}16, ASM^{+}16, AAH^{+}19, BB05, BB06, BD20, CDPT19, CKP12, DP05, Dem05, Fah02a, Fay02, FC09, Hea02, Hea03, Hea06, KH16, KMG06, Lan19, LCW^{+}17, LOCMS13, MS14, MBS17, NS07, OO05, OMN16, ÖK17, Ree04, RP03a, RM16, SAZ17, SH07, SH09, Tak17a, Tay11, Wal03, Wal05b, Wal15, Wal16a, WS16a, Wal17b, ZD10]. Johnson [BPZ12, GR11]. Joinpoint [KKT13]. Joint [AGKJ19, ÇE14, Seo19]. Jointly [Lee13]. Jöreskog [GZ07]. Joseph [SC06]. Journals [Kel05]. Journey [LFG19].

Kalman [EO09]. Kappa [BS09, Heo08, HXH06]. Kendall [Wal03, Wal16a]. Kenward [PA07, PMZ08]. Kerala [VA16]. Kernel [EB04, Eid04, EME10, EAS13, Mug12]. Kernel-Based [EB04]. Key [LC20]. Kim [Sah05]. Kindergarten [Weh03]. KNN [ML24]. know [NT25]. Knowledge [BB07]. Known [Mic11, SGM16, SK13, SP14a, SP14b, TS11]. Kohonen [DEH^{+}06]. Kolkata [RRS07]. KPSS [KE12]. Kruskal [BB04, Hea03]. Kumaraswamy [AC15, AA20, kAEAE21, FEB13, MA24]. Kurtosis [SP14b].

L2E [Dye11]. Lack [LT07]. Lack-of-Fit [LT07]. Lacks [Kra18]. Lag [DY20, RB09]. Lag-One [RB09]. Lagos [AA12]. Lambda [Su05, Su11]. Language [GJ17]. LAO [Nav07]. Laplace [AT09, DGJ16, HZ16, JA18, Ots05]. Large [BAN06, CF03, Fah02b, MC11, Nav07]. Large-Scale [CF03]. Largest [Wil19b, Wil19e]. Lasso [DY20, HWS^{+}24]. Latent [CB12, Eic05, HB08a, Koc16, KCCD10, QWL^{+}16, WWB12]. Latin [JPH23]. Lattice [OO18]. LAV [Die11]. Layouts [SJ25]. LCA [ZD10]. Lead [OL03]. Leading [Saw05]. Leaf [LM17]. Learners [WZS11]. Learning [ML24, NT25, PW24]. Learning-based [ML24]. Least [AKQ10, Die09b, Kha12b, Kha13, NW09, OMN16, Pas09, SAZ17, Tof08]. Leave [KCCD10]. Leave-One-Out [KCCD10]. Left [CGH15, SFA16b]. Legendre [HR16]. Leiman [Bea17]. Leiman-Based [Bea17]. Length [Hur08, MC07, MC08, MK23, RNN12, RNN13a, RNN13b]. Length-Biased [RNN12, RNN13a, RNN13b]. Leone [AGKJ19, OON24]. Less [Leh15]. Let [Kes15]. Letter [Con18, Weh03]. Lev [OS10]. Level [CM02, EK03, JPH23, Kon13, NW09, PM14, Udo20, Wal08, ZK05, DW19]. Level-2 [JPH23]. Levels [FF16, Goo13, Lip12, Man13]. Levene [OYK^{+}12]. Leverage [EN23]. Leveraging [NSJ^{+}24]. liberalization [PW24]. Life [AOCAN19, BK09a, Ebr06, HA14, KR16, LHO^{+}18, LRT20, PP17, Pra10a, Rao11, RK13]. Lifestyle [PH25]. Lifetime [AKA20, AFL19, MSS20b, RAJ18, SSMK22]. Likelihood [AC15, AOAH11, AA21, Bak07, BII11, BLNA^{+}22, CC09, EG11, Eic05, kAEAE21, FF12, FZ09, Gil24, HNG19, KPR14, KS09, Kra18, LSAA02, MMH19, Mey16, Mic11, Ola20, PTC18, PI08, Su11, Wil10, YSVS19, Wil04]. Likert [LB10, MOD07, Nan02, ZGZ07]. Likert-type [LB10]. Lilliefors [AS15a]. Limitations [GL06, RMS17]. Limited [KR16, Lip17a]. Lindley [AOCAN19, GZ17, JAJ24, Lov19b, MZ17, RS20, SS23]. Line [AE12, Eid04, EAS13]. Linear [AS17, ANAS16, AN04, ASR^{+}16, AW16, AO08, BB16, BB19, CAM08, DEE17, Dor16, Dub04, EAN05, Ele17, FBR^{+}08, Gau17, GV03, HSA18, HMNW07, JK14, KPR14, KOW16, Kib06, Kib08, KCA^{+}17, LeB16, LH09, Lip17b, LWZK16, LOA15, Mic11, MAP08, NS07, OMN16, PM14, QS12, RTA22, Raz13, RBAR18, Sah09, SAZ17, SAA12, Udo13, WF09, YS14, YSVS19, RRB08]. Link [Raz13, SWM19]. Linkages [LBC^{+}19]. Liouville [SC06]. Literacy [Weh03]. Liu [Asa16, MKS19]. Liu-Type [Asa16]. Living [Weh03]. Loading [FF16]. Loadings [BS14, FF08a]. Local [ADARAM05, EM14]. Localized [ST14]. Locally [Con02]. Locating [FF08b]. Location [ADAR09, AAK10, Bak03, Bla02, Fay02, Fay06, KK20b, MC07, NW10, Per03, RJR14, Saw02c, Saw05, SY19, WK02b, WK04, Wil16a, WH16]. Location-Scale [AAK10]. Log [BJJ22, BJJ23, Gau17, LAMB17, RRB08, RRB12, Su11]. Log-Linear [Gau17, RRB08]. Logarithm [AAK22]. Logarithmic [GRM20, NE16]. Logic [Boo19, CMR02, Kle19]. Logistic [ADARAM05, ARM10, ASAIJ12, AL15, AZ17, AASIJ10, AGS14, Asa16, BJJ22, BJJ23, BII11, CCH19, Gan12, KRR13, KRR14, KKST05, Lip06, LK12, LK13, Liu14, LAMB17, Pan18, PS02a, PN03, PS17, RRS07, RRB12, SRW21, SIA11, TZZD08, WS16a, WS19, Wil18b, YS16]. Logit [HS08, Min13]. Loglinear [Mos12]. Lognormal [BE06, Ebr06, RD16, RNN12, RNN13b, SA13, VA16]. Lomax [AK20, AF23, MESSE16, NH12, OON24, Rao11]. Long [MDS03, MŞ09, Qu05, SSMK22]. Long-Tailed [MŞ09]. Longitudinal [AKO03, AK04a, AK04b, BZBS03, Dem04a, FBR^{+}08, LX06, LEKG10, LOCMS13, LBC^{+}19, MBS02, OD07, Pér07a, Seo19, Ste16, ZS17]. Look [BPWR10, Lec07]. Looking [NR05]. Loss [GK04, OGKH17, SSM18]. Loss-Based [SSM18]. Lost [BB07]. Lot [SU16b]. Lovric [Saw16, ZK16]. Low [Weh03]. Low-Income [Weh03]. Lower [Ben10, Ben17, KDMAA18]. LOWESS [Wil17a]. Lp [Lip07]. Lp-Metric [Lip07]. LQ [BB11, SJ07].

LQ-Moments [BB11, SJ07]. LQL [KIR12]. LT [AAK22]. Lui [Whi07]. Lyapunov [JR16].

M [MADT03]. M/D/2 [MADT03]. Machine [DK19, ML24, PW24]. Machine-based [DK19]. Macro [LCW^{+}17]. Macroeconomic [SS18, ZC13]. Maddala [Ahm20]. Mahalanobis [Bos09, HSA18]. Main [PDC13]. Maintenance [AT09]. makers [WB06]. Making [Cam09b]. Management [PH25]. Manifestation [ZK05]. Manifolds [RNN13b]. Mann [HH16, Saw05, Wil12, Wil19a, NT25]. MANOVA [FF13, HM10, RRHA11, Tay11]. Manufacturing [Nan10]. Many [LM19, Wal08]. MAPD [KIR12]. Maps [DEH^{+}06, SLC17, SG21]. Marginal [AGKJ19, IT18, TZA08]. Markers [Bos09]. Market [BA11, SS18]. Markov [AK10, BII11, BA11, CCH19, CT09, ICS07, Lip13b, NAK13, OR02, ORSCMG03, SIA11]. Marshall [JAJ24, Rao11]. Mass [MBS02, Par10]. Matched [Gan12, Lee11b, Mal02, Ree09a]. Matched-Pair [Lee11b]. Matching [ÇE14, LMRM^{+}20]. Mathematical [Rod10, ZZ15, SC06]. Mathematics [LOK11, NT25]. Mathmatics [Saw07]. MATLAB [BD20, ÖK17, ATvBB17, Alo06]. Matrices [CFSF10]. Matrix [Bha08, Cam07, DT04, GZ07, HSA18, Wal05b]. Matter [BHH12]. Matters [Wal04]. Maximal [MA17]. Maximize [Opd03]. Maximum [AC15, AOAH11, CC09, Eic05, kAEAE21, FF12, Gil24, HNG19, KI10, KS09, MS14, MBS17, MMH19, Mic11, Neu09, NH06, PTC18, PI08, RC17, Su11, DSK18]. Maxwell [IS24, SK07]. May [Sha06a]. Mckean [NS07]. MCMC [MESSE16]. McNemar [Day17, DT04]. Mean [ADAR09, ADAAM09, ASS19, AO15, AOBCP16, AKK24, BM05, Bou10, BS02, Cam07, Cam09a, CBP18, CGH15, CERA24, CEI24, Dan25, GSP24, GCP24, Kad16, KSS10, KN05, KS06, KV14, KC20, MBS16, MC08, MK10, OS10, PS18, PPC18, RL16, SLLL11, ST02, Sha08a, SS16, SF19, SKB11, STJ11, SVS16, SSS16, SD17, SK18, SY19, SSS21, SB09, Sub13, SP14b, SRA17, TS11, TL14, VZ20, WWB12, Wil10, YTP24, ZA08, ZA11, ZK16, DSK18, IB24]. Mean-Variance [SRA17]. Means [BC02, CCC19, DRTW17, GJ17, KKW^{+}06, MT24, NKdG^{+}16, PS02b, RRB12, RRS23, RL02, Saw02b, TC12, Wal05b, WWB12, WK02a]. Measure [Aly07, ÇE14, Heo08, IT18, MJ20, Saw02c, SPB19, TW07, Wal05a, Wil16a, WH16, Wil18d]. Measurement [Fan10, Han03, KS10, KBKBG07, Kel05, KZ08, LHH11, SS12, SS13, SPT12, SSAH18, SK10, SS15, WFC15, YTP24, ZK05]. Measurements [ATvBB17, KS10, KZ18, LL07, VSS16]. Measures [Day03, FBR^{+}08, KJI10, Kin03, KMA21, LHB04, LH04, LZ07, MBWZ07, NW10, OL03, ON14, PJ07, PS22, PW04, RSQ20, SLLL11, SL17, SB09, SLA16, WK02b, WZ17]. Measuring [Fer10, SLC17, TZZD08, XGY^{+}08]. Meat [PK06]. Mechanical [Saw02a]. Median [AS10, AOY11, Bic11, MC08, OAC03, RD16, ST02, SQ15, SK13, SP14a, Tho10, Wil18d]. Median-Unbiased [Tho10]. Medians [RM07, Wil05]. Mediator [WFC15]. Medical [Gan14]. Medication [Sha11].

Medicine [KHC05]. Medium [Mar12]. Melanoma [ST14]. Member [TL15]. Membership [CKPJ18]. Memory [Qu05]. Mentoring [Hal04]. Meta [FB10, ML03, Saw03c, Tho16, TSK04, Wal03, WBN07, XGY^{+}08]. Meta-Analyses [XGY^{+}08]. Meta-Analysis [ML03, Saw03c, Tho16, TSK04, WBN07]. Meta-Analytic [FB10, Wal03]. Metabolomics [LWZK16]. Meter [KS10]. Method [AAK22, ANAS16, Bak07, CS04, DEE17, Die09a, EO13, kAEAE21, FZ02, FR04, GVCM18, Hea02, Hea06, JADB10, KJMA18, KS25a, KH16, KKZB16, LM17, LS13, LA14, LOA15, MC11, NRSGRN02, PM14, Per06, RWS09, SK09, SA07, ST14, SW08, SS15, Smi09, WWB12, Wil18a, Wil18b, Wil19a, NT25]. Methodology [ASM^{+}16, MS03, Pér07b, PE08, SLC17, Sen02, Tur09]. Methods [AM14, AKQ10, AK04b, AZN20, ACP05, AŞ15b, Ber05, Che14, CC19, CGH15, DW19, Fay06, Fin11, FF12, Fin14, FR04, HSA18, HL16, IJ06, KS14, KSM13, KM14, KCCD10, Lan13, LB16, LHO^{+}18, LMS15, MYY^{+}06, NW09, OR02, PZ11, PNG08, PNZM24, QS12, Rao17, Rus08, SKB15, SAZ17, Tho09, Wan03, WA03, YHPR17]. Metric [Lip07]. MEWMA [SV12]. MicceriRD [Lan19]. Michelson [Con18]. Microarray [DGJ16]. Midpoint [LM17]. Migration [TL15]. MIMIC [GZ07]. Minimized [Dye11]. Minimizing [Die06, Lip07]. minimum [DSK18]. Mining [AGHL08, Raz13, Xu05]. Minitab [PS02a]. Misapplying [Ber05]. Misclassification [LH13]. Misclassified [Lee11a, Lee11b, Lee13, Lee15, SBY07, TP16]. Misconceptions [Saw05]. Misguided [Fra19]. Misleading [Sha06a]. Missing [AKO03, AK04a, AK04b, Che14, CTDVD05, Kna19a, LB10, LHO^{+}18, LOCMS13, OO18, PA04, PA07, PMZ08, PC18, PC14, SS24, SVS16, ZS17, ZCB^{+}17, Koc18]. Missingness [QWL^{+}16]. Misspecification [HL16, LeB16, OYO16, SLLL11]. Misspecified [Bea11, Fan10, YOO14]. Misuse [RL02]. Mixed [ANA12, DBH02, Eic05, GV03, HMNW07, HL16, KZ08, KIR12, LeB16, LLHT03, Sah05, SLA16, TGhC^{+}05]. Mixed-Effects [HMNW07, HL16, LLHT03]. Mixture [AK18b, CC19, FA13, FF12, FZ09, LA18, LHM14, LHH11, PD11, SFA14, SFA16a, WZS11, ZD10]. Mixtures [HNWD05, Su11]. ML [GZ07, KS25b]. MLE [Hur09, PI08]. MLEs [Hur03]. MNAR [Kna19a]. Model [AE12, AAI15, ANAS16, ANM15, AK18a, ASR^{+}16, AAH^{+}19, AK18b, AW16, AO08, BS04, BJJ22, BJJ23, BII11, Boo19, CE15, DEE17, Day03, DY20, Ebr06, EAM09, EM14, ESAH20, Eni09, EN23, FR03, Fan10, Fin19, FF16, Fuk05, GD02, GVCM18, GG09, GS06, HNWD05, HL16, HXH06, HK10, HY03, HS08, Im20, JK14, Jay17, KKT13, KC06, KOW16, Kle19, KCA^{+}17, KK20a, Lan09, LeB16, Lee12, LH09, Liu09, LOK11, LRT20, Lor19, LOA15, MMH19, MF11, Man05, MSS20b, Min13, MRKM^{+}17, MAP08, NS07, NAK13, OL11, OS10, PD11, Pan18, Pap16, PK06, PP17, PS22, Pra14, QWL^{+}16, RRS07, RRB08, RAJ18, RMS17, Saa19a, SKO09, SSMS10, SS14, SL16, SSS21, Tig03, Tor18, UAD16, Udo20, WW10, WF09, Wil06b, Wil08a, Wil04, Wu04, WZS11, XZ05]. Model [YT10, YOO14, bZbWA13, PW24, OGKH17, PK06]. Model-Based [SL16]. Model-Robust [EM14]. Model-Selection-Based [Fuk05]. Modeling [ANAS16, ACP05, AE11, ANA12, BD20, BOO10, BBB12, BPZ12, CT05, CE15, CCH19, CID^{+}14, CT09, CCC19, Dem04a, FB10, Har10, LS16a, LA18, Lip13b, LM15, Lip19, LHH11, LZW14, Min13, OD07, Oya11, PS02a, Pos02b, PM08, PC14, RS17, Res08, RL10, Rod10, SA02, SA14, Tig03, TL15, Wil04, YD17, ZC13, bZbWA13, Koc18]. Modelling [RM16]. Models [AK20, AFL19, ANA12, BB16, BZBS03, Bea11, BA11, Cam07, Cam09a, CC19, CKPJ18, CB12, CID^{+}14, CAM08, Die09b, Dub04, Eic05, FBR^{+}08, FF12, Gau17, GV03, GCH08, GG09, GS19, HMNW07, HB08a, HL16, ICS07, IJ08, KS14, KJ14a, Kib08, LLHT03, LHM14, LCW^{+}17, LK12, LK13, Liu14, LB19, MSS07, MEH08, MW07, MMS08, Mos12, MLS10, NSSO14, OL11, OMN16, OYO16, OCB13, OAA10, PH25, PM14, PN03, PM03, QS12, RK13, Raz13, RBAR18, SW06, Saf08, Sah09, Seo19, SH07, SH09, SE12, SIA11, SA11, SAA12, Su16a, SLA16, Tor18, UAD16, Wal16b, WS16a, WWB12, WF09, YS16, YD17, Zha05, ZD10, YS14]. Modern [AZN20, ZS17]. Modification [BS04, DT04, Fan10]. Modifications [OYK^{+}12, Pet02]. Modified [AS15a, AOAH11, ASAIJ12, CEI24, DK19, GSP24, Kho07, KC20, LL16, MKA23b, SHA06b, STJ11, Sub13, SP14a, SP14b, ZL04]. Modulated [FRJ02]. Modulo [DS25]. Moment [ASK16, AGKJ19, AF23, HW11, KDMAA18, Wal17b]. Momentary [BKS13]. Moments [BB11, CS12, HB12, MAP08, SJ07]. Monitoring [Fuk05, MW07, SLA16]. Monte [AŞ15b, BA11, CT05, CC09, CCH19, Dye16, El 11, Ele17, FBR^{+}08, FC04, FF12, FF13, Har18, Har19, HS11, Kan05, LAA17, Mic13, MSK^{+}11, MŞ09, NKW^{+}19, Opd03, RM16, Tho09, YOO14, YS14, YS16]. Months [Sha11]. Mood [LS17]. Morley [Con18]. Mortality [HPSR23, KKT13, LEKG10, MMS08, PM08]. Most [Con02, Wil18a]. Mounting [Sen02]. Moving [ADAR09, AHAO09, KQC06, KS06, MF11, Rah14, RR24, ST07, ST08, SA11]. Mplus [LHM14]. MR [EC08]. Multi [KJMA18, KJ14a, KZ08, SH09, ZK05, SJ25]. Multi-Group [KZ08, ZK05]. Multi-objective [KJMA18, SJ25]. Multi-unidimensional [SH09]. Multicollinearity [BB16, BB19, Dor16, DS16, HWS^{+}24, Kha12a, KI16]. Multicomponent [BJJ22, BJJ23, Rao14]. Multicore [Fog15]. MultiFactor [Goo09]. Multilayer [KK16]. Multilevel [CD09, CKPJ18, JPH23, LZW14, LOCMS13, Lor19, RS17, SSAH18]. Multinomial [CS12, MW07, PN03, PM08, RRS07]. Multiple [Abe15, ANAS16, AN04, AKP08, ACB06, ASR^{+}16, BB16, BB19, BBW11, BH06, BJ11, CTDVD05, CKPJ18, Dan25, DMBR16, FF16, FB10, KKW^{+}06, LHB04, LB10, LS16b, LB16, LMRM^{+}20, Lip09, LHO^{+}18, LOCMS13, MS03, MLS10, Nav07, Opd03, OE12, PNG08, Rah17, RM07, SAZ17, Tak17a, Tak17b, Tur09, WFW02, WK02b, Wil07a, ZCB^{+}17, ZL04]. Multiple-Comparison [LS16b, ZL04]. Multiple-Sample [Rah17]. Multiplicative [ANM15, Zha05]. Multiplicative-intercept [Zha05].

Multiplicity [Fra19]. Multistrata [CS04]. Multivariate [AA12, BS04, Bha10, Bos09, BLNA^{+}22, CE15, CERA24, CS04, Dem04b, FBR^{+}08, FRJ02, FR04, Gao04, Hea04, KN05, Kho05, Kib08, KH16, LH04, LL07, Mak17, MIN18, MBWZ07, NXFN19, NW10, NM17, NM19, PDC13, RS17, Rao17, SL17, SV08, SLD14, VSS16, WK04, WH16, WBN07, ZC13, Zha05, Lip13a]. Multiwavelet [HR16]. Mutation [Jay17].

n [KIR12, Neu10, DiS02]. Nadarajah [AEOD24]. naïve [FKP^{+}25]. Nakagami [KMA21]. Nanosensor [NSJ^{+}24]. NASDAQ [RR24]. National [OAA10]. National-100 [OAA10]. NDHS [HPSR23]. Nearest [DNUH22]. Need [BBW11, Ber17]. Needed [Wal08, WB06]. Neether [Kes15]. Negative [FA11, HA07, HB08b, HB12, Mir08]. Neighbor [JVA10]. Neighbors [DNUH22]. Nested [SM17, SLA16, Tor18]. Net [AL15, HWS^{+}24]. Network [KK16, PNZM24, RTS16, SSM18, UAD16, Udo20]. Networks [BD20, LSAA02, Lan09, MYY^{+}06, SJ25]. Neuhäuser [NKC18]. Neural [BD20, LSAA02, Lan09, MYY^{+}06, PNZM24, RTS16, SSM18, UAD16, Udo20]. Next [Sha11]. Neyman [DR18]. NHL [Hur08]. NHST [BD16, Rod10]. Nigeria [AA12, HPSR23, SA14]. Nine [WS16a]. No [OL03, Rod10]. Noise [Bha08]. Nominal [BS09]. Non [AAI15, AA12, Alk13, BS11a, CS09, DP21, DWT19, DK19, EO13, Gan14, GK20, GT18, Hea04, Hol06, JR16, KRP02, KSS10, KY06, KY08, KOW16, Koc16, KH16, KSB10, KV14, LH13, LZ07, LS05, MSS19, MIN18, NKdG^{+}16, NKW^{+}19, PMZ08, PS18, RTA22, Saw13, SS16, SKB11, SS15, SD17, SY19, SJAZ14, Tor18, TX11, Wal05a, WPN^{+}18, Wil15b, ZA08, ZJ08, Con18]. Non-Central [ZA08]. Non-Communicable [AA12]. Non-Conventional [SY19]. Non-differential [LH13]. Non-Existent [KSS10]. Non-Gaussian [SJAZ14]. Non-homogeneous [TX11]. Non-identically [JR16]. Non-Independence [CS09]. Non-Independent [KRP02]. Non-Inferiority [GT18]. Non-Linear [RTA22]. Non-Nested [Tor18]. Non-Normal [Alk13, BS11a, DK19, Gan14, KOW16, LS05, NKdG^{+}16, PMZ08, Hea04, Hol06, KH16]. Non-Normality [KY06, Koc16, MSS19, NKW^{+}19, WPN^{+}18, Wil15b]. Non-Normally [SD17]. non-null [Con18]. Non-Overlap [Wal05a]. Non-Parametric [DWT19, EO13, GK20, LZ07, MIN18, ZJ08]. Non-Proportional [AAI15]. Non-Regular [KY08]. Non-Respondents [KV14, SKB11]. Non-Response [KSB10, PS18, SS16, SS15]. Non-Statistically [Saw13]. Noncentral [Lec07]. Nonlinear [BOO10, HL16, Lip10, MB05, Sah09, YS14]. Nonnormal [CWBK07, LH14, WS16b]. Nonnormality [CCC19, KWOF02]. Nonparametric [ACB06, ACP05, AK18b, BB04, BB05, Bla02, BE17, CGH15, CS04, Fah02b, Fay02, GMS05, HL03, KBKBG07, LLHT03, Opd05, VSS16, Wil18d, Wil19c, CS04]. Nonresponse [Dan25, MLS10]. Nonrigorous [Kho05]. Nonsignificant [LR03]. Nonstationary [El 11]. Normal [AHAO09, ASS04, Alk13, AK13, BB06, BS11a, CFSF10, DP21, DK19, Dye16, FLE04, FZ09, Gan14, Hur03, KOW16, KK13, KMG06, LS05, NKdG^{+}16, NM17, OY17, ÖK17, Oya11, PMZ08, RL16, RC17, SP16, TGhC^{+}05, ZK16, Hea04, Hol06, KH16]. Normal-Based [Dye16]. Normality [CE15, Eni09, KY06, KOW13, Koc16, LAM16, MSS19, NKW^{+}19, Ots05, RR24, WPN^{+}18, Wil15b]. Normalizing [SS09]. Normally [AO08, SD17]. Normals [HNWD05]. North [BB11]. North-Bank [BB11]. Not-So-Quiet [RL10]. Note [Die09a, Hur03, Hur07, RNN13b, Rav17, Wil19d, ZL04]. Notes [Lui10]. Notion [Lov19b]. Novel [LS16b, NXFN19]. Nuisance [ADARAM05]. Null [BH16, LSAA02, RL16, Saw16, ZK16, Con18]. Number [ANA12, BK09b, Dem04b, Dem05, Fog15, HS11, JPH23, Kan05, Kan06, LIT24, Mar03, MSK^{+}11, RB11, SHA06b]. Numbers [DS25, TZA08, Tho16]. Numerical [Hea06]. Nyther [Kes15].

O [Saw02a]. Oaxaca [LF11]. Objective [RMA14, KJMA18, SJ25]. Objectives [Saa19a]. O’Brien [OYK^{+}12]. Observational [AZN20]. Observations [AK04a, DRTW17, DWT19, JR16, KRP02, Lip19, Mug12, SVS16, Wal08]. Observed [LMRM^{+}20]. Obtain [Wal08]. Obtained [DDM07, MMM07]. Obtaining [OAC03]. Occasion [Saw04a, SS16, SSS16, SK18]. Occurrence [Raz02]. Occurrences [SIA11]. Odds [Liu09, LK13, OL11]. Odor [HMNW07]. Off [Kes15]. Oil [RNN12]. OK [Rah14]. Old [PM08]. Oldest [PM08]. Oligonucleotide [HY03]. Olkin [JAJ24, Rao11, Wal17b]. OLS [Min13, SW06]. Omega [EYS^{+}07, PD13]. Omnibus [Wil06a, Wil07a]. On-Site [MEH08]. One [BB04, BH16, BBW11, BS11b, GT18, Hea03, KCCD10, LCW^{+}17, MW07, NS07, NKC18, NKW^{+}19, PA04, RB09, WC09, ZA08, ZA11]. One-Factor [LCW^{+}17, NKW^{+}19]. One-High-Threshold [MW07]. One-Sided [BS11b]. One-Way [BB04, Hea03, NS07, ZA08, ZA11]. Online [BA11]. Only [Leh15]. Openness [Fer10, Fer12]. Operating [GZ07]. Opposition [Fra19]. Optimal [ATvBB17, Asa16, CEI24, KS10, Lip07, MF11, Pap16, PW24, RR07, RJR14, SP16, Tho10, TTBG06, Tor18]. Optimization [KJMA18, KI10, PH25, SJ25, SRA17]. Optimized [WC05]. Optimizing [KS25b]. Optimum [DR18, DDM07, VS08, XZ05]. Oral [AAH^{+}19]. Orcutt [Die09a, SAA12]. Order [AA21, AK21, BII11, BH06, CCT09, CS12, HB08a, Hea06, KKK19, KC20, Mic13, Ree11b, SIA11, TuAH19]. Ordered [BB16, BB19, BI02, Min13, Neu09, Pos02a, Wal16b]. Ordering [Wil19e]. Ordinal [BII11, CTDVD05, FC04, Heo08, Lee11b, Liu09, LOK11, LK12, LK13, Liu14, LB19, OD07, Rus08, SWM19, TNH17, ZGZ07]. Ordinary [Kha13]. Organizations [BBW11]. Organizing [DEH^{+}06]. Orthogonal [YT10]. Other [HH16, ST02, ZL04]. Outcome [DMBR16, Ree06]. Outcomes [Har18, ICS07, MA06, MRKM^{+}17, Seo19, NT25]. Outlier [LB16, LLB17, RR07, SL16]. Outliers [Dor16, EAN05, Ele17, MRMH17, SK11]. Outlyingness [Mak17]. Ovarian [Bos09]. Over-Dispersed [ANA12]. Overall [El 11, XGY^{+}08].

Overdispersed [Doy09]. Overdispersion [MEH08, bZbWA13]. Overlap [Wal05a].

p [Neu10]. Package [BO13]. Padding [NSSO14]. Pair [ASG16b, Lee11b, Lip09, PD05, PDC13, ZC13]. Pair-Copula [PDC13, ZC13]. Pair-Wise [PD05, ASG16b]. Paired [DRTW17, DWT19, DBH02, EB04, FKL^{+}03, RYZ16]. Pairs [Gan12, Mal02]. Pairwise [BM05, GDG20, Opd03, ON14, RRHA11]. Panel [KCA^{+}17, Raj21, SH15]. Paoli [CS06]. paradigm [SJ25]. Paradox [Lov19b]. Paradoxical [HXH06]. Parallel [Wat06, GJ17]. parallelism [Cha22]. Parameter [AJ17, AOCAN19, ACB06, Asa16, AŞ15b, Bak03, BAN06, BK22, Fan10, FEB13, FF12, Geo14, HM08, JA18, JPH23, Kha12b, KI16, LS13, LAA17, MN05, MN17, NE16, NAK13, OHQ^{+}20, PS18, PP17, Qu05, SMSN20, Saw05, SGM16, SP14b, Wal06, SKSV25]. Parameter-Free [LS13]. Parameterization [Lip10]. Parameters [ADARAM05, ADAR09, AOAH11, AM19, BE06, Cha24, Die06, Ebr06, EAM09, Eni09, GR11, GK20, Hur03, KMA18, KM14, KK20b, KMA21, Mic11, OGKH17, SQ15, SFA14, SY19, SBY07, SA13, TS11, TSS20, TP16, YAS14]. Parametric [AAI15, AA20, BLEL02, CGH15, DEE17, DWT19, EO13, EAS13, GK20, KBKBG07, LZ07, MIN18, NKdG^{+}16, PJ07, SKO09, SK09, SPB19, Su16a, VO03, ZJ08]. Pareto [AW05, AFL12, HHN20, PTC18, Pra14, Pra20, VA16]. Parity [Opd05]. Parsimony [Wal04]. Part [FZ02, FZ09]. Partial [BG15, BH06, EG11, FF16, Heo02, Hur03, Kna15, OL11, PS22]. Partially [LMRM^{+}20, LRT20]. Participatory [Kle19, Saa19a]. Partition [Su11]. Passengers [LIT24]. Patel [Wil19a]. Paternity [CMR02]. Paths [Tur09]. Patient [TSK04]. Patients [MMM07, ST14]. Patterns [GCH08, Odi12]. PBIB [GDG20, SG21]. PCI [PE08]. PCIC SAS [DP05]. Peak [KS25a]. Pearson [HW11, Wal17b]. Peer [And13b]. Penalized [AL15, LX06, LWZK16, Ola20]. Penalty [Mey16]. Per-Family [Fra15]. Percent [Lip17b, Wal05a]. Percentage [Tof08]. Percentile [KH16, ZA08]. Percentile-Based [KH16]. Percentiles [BB06, DEE17]. Perceptual [SLC17]. Perfect [Bea17]. Perform [RP03a]. Performance [EM14, Eni09, Gil24, HWS^{+}24, Hus07, Im20, KN05, LB10, LKW^{+}19, TNH17, ZS17, ZA08, FKP^{+}25, PW24]. Performances [KB16]. Performing [RP03b]. Period [Ree12, Sha11]. Periodic [BC05, WCW08]. Periodically [SAA12]. Periods [GS19]. Permutation [And13a, And13b, CDPT19, CM02, Gao04, HH16, MPS11, NH06, OO05, Odi12, Opd03, PB13, RM07, Web07, WS09]. Permutation-Based [HH16]. Permutations [DiS02]. Persian [AS16]. Personal [Hal04]. Perspective [Hal04, JPH23, RL16, Rav17, WFC15, ZK16]. Perspectives [Les11]. PERT [MEP09]. Phase [Dye11, KV14, Lui10, MBS16, SZH^{+}18, VZ20]. Phases [WZS11]. Phasor [KS10]. Phenomenon [Par10]. Physical [WBN07]. PI [ZD10]. PI-LCA [ZD10]. Pickands [Fer17]. Piecewise [Res08, WZS11].

Pietro [CS06]. PINES [OCB13]. Pinus [OCB13]. Placement [KS10]. Plan [HBA20, KR16, KIR12, MSS19, SSK13, SP16, SU16b]. Planning [WBSMI09]. Plans [AOCAN19, Rao11, SS12]. Plant [LM17]. Plot [BZ09, Kan05, PA07, PMZ08]. Ploya [HA07]. PLS [Koc16, Koc18]. PLS-based [Koc18]. PLS-SEM [Koc16]. Plus [Wil18d, Wil19e, YT10, AAK10, AKA11]. PMUs [KS10]. Point [Ela06, GV03, IT18, Pra14, QT06, RL16, Saw16, ZK16, ZD10]. Point-Referenced [GV03]. Point-Symmetry [IT18]. Points [Wil06a]. Poisson [AA12, ANA12, CS08, FR03, GZ17, KM14, KJ14b, Lee12, MMS08, PS22, RS20, RNN13a, SS23, SF19, SA12, TX11, XYJ03, bZbWA13]. Pollution [KKZB16]. Polychoric [CC09]. Polynomial [Hea04]. Polynomials [Hea06]. Polytomous [Eic05, Gan12]. Pooled [Bic11]. Pooling [GMS05]. Popular [HBA20, ÖK17]. Population [ADAR09, ADAAM09, AS10, ASS19, AOBCP16, BKS12, CERA24, CEI24, Dan25, Ela06, GSP24, GCP24, GS19, Heo02, Hit09, IB24, Kad16, KKZB16, Kum13, KV14, KC20, LBC^{+}19, MBS16, MPC^{+}16, MT24, Mic13, NKdG^{+}16, PH25, PPC18, RK13, SS13, SS16, SS24, SKB11, STJ11, SGM16, SVS16, SSS16, SK18, SY19, SSS21, Sub13, SP14b, TS11, TC12, TL14, TS16, VZ20, Wil10, ZRB23, DSK18]. Population-Based [LBC^{+}19]. Populations [ACB06, BF14, BG15, CFSF10, Gao04, GG09, Kho07, NKC18, RN18, Kna19b]. Porridge [BB19]. Portfolio [KI10, SRA17]. possible [Con18]. Possibly [SBY07]. Post [HM10, MT24]. Post-Stratification [MT24]. Posterior [SA12, SA13]. Posteriori [CC09]. Potential [HSA18, RTS16, Saw02a]. Poverty [HP03, Wil04]. Power [Abe15, ADARAM05, BW05, BE17, CID^{+}14, CM02, CS09, Doy09, Ebr06, FBR^{+}08, Fay06, FKL^{+}03, GP04, Hea06, IS24, JAJ24, KBKBG07, KWAO04, KH16, Kra18, LLB17, LW05, LS17, MBWZ07, OON24, Opd03, OE12, PM14, RRHA11, RBAR18, SHA06b, SM04, Smi09, Tay11, TGhC^{+}05, VM09, Wal15, Web07, WS09, WK02a, Wil08a, ZZ15]. Powered [Lip07]. Powerful [Con02, Kon13, Opd05]. PPS [VA16]. Practical [BE17, Fuk07, LMRM^{+}20, LOCMS13, Wil19f]. Practice [KOW13, LS06]. Practices [LOCMS13]. Pratt [Wal17b, WZ17]. Pre [NRSGRN02, TTBG06]. Pre-assigned [TTBG06]. Pre-Testing [NRSGRN02]. Precision [LH14, Udo20]. Preclinical [WCW08]. Predication [kAEAE21]. Predict [PN03]. Predicting [LOK11, ST14]. Prediction [AS17, Alo04, BA11, CdlTE22, CB12, Fin14, IS24, Lip17b, Lip19, RJR14]. Predictions [FA13, LMME09]. Predictive [Kib08, Wal04]. Predictor [AZ17, WFW02]. Predictors [BS14, Bea17, Lip06, LZW14, Wil07a, Wil11, Wil18b]. Preference [EG11, GMS05]. Pregnancy [PC14]. Preliminary [AW16, KOW13, LAM16, PS02b, Per03]. Preparing [WB06]. Preprocessing [NSSO14, SL16]. Presence [ADARAM05, BB19, EAN05, GRM20, Kha12a, KSB10, SW06, Saf11, SH15, SS13, SPT12, SS15]. President [Jin04]. Pressure [SLA16]. Pretest [NR05]. Prevalence [Kit10].

Prevention [QWL^{+}16]. Price [XZ05]. Prices [DR14]. Primer [Rao17, ZLB19, Kna19b]. Principal [AW16, Kan05, NLQ10, PNZM24]. Prior [BY11, OY17, PP17]. Priorities [Lip13b]. Priors [KJ12, RMA14]. Probabilistic [HW11, Sha11]. Probabilities [AZ17, HM08, KHC05, Lip13b, SIA11]. Probability [AT09, HNG19, Hur07, LA18, LM15, MEP09, MC07, MC08, MAP08, PS22, Raz02, Wil17b, Wil17a, Wil19c, Wil19b, Wil19d, YOO14]. Probable [Saw02b]. Probably [BPZ12]. Probit [Raz13]. Problem [Dan25, DS16, Fra19, HL03, Jin04, KI10, MMM07]. Problems [KJMA18, WE05]. Procedure [AOY11, BC05, BC19, Bou10, Dub04, HM10, HS18, LS16b, MBS16, MBWZ07, PI08, SSS16, SAA12, Tay11, ZL04]. Procedures [BM05, BJ11, Die09b, DBH02, Ela08, KKW^{+}06, LL07, LAMB17, OE12, SS24, TS16]. Process [KN05, KS06, MN05, SV08, ST07, ST08, SJAZ14, TX11]. Processes [Alk13, DAC02, Nan10, OR02, ORSCMG03, Saf11, TZA08, WSN15]. Processing [Bha08]. Processors [Fog15]. Product [HW11, KC20, MT24, PS18, SK07, SK10, STJ11, VZ20, Wal17b]. Product-Moment [Wal17b]. Production [Saw03c]. Proficiency [LOK11]. Proficient [TS18]. Profile [GMS05, VSS16]. Prognostic [AAH^{+}19]. Program [ATvBB17, Boo19, DiS02, Fra17, Les11, Wal15, Wal17a, ZD10]. Programming [KI10, OMN16]. Programs [Abe15, CBP18, Kle19]. Progress [AK10]. Progressive [BP16, FEB13, GRM20, PPS14, PS17]. Progressively [KJ06, KJK09, LRT20, Pra15]. Projective [DDM07]. Promotion [WW10]. Propagation [Koc16]. Propensity [Bar15, LMRM^{+}20, LKW^{+}19]. Properties [ANO^{+}25, AAA15, AALF14, AF23, Dye16, ESAH20, EK03, FLE04, HB08b, HNG19, KC06, KLS07, LS16a, LFO07, MA24, MA06, NAY^{+}18, RAJ18, SMC16, SPB19, Udo13, WC05]. Property [ZT12]. Proportion [FB04, GT18, TS16]. Proportional [AAI15, Liu09, LK13, OL11, PS22]. Proportionality [GCP24]. Proportions [CS12, Ree04, Ree09a, Ree09b, Tur17, VM09, Wal05b, WC05]. Proposed [Kha12b, NW10, SF19]. Proprietary [Wan03]. Protected [BM05]. Protocol [CPT04]. ’Proven [GJ17]. Provinces [HP03]. Pseudo [AOY11, Dem04b, Dem05, Fog15, HS11, WS16a]. Pseudo-Median [AOY11]. Pseudo-Random [Dem04b, Dem05, Fog15, HS11]. psychological [WBN07]. Public [OGKH17]. Publication [ML03].

Q [Kho04, NRSGRN02]. Q-Sort [NRSGRN02]. QSS [KIR12]. QSS-1 [KIR12]. Quadratic [RC17]. Qualitative [RR03]. Quality [Cha24, Cox13, Rav17]. Quantifying [FZ02, FZ09, FB04]. Quantile [WC09, Wil12, Wil16b, Wil19f, ZJ08]. Quantiles [KS25a, RK13, Wil08b, WC09, Wil16a]. Quantitative [Bou10, Cli11, HS18, ÖÇ17, SS14]. Quantum [Par10, Saw02a]. Quartiles [SP14a]. Quasi [Eic05, GZ17, HB08b, HB12, HXH06, Wil04, YT10, YSVS19]. Quasi-Independent [HXH06]. Quasi-Likelihood [YSVS19, Wil04].

Quasi-Lindley [GZ17]. Quasi-Maximum [Eic05]. Quasi-Negative-Binomial [HB08b, HB12]. Quel [Sah09]. Questionnaire [AS16, NRSGRN02]. Queue [MADT03]. Quiet [RL10].

R [HBA20, LMS15, Lui07, OMN16, AK20, AAK10, AKA11, BO13, CDPT19, Dem04b, Dem05, Hit09, MBS17, PM14, RM16, Su16a, SA12, Tak17a]. R. [Saw04a]. R2 [BB19]. RACOG [FKP^{+}25]. RACOG-RUS [FKP^{+}25]. Radium [Lee12]. Radj2 [Wal15]. Rainfall [SIA11]. Random [BZBS03, CPDP18, Dem04b, Dem05, FEB13, Fog15, HS11, Hur03, Kad16, KK13, Kra18, KC20, KMG06, LM15, LMME09, Mar03, RNN13a, Saw04b, SK07, SS24, SK10, SS14, TC12, TL14, TD03, YTP24, YD17, DSK18]. Random-Slope [YD17]. Randomization [FF08a, LFG14, LFG19, Man13, SW08]. Randomized [AKO03, AK04a, BLEL02, Bou10, CS12, GS06, HS08, Kon13, Lui06, MK05, OS10, ÖÇ17, PS22, Sah05, SS14, TS16, TS18, Whi07]. Randomly [CGH15]. Randomness [WB02]. Range [CCT09, KQC06, Rah14]. RANGEN [Fah02a]. Rank [BZ09, Con02, DBH02, Dub04, KM02, LHB04, Mak17, MOD07, Mic13, MCL09, NS07, Nan02, Pet02, Rus08, SS09, TW07]. Rank-Based [DBH02, KM02, MCL09, SS09, Dub04]. Rank-Order [Mic13]. Ranked [ADSBH02, ADAR09, ADAAM09, AHAO09, AOAH11, ASS04, BC19, BK22, Bou10, CBP18, SMSN20, ST02, SS04, SQ15, TuAH19, YAS14]. Ranking [SKB15]. Rankings [EG11]. Rao [BRT17, Saw16, ZK16]. Rare [PS22]. Rasch [Pan18]. Rate [And13a, And13b, DMBR16, KPR14, OE12, PB13, PKRK25, Sha11]. Rated [Min13]. Rater [BS09]. Rates [AOY11, FKL^{+}03, Fra15, Har19, HP03, KM02, LB16, LEKG10, LHO^{+}18, PA04, PA07, PMZ08, RBAR18, SM17, SW08, TMB15, WA03, WA07]. Rating [ZGZ07]. Ratings [Eni09, GMS05, Leh15]. Ratio [CFSF10, CERA24, CEI24, FZ09, GSP24, GCP24, Heo02, KPR14, KK13, Kra18, LSAA02, LOK11, LB19, MKA23b, MT24, PS18, ST02, SS04, SK10, STJ11, Sub13, SP14a, SP14b, TC12, TL14, Tak17a, Tak17b, VZ20, YSVS19]. Ratio-cum-Product [VZ20]. Ratio-Product-Ratio [PS18]. Ratio-Type [GSP24, GCP24, TL14]. Ration [Pér07a]. Rayleigh [AJ17, AF23, BJJ22, BJJ23, EAM09, MA24, MK23, OHQ^{+}20, PPS14, Pra10b, Pra15, Rao14, SFA14]. Rayleigh-Exponential [BJJ22, BJJ23]. Re [QS12]. Re-sampling [QS12]. reaching [PH25]. Reaction [PNG08, Saw02a]. Real [BK09a, Nan02]. Really [BHH12]. Reasons [Wil19f]. Reassignment [BR03]. Recall [Pre04]. Recapture [GG09, GS19]. Recently [NW10]. Recognition [PNZM24, Weh03, WZS11]. Recommendation [ML24]. Recommendations [LMRM^{+}20]. Recommended [PM10]. Reconciliation [KS10]. Record [AGKJ19, AA20, AF23, HNG19, KDMAA18, KD18, MESSE16, NH12]. Records [TSS20]. Recovery [DAC02]. Recreation [MEH08]. Rectangular [OO18]. Recurrence [AGKJ19, Sha11]. Recursive [MDS03]. Reduce [Pan18, YII13]. Reducing [LEKG10, SY19]. Referenced [GV03]. Referents [FF08b]. Reflections [BD16, Hig04, Kna02]. Regarding [Con18, Whi07]. Region [BB11]. Regional [BB11, DNUH22]. Regression [Abe15, ARM10, AKA11, ANAS16, AKQ10, AL15, AZ17, ASR^{+}16, AW16, ANA12, BB16, BB19, BY11, BZBS03, CdlTE22, CB12, Die09b, Die11, Dor16, EAN05, EM14, FR03, GD02, Gan12, Gan14, HSA18, HM08, HWS^{+}24, HY03, IJ08, JK14, KKT13, KK16, KKST05, Kha13, KI14, Kib06, KB16, Lee12, Les11, LC05, Lip06, Lip07, Lip09, Lip12, Lip17b, Lip19, Liu09, LOK11, LK12, LK13, Liu14, LWZK16, LOA15, MMH19, MKA23a, Min13, MB05, MAP08, MŞ09, NW09, NLQ10, NSSO14, OMN16, Ola20, OA11, Pan18, PS02a, PM03, PM08, RRS07, SW06, Saf08, SA10, SAZ17, SRW21, SW08, SIA11, SAA12, SWM19, Su16a, TZZD08, Tho09, Tig03, Tof08, Ura20, Wal16b, WS16a, WS19, WFW02, WE05, Wil07a, Wil08a, WC09, Wil11, WC13, Wil16b, Wil17a, Wil18b, Wil19e]. Regression [Wil19f, Wu04, YSVS19]. Regression-Discontinuity [Les11]. Regressions [Lip13a, Lip18]. Regressors [AO08]. Regular [KY08]. Regularized [Lip18, RTS16]. Reinterpretation [Day17]. Rejection [RL10]. Rejoinder [RH03]. Related [Ng06, RS17, RNN13b, TZA08, ZZ15]. Relation [AS05]. Relations [AGKJ19]. Relationship [KRP02, KS09, Mic11, Oya11, XYJ03]. Relationships [AN04, Cha24]. Relative [Dye11, LZW14, Saw02c, TZZD08]. Relevance [Lov19b]. Relevant [Fra15]. Reliability [AK18a, AA20, BP16, BJJ22, BJJ23, Ben10, CT05, Cha22, EYS^{+}07, Kan06, KBKBG07, KJ14a, KJ14b, KMA21, NRSGRN02, PD13, PJ07, Rao14, RWS09, SSAH18, Spe04, Wal06, Wal08, ZZ15]. Remains [Fra19]. Remark [BRT17]. Remarks [ZK16]. Remittances [Fer12]. Removals [FEB13]. Renewal [TZA08]. Repairable [AE11]. Repeated [ATvBB17, AM14, BLNA^{+}22, FBR^{+}08, ICS07, LHB04, LH04, LL07, MBWZ07, ON14, PW04, SLLL11, SB09, SLA16, TW07, VSS16]. Replenishing [BPWR10]. Replications [MSK^{+}11]. Reply [Lui07]. Reported [OL03]. Reporting [Spe04]. Required [KBKBG07, MSK^{+}11, Wal08]. Resampling [BPWR10]. Rescue [SJ25]. Research [BB07, BBW11, Che14, CID^{+}14, Gan14, Har10, LS06, Liu14, Saw03b, Ste16, WA07, Xu05]. RESET [SM04]. Residual [JPH23, Lip13a]. Resolves [HXH06]. Resolving [Fay06, ZZ15]. Respondent [CPT04, Pre04]. Respondent-Generated [CPT04, Pre04]. Respondents [BBW11, KV14, SKB11]. Response [ASM^{+}16, ACP05, And13b, Bou10, CBP18, CS12, DW19, FF12, GCH08, GS06, HS08, KSB10, LH09, Liu14, LB19, MS03, Nee04, OD07, OS10, Ola20, ÖÇ17, PS18, PS22, PB13, Raj21, Sah05, SS16, SH09, SS14, TS16, TS18, TMB15, Wal17a, XYJ03, SS15]. Responses [Im20, NKC18]. Restricted [HXH06, MKS19, Man05, Sha11]. Restrictions [Sah09]. Resubstitution [KCCD10]. Resulting [Bea17]. Results [El 11, MK10, TSK04, Wil19e]. Retain [Kan06]. Retained [BS14]. Retained-Components [BS14]. Retest [DT04]. Retirement [Saw04a]. Retrospective [NR05]. Returns [HNWD05]. Revenue [IS24]. Review [RYZ16, Rah17, Rao17]. Revise [Ela08]. Revised [Boo19]. Revisited [Kna07]. Revisiting [PI08]. Revolution [RL10, RM16]. RGI [CPT04, Pre04]. RGLM [CAM08]. Rice [KK13, SK07]. Ridge [AKQ10, Dor16, HWS^{+}24, Kha12b, KI14, KI16, KB16, LOA15, LAA17, NSSO14, SA10]. Right [LS05, MSS20a, MRKM^{+}17, Ser10]. Right-Censored [MRKM^{+}17]. Right-Skewed [MSS20a]. Right-tailed [LS05]. Rising [Weh03]. Risk [GVCM18, IJ06, KRP02, Lui06, MMM07, Ozg16, PN03, RRS07, Zha05, Whi07]. Risks [KJ06]. River [BB11]. RMSE [Har18]. Road [Kor16]. Robust [ASS19, ARM10, ANAS16, ASR^{+}16, BS11b, BLNA^{+}22, CA08, CB12, CCC19, CAM08, DY20, EM14, FF13, FKL^{+}03, FRJ02, Gan14, HSA18, JK14, KWAO04, KAF05, Kho04, KS06, Kib08, Kin03, LHB04, LOA15, MRMH17, MN05, MN17, Mic13, MK08, NSJ^{+}24, NW09, OYK^{+}12, OA11, SL17, SA10, SAZ17, SH15, SZH^{+}18, SRA17, WK02b, WK04, WE05, Wil06a, Wil11, WC13, Wil15a, Wil16a, WH16, Wil17b, Wil18c, Wil18d, ZA11]. Robustness [Alk13, AO08, CS09, LS17, MSS19, RRHA11, SJAZ14]. ROC [Bak07, LC20, PA14]. Roger [PA07, PMZ08]. Role [Har10]. Root [ANM15, AO15, Fuk07, KCA^{+}17, Raj21, SH15, ZA08, ZA11]. Roots [El 11]. Rotation [Fin11, KS09]. Rotterdam [PK06]. Routine [Alo04]. RSREG [ASM^{+}16]. RSS [ASAIJ12, AASIJ10]. Rule [Ebr06]. Rules [Kan06, Kho05, MA17, RR24, Saw09]. Rulon [Wal06]. Run [Ela08, EC08, KN05]. RUS [FKP^{+}25]. Ruxton [NKC18].

s [AZ17, SPT12, SGM16, AAK10, AKA11, BPWR10]. S-PLUS [AAK10, AKA11]. S-SMART [BPWR10]. S. [PM03]. Salaries [LF11]. Salary [LF11]. Sale [Saw03c]. Sales [WW10]. Salomon [DSCS08]. Same [Heo02, Man05]. Sample [AB11, AOY11, BAN06, Bic11, BW05, BJ11, BE17, CCT09, CKPJ18, DBH02, EYS^{+}07, Fah02b, FR04, GT18, HW11, Heo02, Hit09, Hur03, IB24, KBKBG07, KK20b, Lip10, LW05, NW10, NM19, OO05, OYO16, PD05, Pas09, PM10, Per03, Pre04, Qum13, RYZ16, Rah17, RTA22, Rav17, Raz02, Ree09a, SH14, SKO09, ST02, SS04, Saw02c, SM02, Sha08a, SGM16, TD03, VS08]. Sampled [EO13]. Samples [ADSBH02, BPWR10, BE06, BK22, DRTW17, DWT19, Fay02, FKL^{+}03, KS14, Kho07, LAM16, LFG19, MS14, MBS17, Mal02, MOD07, OY17, SMSN20, ST02, Sha08b, Wal16a, NT25]. Sampling [ADAR09, ADAAM09, AHAO09, AOAH11, AOCAN19, ASS04, BB19, BC19, Bou10, CBP18, CS12, CF03, CERA24, Eid04, GSP24, GCP24, GW09, HS18, Kad16, KR16, KSB10, KIR12, KV14, KC20, LK13, Lu16, MBS16, MC11, MEH08, MSS19, Odi12, Opd03, PS18, Pap16, Rao11, RNN13a, SQ15, SS12, SS16, SS24, SH07, SH09, SK10, SKB11, STJ11, SSK13, SSS16, SK18, SU16b, TC12, TL14, TuAH19, VZ20, YTP24, YAS14, Dan25, DSK18, QS12]. SAS [ASM^{+}16, ANAS16, ASR^{+}16, CKP12, KH16, LCW^{+}17, Liu09, Lu16, NS07, PS02a, RP03a, SAZ17, ZD10]. SATLINS [Udo13]. Saturated [Udo13]. Sawilowsky [RH03]. SBIBDS [DDM07]. Scale [ADAR09, AAK10, BAN06, BK22, CF03, CC19, Geo14, GK20, HNWD05, Heo08, JA18, Kho04, MOD07, MC07, OHQ^{+}20, ZK05]. Scale-Level [ZK05]. Scaled [PS17]. Scales [GZ07, Lip17a, ZGZ07]. Scaling [Lip13b, WWB12]. Schedule [MADT03]. Scheffé [ZL04]. Scheme [BP16, KSB10, VZ20]. Schmid [Bea17]. School [Kle19]. Schubert [Saw02b]. Science [Fra15, LS06, Saw03b, bZbWA13]. Sciences [HKP17]. Scientific [Har10]. Scientists [ODLR17]. Score [Bea11, BS14, Bea17, LMRM^{+}20, LKW^{+}19, LZ07, Cha22]. Scores [Bar15, Cha22, SS09]. Scoring [IJ08]. Scrambling [CS12]. Scree [Kan05]. Screening [BC05, Bos09, EK03, Ura20, WCW08]. Scripta [Saw07]. Search [SJ25, Tur09]. Searches [GT10]. Seasonal [El 11, KE12]. Seasonality [NIN16]. Second [BH06, HB08a]. Second-Order [BH06, HB08a]. Sectional [SH15]. Security [DNUH22]. Seed [HS11]. Segments [LC05]. Segregation [BR03]. Select [EAS13]. Selected [AA12, PKRK25, Wil06a]. Selecting [WF09]. Selection [AL15, DK19, EME10, EK03, FR03, Fuk05, GG09, HS11, KK16, KY08, KIR12, KHC05, LEKG10, ODLR17, PD05, PK06, PP17, RWS09, SKO09, SSMS10, SBY07, Tor18, UAD16, WF09, ZJ08]. Self [DEH^{+}06, Min13]. Self-Organizing [DEH^{+}06]. Self-Rated [Min13]. Sell [RR24]. SEM [Koc16]. Semi [AAI15, CGH15, DEE17, KI10, KMA18, PJ07, SKO09]. Semi-Bayesian [KMA18]. Semi-Parametric [AAI15, CGH15, DEE17, SKO09]. Semiparametric [HY03, SLA16]. Semipartial [AKP08]. Sensitive [Bou10, HS18, PS22, SS14, Wil08b]. Sensitivity [LC20, SR07, TS16, WCW08]. Sensitivity-Specificity [LC20]. Sensitizing [Kho05]. Sensory [Nan05, SLC17]. Separate [AB11, SK10, TL14]. Separate-Variances [AB11]. Sequences [VJ11]. Sequential [BA11, LS16b]. Serial [LeB16]. Series [ANM15, BD20, BP16, BOO10, DY20, DDM07, EN23, GP04, HML07, KC06, KJ14b, KCA^{+}17, MDS03, NXFN19, NE16, NIN16, RAJ18, RB09, Saf08, SSMS10, SHA06b, SK11]. Server [MADT03]. Service [WBSMI09]. Set [ADSBH02, ADAR09, ADAAM09, AHAO09, AOAH11, ASS04, BC19, BK22, Bou10, CBP18, SMSN20, ST02, SS04, SQ15, TuAH19, YAS14]. Setting [XZ05]. Setup [Pra10a]. Several [BZ11, BF14, GVCM18, KBKBG07, SJAZ14]. Severity [IJ06, Sie05]. Shape [Geo14, LM17, PA14, SE12]. Shape-Based [LM17]. Shaped [MSS20a]. Shapley [Lip06]. Shift [MC07, Saw05]. Shifted [AM10]. Shifting [HB08a, Sen02]. Shiny [MBS17]. Shootouts [Hur08]. Short [Ela08, EC08, GZ07, KN05]. Short-Run [Ela08, EC08]. Should [Kna18]. Showing [PD11]. Shrimp [CdlTE22]. Shrinkage [Asa16, JK14, OHQ^{+}20, Pra10a, Pra10b, RWS09, Wal07]. Sided [BC02, BS11b]. Sides [Man05]. Siegel [LS17]. Sieve [Alo04, GW09]. Sigma [RB11, Rav17]. Signal [Bha08]. Signals [BK09b]. Signed [MOD07]. Signed-Rank [MOD07]. Significance [Goo13, Hei06, Kna02, Man13, RL02]. Significant [HM10, Saw13]. Simple [ASM^{+}16, BB16, BE06, Bak07, BY11, BC19, BH06, CBP18, ÇE14, DEE17, EAN05, EAM09, FR04, Kad16, KC20, Lui06, MMH19, NKC18, PW04, RN18, TD03, Wil08a, YTP24, Whi07]. Simulating [Hea02, Hea04, KH16]. Simulation [AAK22, ARM10, AKQ10, AZN20, BC05, CC09, DAC02, Dye11, Dye16, Fuk07, GS19, GT18, HL16, HKP17, HWS^{+}24, Hol06, KSS10, Les11, LMRM^{+}20, Mic13, MK10, MSK^{+}11, OON24, Rus08, SA02, Tak17b, YOO14]. Simulations [BE17, HS11, Koc16, RM16]. Simultaneous [Oya11, Wil17b, Wil17a]. Simultaneously [SS15]. Singh [Ahm20]. Single [BO13, CKP12, Jay17, Koc18, LR03, LFG14, LFG19, Lui10, NS07, Opd05, PC15, PC18, SS12, Saw02c]. Single-Case [BO13, LFG14, LFG19, PC15, PC18]. Single-Stage [Lui10]. Single-Study [LR03]. Single-Subject [CKP12]. Singular [Lip16]. Site [MEH08]. Situations [Gan14, ZL04]. Six [Fay06, Pet02, RB11, Rav17]. Size [AKP06, BH16, BJ11, BE17, CdlTE22, CKPJ18, EYS^{+}07, GT18, Hit09, KAF05, LW05, MPC^{+}16, Mir08, PD05, PS22, PM10, Qum13, Rav17, SH14, Saw09, SHA06b, SM04, TP16, VA16, Wil18d, ZA08, ZA11, ZRB23]. Size-Biased [Mir08, SHA06b]. Sizes [AB11, BF09, BF11, BF13, BF14, CP13, KBKBG07, LR03, Mec03, MC07, SM02, Ste16, Wal05b]. Skew [AT09, BB06, BC19]. Skew-Normal [BB06]. Skewed [AZ17, HZ16, LA18, LL16, MSS20a]. Skewness [SP14b]. Skip [SU16b]. Skipped [Wil15b]. SkSP [HBA20]. SkSP-R [HBA20]. Slightly [Bea11]. Slope [BY11, EAN05, MK08, YD17]. Slopes [KM02, LMME09, WC09]. Small [AB11, BPWR10, Mal02, Mar12, MOD07, NW10, Ree04, Ree09a, RB09, SAA19b, Sha08b, VM09, Wal16a, WBSMI09]. Small-Sample [NW10]. Small-to-Medium [Mar12]. Smaller [AO15]. Smallest [Wil19e]. SMART [BPWR10]. Smoked [ANA12]. Smooth [YS16]. Smoother [WE05, Wil15a, Wil17a, Wil19c]. Smoothers [Wil08a, Wil11, Wil16b]. Smoothing [Die06, IB24, LIT24, Tho10]. SMOTE [KS25b]. Social [Fra15, LS06, Saw03b, WBSMI09]. Socioeconomic [RRB08]. Socratic [Ber09]. Software [AAK10, BO13, LMS15, ODLR17, SA12]. Solace [Die09a]. Solution [DS16, Lip09, LM19, LZ07, Lov19b, Mic11, Ola20]. Solving [KJMA18]. Some [ADAAM09, ACP05, AO08, BE06, BK09a, Che14, Con02, Dem05, Eid04, El 11, HB08b, HB12, HNG19, Hus07, KS14, KY08, Kib06, KB16, Kna02, LFO07, LAA17, NW10, NIN16, ÖK17, PJ07, Pra20, RK13, SA10, Udo13, VSS16, Wil17b, Wil17a, Wil19f, ZK16, AZN20, DS16, MN21]. Sort [NRSGRN02]. Source [BD20, MW07]. South [SA14]. Sp [DK19]. Space [BS14, Odi12, Ura20]. Sparse [WS19]. Sparseness [HL16]. Spatial [GV03, KJI10, Kor16, RJR14, WB02]. Spatially [OA11]. Spatio [CS08]. Spatio-Temporal [CS08]. Spearman [Rus08, Wal06]. Special [FS15]. Specific [LH09, Tur09]. Specification [GT10]. Specificity [LC20]. Specified [Wil17b, Wil17a]. Specifying [YS14]. Spectra [Res08]. Speed [SA14]. Spent [WCW08]. Spherically [KK20a]. Spline [LT07, LH09]. Spline-Based [LT07]. Splines [LX06]. Split [BZ09, Kra18, Lip12, PA07, PMZ08, Wal06]. Split-Plot [BZ09, PA07, PMZ08]. Spread [OYK^{+}12]. Spring [BPZ12]. SPSS [Abe15, FZ02, Liu09, PS02a, Tay11, Wal05b, Wal16a, WS16a, Wal17b, ZLB19]. Square [ANM15, DW19, Dye11, Kha12b, SAZ17, WY03, YT10, ZA08, ZA11, Web07]. Squared [AO15, AKP08, Die06, LH14, SY19]. Squares [AKQ10, Die09b, Kha13, NW09, Tof08]. SRS [ASAIJ12, AASIJ10]. Stability [Ste16]. Stabilized [GW09]. Stabilizing [GP04]. Stable [MSS07]. Stage [Ela08, EC08, Lui10, LOA15, NRSGRN02, TZA08, TS18, TX11, RWS09]. Stan [AAK22, AK20]. Standard [MK08, TK06]. Standardized [SB09, Wal05b, ZA08, ZA11]. Standby [BP16]. STAR [YS14]. STATA [MEH08, PS02a, Liu09, Liu14]. State [AA12, DS02, Lip13b, MADT03, WCW08]. States [PKRK25, TL15]. Statistic [DK19, LSAA02, LL16, Opd05, SV12]. Statistical [BH16, Ber05, CMR02, CID^{+}14, Har10, Hei06, HBA20, KBKBG07, KKST05, KK20b, LS16a, LMS15, MYY^{+}06, NH12, OL03, ODLR17, Pos02b, PC14, PKRK25, Qum13, Rao17, Rod10, Sen02, SMC16, SJ07, SK09, SA07, SE12, SL16, TL15, UAD16, Udo20, Web07, XZ05, ZZ15, Kna19b]. Statistically [LR03, Saw13]. Statisticians [Ber05]. Statistics [AA21, AK21, CCT09, DRTW17, EM14, Fah02b, FF13, Har18, Har19, Hig04, KBKBG07, KWOF02, KWAO04, KKW^{+}06, KKK19, KC20, LL16, Lov19a, PW04]. Status [SZH^{+}18]. Statuses [Saw13]. StatXact [Pos02a]. Steady [Lip13b, MADT03]. Steady-State [Lip13b]. Step [EAM09, LRT20]. Step-Stress [EAM09, LRT20]. Stepwise [HM10]. Stereotype [Liu14]. Stochastic [AO08, BZ09, Jay17, KI10, Lip13b, SS14]. Stock [BA11, DR14, PDC13, SS18]. Strategies [MK08, PS02a]. Strategy [Har18, TS18, Zum02]. Stratification [CERA24, DR18, MT24]. Stratified [CEI24, GSP24, GCP24, SS04, SK10, TC12, TL14, TS18, DSK18]. Stratum [RYZ16]. Street [MPC^{+}16]. Street-Dwelling [MPC^{+}16]. Strength [BJJ22, BJJ23, KJ14a, Rao14, Wil11, Wil15a]. Stress [BJJ22, BJJ23, EAM09, KJ14a, LRT20, Rao14]. Stress-Strength [KJ14a, Rao14]. Striving [NKC18, RN18]. Structural [AAA15, CE15, CB12, CID^{+}14, Fan10, Fuk05, FB10, Kib08, Mic11, RM16, RMS17, Koc18]. Structure [BII11, Hit09]. Structure-Based [BII11]. Structured [WWB12]. Structures [YD17]. Student [AK10, ASS19]. Student- [ASS19]. Students [Hal04]. Studies [AOBCP16, AZN20, FF08b, HKP17, Har18, Har19, KRP02, Lee11b, Lee13, LH13, Lee15, LF11, LLHT03, LH09, LX06, LP06, MSK^{+}11, PC15, PC18, RNN12, TSK04, Wal04, XGY^{+}08]. Study [ARM10, AK10, BB11, DAC02, Dye11, EME10, Ele17, GG09, GS19, GT18, HL16, Kan05, LR03, LMRM^{+}20, LHH11, LHO^{+}18, LAA17, ML24, Mar12, MSS20b, MK10, MIN18, NKW^{+}19, NM17, NM19, PH25, RSQ20, Rus08, SKB15, SM17, SA02, SFA16b, WSN15, ZT12, AZN20]. Studying [Har19]. Sub [Dan25, KV14, SKB11]. Sub-Sampling [SKB11, Dan25]. Subband [LS16a]. Subject [CKP12]. Subjects [PA04]. Subset [DK19, Nan05, SW08, SBY07]. Subsets [DP05, Sha06a]. Substantive [OL03]. Success [AK18a, Raz02, Wil19c, Wil19b, Wil19d]. Success/Failure [AK18a]. Successive [SS16, SS24, SKB11, SSS16, SK18]. Sultan [LIT24]. Sum [Heo02]. Sums [Die06, Hur03]. Sun [LHO^{+}18]. Sunspots [AS05]. Super [SSS21]. Superpopulation [Pap16]. Support [DK19]. Supporting [OL03, WB06]. Surface [ASM^{+}16, ACP05, MS03, Raj21]. Surveillance [RRB08]. Survey [BBW11, FR04, LHO^{+}18, Lor19, Pap16, Pér07a, Pér07b, TD03, VS08, ZLB19]. Surveys [Kum13, Pre04, RTA22, Sah05, SGM16, Wan03, YHPR17, ZS17]. Survival [AK20, AM14, AFL19, AAH^{+}19, FR04, ST14, SSMK22]. Symbols [DSCS08]. Symmetric [Gan14, KK20a, MŞ09, NSSO14, Udo13, YS16]. Symmetry [IT18, Mic13, MPS11, OKW^{+}02, PS02b, YT10]. Syndicated [Wan03]. Synthesis [MSK^{+}11]. Synthetic [AKK24]. Systat [PS02a]. System [AE11, AT09, BP16, Bos09, GR11, IJ08, Leh15, PK06, RAJ18, SL16]. Systematic [LHH11, RMS17, STJ11]. Systems [Hea02, KJ14a, Leh15, ML24, SM04].

t [Wil10]. T2 [MIN18, Nan02]. Tables [HL07, IT18, JADB10, KRP02, LA14, PD11, Pos02a, Sha08b, WY03, YT10]. Tailed [MŞ09, LS05, MSS20a]. Tangent [Udo13]. TANH [Udo13]. TAR [EN23]. Task [MEP09]. Tau [Wal03, Wal16a]. Tax [IS24]. Teachers [ODLR17]. Teaching [DSCS08, Saw04b, Sto04]. Technical [ANAS16]. Technique [AKQ10, Dan25, IB24, MESSE16, MMM07, OY17, Sah05, SSMS10, SVS16]. Techniques [JA18, Mar12, MC07, MC08, Nav07, Oya11, ÖÇ17, ZS17]. Technology [Hig04]. Tehran [PDC13]. Telecommunications [Opd05]. Temperature [AS05, MF11]. Template [CKP12]. Temporal [CS08]. Tendency [NKC18, RN18]. Term [KCA^{+}17, Mey16]. Terms [AO08, DK19, Rah14, SSMK22]. Terry [Mey16]. Test [AS15a, AW16, BR14, BK09b, BF11, BF13, BF14, BG15, Bic11, CS08, CFSF10, Cli11, CS09, Day17, DRTW17, DT04, FF08a, FF13, FZ09, Gao04, GJ17, HH16, HL03, KPR14, KR16, KWAO04, KE12, Kon13, Kra18, KCA^{+}17, KK20b, Lee11a, LFG19, LT07, LL16, LRT20, MS14, MBS17, Mal02, Man13, MSS07, MOD07, Mic13, MPS11, MK08, NS07, Nee04, Neu09, Neu10, NKC18, OKW^{+}02, PS02b, Pet02, PTC18, RK13, Sah09, SM02, Saw05, SH15, SM04, SS09, TGhC^{+}05, VO03, WS16b, WA07, Wil06a, Wil07a, Wil08b, Wil12, Wil16a, XYJ03, ZRB23, AB11, BRT17, NT25, PA07, PMZ08, Wil19a]. Testing [AS05, And13a, And13b, BH16, BKS12, BC02, CMR02, Die11, EAM09, FF16, GMS05, GK20, Har10, HA14, JADB10, KRR13, KRR14, KOW13, Kna02, KZ08, LH13, LHB04, LCW^{+}17, LS05, Lui10, Mec03, MK05, NRSGRN02, Nav07, OY17, Opd05, Ots05, Ozg16, PNG08, Pau05, PS02b, Pra10a, RL16, RL10, RL02, Saw03a, SU16b, WWB12, Zha05, PB13, ZK16]. Tests [ADARAM05, AM10, AOCAN19, ASAIJ12, AZ17, AASIJ10, BM05, BZ09, Ben10, BI02, Ber17, BS11a, BZ11, Bla02, BW05, Cha24, CDPT19, Con02, CS04, CM02, CWBK07, DWT19, Die09a, Ebr06, Fay02, Fay06, FKL^{+}03, GK04, GCH08, GK20, GCAC07, GT18, HL07, Hei06, HK10, HH16, KKST05, KM02, LAM16, LHB04, LFG14, LS17, MK10, MBWZ07, NH06, NKdG^{+}16, NKW^{+}19, NIN16, ÖK17, Opd03, OYK^{+}12, Per03, Pos02a, Rah17, Raj21, RRHA11, Rao11, RM07, RBAR18, SLC17, SRW21, Sha08b, TW07, VSS16, Web07, WS09, WA07, Wil07b, YSVS19, ZK05, Cha22, El 11, RR24]. th [KDMAA18]. Their [Fah02b, Goo02, HNG19, KB16, ÖK17, BK09a, BPZ12]. Themselves [LZ07]. Theoretical [El 11]. Theory [CF03, DW19, FF12, Kum13, LS06, Nee04, SH09, SFA16a, Tak17b, SJ25]. theory-based [SJ25]. There [NR05, WC13, Wil16a, Wil17b, Wil18a, Wil19f, Rod10]. Theta [EYS^{+}07, ZGZ07]. Thing [Kes15]. Think [Saw03d]. Third [Hea06]. Third-Order [Hea06]. Thornton [HS18]. Three [AOCAN19, Bea17, KSM13, KM14, LAM16, Mic11, PM14, Wal05b, WA07, WC09, SG21]. Three-Level [PM14]. Three-Parameter [AOCAN19]. Threshold [LH09, MW07, SA11]. Thresholds [KS25a]. Thumb [Saw09]. Thurstone [Lip13b]. Tied [WC13]. Ties [EO13, Fay06]. Tightened [SP16]. Time [AAI15, ANM15, AKO03, AE11, BD20, BS04, BOO10, CGH15, DEE17, DY20, FRJ02, GP04, HML07, KC06, KCA^{+}17, LW05, MRKM^{+}17, NXFN19, NIN16, PNG08, PP17, RB09, Saf08, Seo19, SSMS10, ST14, SK11, TZA08, Tig03, WCW08, ANO^{+}25]. Time-Averaged [LW05]. Time-Modulated [FRJ02]. Time-Series [HML07]. Time-to-Failure [DEE17]. Time-Varying [MRKM^{+}17, Tig03]. TIMSS [Lor19]. Tools [WB06]. Topp [AGKJ19, OON24]. Tournament [Saw16]. Toxicity [LH09]. trade [PW24]. Trading [RR24]. Traditional [FBR^{+}08, MSS20b]. Traffic [Kor16]. Training [Ber05]. Trajectories [LLHT03, LBC^{+}19]. Transect [AE12, Eid04, EAS13]. Transfer [Udo13, Udo20]. Transform [Nan02, NSSO14, Pet02]. Transformation [BOO10, BS14, Bea17, GP04, SSMS10]. Transformations [Hea04, QS12, SS09]. Transformed [AAK22, ANM15, DGJ16, GRM20, KKW^{+}06]. Transformer [NSJ^{+}24]. Transforming [KWOF02]. Transition [DT04, SIA11, YS16]. Transmuted [AAA15, NAY^{+}18]. Treating [TD03]. Treatment [AK04a, AZN20, CKCL18, CWBK07, KOW16, MA06, MK05, Par10, TX11]. Treatments [Saw13, VJ11]. Tree [Oya11]. Trees [Hur07]. Trend [CM02, CS09, KCA^{+}17, Tho10]. Trends [Tho09]. Trial [AKO03, Lui10, SZH^{+}18]. Trials [BC05, BLEL02, BW05, HL07, Lui06, Ree06, Tho09, Whi07]. Triangle [DS25]. Trigonometric [BOO10]. Trimmed [BLNA^{+}22, BS02, HM08, KKW^{+}06, MMH19, MC08, WK02a]. Trimming [KWOF02, RR07, SR07]. Trip [MEH08]. Triple [LIT24]. Triples [PS02b]. Triply [Kit10]. Trivials [SY02, Saw03c, Saw03d]. Triwizard [Saw16]. Trouble [LR03, SY02]. True [Cha22]. Trump [LOCMS13]. Truncated [ASK16, AOCAN19, AA14, AT09, kAEAE21, KJ12, Rao11, Tor18]. Truncation [ML03]. Truth [RL16, ZK16]. Tukey [LS17]. Tumor [Sha11].

Twelve [Sha11]. Twenty [Fah02b]. Two [Abe15, AOY11, AKO03, AK04a, AŞ15b, BG15, Bic11, Bla02, BC02, DRTW17, DWT19, DBH02, EO13, Ela06, Ela08, EC08, EK03, Fay02, FA13, FZ09, Gao04, Geo14, HH16, IT18, JADB10, JVA10, KM14, KY08, Kel05, KAF05, KJI10, Kon13, KV14, KK20b, LFG19, LZ07, LOA15, MS14, MBS17, MBS16, MKA23a, MKA23b, Man05, MT24, MC08, MW07, NKdG^{+}16, NKC18, OR02, OO05, OS10, OY17, OHQ^{+}20, PM14, Per03, RYZ16, RSQ20, RTA22, RWS09, Ree04, Ree09b, RP03a, RP03b, RN18, RL02, Sah09, Saw02b, Sha08b, SFA14, SFA16a, SSS16, SK18, SPB19, SP14b, TC12, TZA08, TS18, TW07, VZ20, VM09, Wal05a, Wal15, WC09, Wil11, Wil12, Wil16b, Wil17b, Wil18a, Wil19f, ZD10, PS18]. Two- [DBH02, PM14]. Two-class [ZD10]. Two-Component [FA13, SFA14, SFA16a]. Two-Factor [KY08]. Two-Group [AKO03, AK04a, KAF05, MC08]. Two-High-Threshold [MW07]. Two-Level [EK03, Kon13]. Two-Occasion [SSS16, SK18]. Two-Parameter [AŞ15b, Geo14, PS18]. Two-Parametric [SPB19]. Two-Phase [MBS16, VZ20]. Two-point [ZD10]. Two-Sample [AOY11, Bic11, OO05, Per03, RYZ16]. Two-Sided [BC02]. Two-Stage [Ela08, EC08, TZA08, TS18, RWS09]. Two-Wave [LZ07]. Two-Way [IT18, JADB10, RP03a, RP03b, Sha08b]. Type [AB11, AOY11, Asa16, AM19, Bak07, GSP24, GCP24, Kad16, KPR14, KR16, KSB10, LFG19, MKS19, MT24, MMM07, OE12, PJ14, PB13, Rah17, SPT12, SVS16, TC12, Tig03, TSS20, YTP24, LB10, TL14, AA21, And13a, And13b, BM05, BP16, Fay06, FEB13, FKL^{+}03, Fra15, GRM20, Har19, KRR13, KRR14, Kes15, KM02, KJ06, KJK09, LS06, LB16, LRT20, OON24, Ozg16, PA04, PA07, PMZ08, PPS14, PP17, PS17, Pra14, RBAR18, SM17, SW08, SFA16b, WA03, WA07]. Type-I [AA21, LRT20]. Type-II [GRM20, KRR14, KJ06, KJK09, PPS14, PS17, Pra14]. Types [Bea17]. Typical [Wil18d].

U. [PM03]. U.S. [PM08]. U1 [Wal05a]. Ultrahigh [Ura20]. Umbrella [XYJ03]. Unbalanced [CBP18, WA07]. Unbiased [AKQ10, Ben17, Ela06, Kha13, MBS16, SGM16, Tho10, Wal17b]. Uncommon [SM02]. Unconditional [Mal02, OO05, RBAR18]. Underlying [Mos12]. Understanding [DEH^{+}06]. Underwriting [ZT12]. Unequal [BF09, BF11, BF13, BF14, BC19, Ozg16]. unidimensional [SH09]. Uniform [YT10]. Unimodal [FZ09]. Unique [Odi12, OAC03]. Unit [El 11, Fuk07, KCA^{+}17, RB11, Raj21, SH15]. Unit-Root [Fuk07]. Units [KS10]. Univariate [BD20, Dem05, MRMH17, MBWZ07]. Universe [GJ17]. Universitatis [Saw07]. Unknown [KS25a]. Unpaired [DWT19]. Unprotected [BM05]. Unreliability [CID^{+}14]. Unreplicated [Per06]. Unrestricted [Man05]. Unstable [WSN15]. Unsupervised [LS13]. Upper [HNG19, KD18]. Upside [MSS20a]. Urban [SJ25]. Use [CD09, DSCS08, LP06, LMS15, OS10, Per03, PE08, Smi09]. Used [CE15, Dem04b, SWM19]. Useful [GT10, SPB19]. User [CAM08]. Using [ADSBH02, ADAAM09, AAK22, AM14, AAK10, AHAO09, AKQ10, AOAH11, AOCAN19, ASS04, AS05, ASM^{+}16, ACP05, ASG16a, Bar15, Bos09, Bou10, BA11, CT05, CDPT19, CC09, DW19, Day03, DP05, Ebr06, Eid04, EAS13, FF08b, FZ02, FR04, Fuk07, GD02, GZ07, GCH08, HM08, HNWD05, Har18, Hei06, HML07, HS08, Kib08, KV14, KC20, KHC05, KMG06, LM17, Leh15, LA18, LF11, LLHT03, LHM14, Liu09, LOK11, LHH11, LK12, LWZK16, MAP08, Oya11, PZ11, PD05, PK06, PS22, PN10, PN03, Pér07b, Per06, PM03, PNZM24, Qum13, RRB08, RP03a, RM07, SAA19b, SKO09, SQ15, SM02, ST14, SGM16, SVS16, SY19, Sta02, SBY07, SK13, SA12, SP16, SS18, TS11, TS16, VS08, WA07, Wil07a, WH16, Wil19c, WZ17, Xu05, ZCB^{+}17, ZLB19, BK22, BE17]. using [CS12, DSK18, EN23, GSP24, IB24, IJ08, JR16, KK16, KS25a, Lor19, MRMH17, MKA23a, MKA23b, MEP09, NXFN19, OY17, PS18, RSQ20, RTA22, SK18, SG21, Su11, SKSV25]. Usual [LL16]. Utility [Heo08]. Utilization [Dan25]. Utilizing [BY11].

Vacations [MADT03]. Validation [Wal04]. Validity [AS16, CS09, NRSGRN02, WFC15]. Value [And13a, And13b, CdlTE22, Die09b, Lip06, Lip16, RRB12, SGM16, Wil19c, PB13]. Values [AGKJ19, AA20, AF23, BB04, BB05, CCT09, Fay02, Goo13, Hea03, HNG19, KDMAA18, KD18, MESSE16, Man13, NH12, OO18, PA04, PA07, PMZ08, Raj21, Smi09, WC13, ZCB^{+}17, DSK18]. VAR [MSS07]. Variability [BS11b, Ele17]. Variable [AZ17, BB16, BB19, CBP18, Eic05, FR03, KK16, LT07, Lip12, Ola20, QWL^{+}16, SL17, SV12, SS14, SK13, SP14a, TS11, TuAH19, Wil19e]. Variables [AL15, Bou10, CS12, CD09, CB12, DR18, Dan25, Gan12, HS18, Hol06, KK13, Koc16, KMG06, Lee13, Lip17b, LM19, Liu14, LB19, MKA23a, MKA23b, OD07, OS10, RSQ20, RTA22, RC17, SS12, SK07, SSK13, SP14b, SS18, TZZD08, Wil18a, Wil19e, XYJ03, ZC13, ZD10]. Variance [AHAO09, Cam03, Cam09b, Ela06, GA17, GL06, GP04, KSS10, LCW^{+}17, LS05, MC11, Pér07b, RP03a, RP03b, Saf08, Sha08a, SS13, SS24, SK13, SRA17, TTBG06, Wal06, WPN^{+}18, ZT06, Cha22]. Variances [AB11, AKP06, CFSF10, MN05]. Variants [OYO16, Wal05a, YS16]. Variation [BKS12, Cam12, LH09, SK13]. Varied [WFW02]. Varimax [KS09]. Various [Fay06, Mic13, OGKH17, Wal06]. Varying [AAI15, MRKM^{+}17, Tig03]. Vector [DK19, Fog15]. Vehicles [NSJ^{+}24]. Verifying [NM19]. Version [MA18, VO03]. Versions [ZGZ07]. Versus [Cam07, Cam09a, Cam09b, Die06, PS02b, RNN13a, WA07, Alk13, Bos09, Goo13, Lip19, Man13, Rod10, Wil10]. Very [AB11]. Veterinary [KHC05]. Via [DP21, LC05, WE05, Wil08a, Wil12, CAM08, Ola20, PD13, Wil08b, Wil11]. Vibrio [RRS07]. Vietnamese [HP03]. View [RNN13b]. Viewpoint [Pra12]. VIF [Ura20]. VIF-Regression [Ura20]. Violating [WWB12]. Violation [Lan13]. Violations [LH04, MK05]. Viral [LLHT03]. Virtual [BPWR10]. Vision [NSJ^{+}24]. Visually [Wu04]. Volatility [DR14]. Volume [Saw07].

Vs [Man05, Die09b, HK10, KPR14, MOD07, Nan02, WA03].

waiting [ANO^{+}25]. Wald [AZ17]. Wallis [BB04, Hea03]. Warde [Sah05]. Warner [GS06, HS08]. Warning [Sha11]. Was [Ser10]. Wave [LZ07]. Wavelet [Aly07, NSSO14, Qu05, Wu04]. Wavelets [AS05]. Waves [LZ07]. Way [BB04, Hea03, IT18, JADB10, KMG06, NS07, OAC03, RR03, RP03a, RP03b, Sha08b, ZA08, ZA11, PH25]. Web [CAM08, MBS17]. Web-based [MBS17]. Weibull [AAA15, AALF14, AOAH11, ALF16, AM19, BK22, ESAH20, Geo14, KMA18, LFO07, MEP09, MA18, NAY^{+}18, PA14, RRS23, SMSN20, SA14, SFA16a, SKSV25]. Weight [BHH12, QWL^{+}16]. Weighted [AE12, AJ17, BPZ12, DY20, Heo08, KS06, MJ17, MMH19, MA18, Mug12, MAP08, ST07, ST08]. Weighting [AZN20, HSA18, MC11, SAZ17, KKZB16]. Weightings [CERA24]. Weights [CF03, Heo08, WY03]. Weisberg [Lip13a]. Welch [AB11, MC07, ZRB23]. Westenberg [LS17]. Western [SA14]. Wheat [PKRK25]. Where [OL03]. Which [DWT19, Lip17a, SK09, Wil18a, Wil19b, Wil19e, ZRB23]. White [Lui07]. Whitney [NT25, Saw05, Wil12, Wil19a]. Whole [Kes15]. Wilcoxon [HH16, MOD07, NT25, Saw05, Web07, WS09, Wil12, Wil19a]. Wind [SA14]. Winner [EG11]. Winsorized [JK14, Wil19e]. Wise [PD05, ASG16b, SM17]. Within [PA04, WK02b, Wil05, WH16, FF16, SV12]. Within-Subjects [PA04]. Without [OL03, GK04, RYZ16]. Witney [HH16]. WLS [GZ07]. Women [Pos02b, PM03, Wil04]. Word [WZS11]. Workable [LZ07]. Workers [Lee12]. Works [Saw04b, SC06]. Would [OL03].

X [EC08, KR16, LS13]. X-Alter [LS13]. xgamma [SMC16]. XII [PJ14].

Year [LFG19]. Years [Hig04]. You’ve [Saw03d]. Yuen [MC07].

Z [Ozg16]. Zero [Doy09, Hus07, MBS02, PPC18, PM03, Raz02, SAA19b, YHPR17, bZbWA13]. Zero-Clustered [Hus07]. Zero-Inflated [Doy09, PPC18, SAA19b, YHPR17, bZbWA13, PM03]. Zettabyte [Lov19b]. Zone [PH25].

References

Ademola:2012:MGP

[AA12] Adewara Johnson Ademola and Mbata Ugochuckwu Ahamefula. Multivariate generalized Poisson distribution for interference on selected non-communicable diseases in Lagos State, Nigeria. Journal of Modern Applied Statistical Methods, 11(2): ??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

             jmasm.com/index.php/jmasm/article/view/631.

Arasan:2014:DBC

[AA14] Jayanthi Arasan and Mohd B. Adam. Double bootstrap confidence interval estimates with censored and truncated data. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/739.

Arshad:2020:PRE

[AA20] Mohd. Arshad and Qazi J. Azhad. Parametric and reliability estimation of the Kumaraswamy generalized distribution based on record values. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1093.

Alam:2021:LEO

[AA21] Intekhab Alam and Firoz Alam. Likelihood estimation and order statistics based on Frechat distribution under Type-I censoring. Journal of Modern Applied Statistical Methods, 20(2):??, July 2021. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1115.

Ahmad:2015:SPT

[AAA15] Kaisar Ahmad, S. P. Ahmad, and A. Ahmed. Structural properties of transmuted Weibull distribution. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/778.

Amir:2019:JCF

[AAH^{+}19] Wan Muhamad Amir, Muhammad Azeem, Masitah Hayati Harun, Zalila Ali, and Mohamad Shafiq. JMASM 54: a comparison of four different estimation approaches for prognostic survival oral cancer model. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1071.

Adeleke:2015:SPN

[AAI15] Kazeem A. Adeleke, Alfred A. Abiodun, and R. A. Ipinyomi. Semi-parametric non-proportional hazard model with time varying covariate. Journal of Modern Applied Statistical Methods, 14 (2):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/774.

Ahmad:2010:BAL

[AAK10] Sheikh Parvaiz Ahmad, Aquil Ahmed, and Athar Ali Khan. Bayesian analysis of location-scale family of distributions using S-PLUS and R software. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/513.

AbuJarad:2022:BIS

[AAK22] Mohammed H. AbuJarad, Eman AbuJarad, and Athar Ali Khan. Bayesian inference and simulation for logarithm transformed (LT) method using Stan. Journal of Modern Applied Statistical Methods, 21(2):??, July 2022. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1128.

Al-Aqtash:2014:GWD

[AALF14] Raid Al-Aqtash, Carl Lee, and Felix Famoye. Gumbel–Weibull distribution: Properties and applications. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/728.

Alodat:2010:ECF

[AASIJ10] M. T. Alodat, S. A. Al-Subh, Kamaruzaman Ibrahim, and Abdul Aziz Jemain. Empirical characteristic function approach to goodness of fit tests for the logistic distribution under SRS and RSS. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/512.

Adusah:2011:TEI

[AB11] Albert K. Adusah and Gordon P. Brooks. Type I error inflation of the separate-variances Welch t test with very small sample sizes when assumptions are met. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/550.

Aberson:2015:SPA

[Abe15] Christopher Aberson. SPSS programs for addressing two forms of power for multiple regression coefficients. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/766.

Adepoju:2015:MLE

[AC15] K. A. Adepoju and O. I. Chukwu. Maximum likelihood estimation of the Kumaraswamy exponential distribution with applications. Journal of Modern Applied Statistical Methods, 14(1): ??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/763.

Ali:2006:NBM

[ACB06] M. Masoom Ali, J. S. Cho, and Munni Begum. Nonparametric Bayesian multiple comparisons for dependence parameter in bivariate exponential populations. Journal of Modern Applied Statistical Methods, 5(1):66–72, May 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/227.

Aggarwal:2006:IGF

[ACBM06] M. L. Aggarwal, S. Roy Chowdhury, Anita Bansal, and Neena Mital. Interaction graphs for 4^{r}2^{n−p} fractional factorial designs. Journal of Modern Applied Statistical Methods, 5(2):475–489, November 2006. CODEN ???? ISSN 1538-9472. URL https:/

/jmasm.com/index.php/jmasm/article/view/265.

Anderson-Cook:2005:SGU

[ACP05] Christine M. Anderson-Cook and Kathryn Prewitt. Some guidelines for using nonparametric methods for modeling data from response surface designs. Journal of Modern Applied Statistical Methods, 4(1):106–119, May 2005. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/173.

Abu-Dayyeh:2009:SEP

[ADAAM09] Walid A. Abu-Dayyeh, M. S. Ahmed, R. A. Ahmed, and Hassen A. Muttlak. Some estimators for the population mean using auxiliary information under ranked set sampling. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/418.

Abu-Dayyeh:2009:BPM

[ADAR09] Walid A. Abu-Dayyeh and Lana Al-Rousan. On the BLUE of the population mean for location and scale parameters of distributions based on moving extreme ranked set sampling. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/425.

Abu-Dayyeh:2005:LPC

[ADARAM05] Walid A. Abu-Dayyeh, Z. R. Al-Rawi, and M. MA. A. Al-Momani. Local power for combining independent tests in the presence of nuisance parameters for the logistic distribution. Journal of Modern Applied Statistical Methods, 4(1):75–80, May 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/169.

Abu-Dayyeh:2002:DFE

[ADSBH02] Walid A. Abu-Dayyeh, Hani M. Samawi, and Lara A. Bani-Hani. On distribution function estimation using double ranked set samples with application. Journal of Modern Applied Statistical Methods, 1(2):443–451, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/50.

Arasan:2011:MRS

[AE11] Jayanthi Arasan and Samira Ehsani. Modeling repairable system failures with interval failure data and time dependent covariate. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/569.

Ababneh:2012:WED

[AE12] Faisal Ababneh and Omar M. Eidous. A weighted exponential detection function model for line transect data. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/593.

Anyiam:2024:NGN

[AEOD24] Kizito Ebere Anyiam, Emmanuel Ifeanyi Ezerioha, Justin Chukwudi Ogbonna, and Marvelous Ebuka Daniels. New generalized nadarajah haghighi distribution: Characterization and applications. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1177.

Athar:2023:MPR

[AF23] Haseeb Athar and Mohamad A. Fawzy. Moment properties of record values from Rayleigh Lomax distribution and characterization. Journal of Modern Applied Statistical Methods, 22(2):??, October 2023. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1144.

Alzaatreh:2012:GPD

[AFL12] Ayman Alzaatreh, Felix Famoye, and Carl Lee. Gamma–Pareto distribution and its applications. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/589.

Aldeni:2019:GFL

[AFL19] Mahmoud Aldeni, Felix Famoye, and Carl Lee. A generalized family of lifetime distributions and survival models. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1047.

Adams:2008:DMC

[AGHL08] Susan M. Adams, Atul Gupta, Dominique M. Haughton, and John D. Leeth. Data mining CEO compensation. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/390.

Anwar:2019:RRM

[AGKJ19] Zaki Anwar, Neetu Gupta, Mohd Akram Raza Khan, and Qazi Azhad Jamal. Recurrence relations for marginal and joint moment generating functions of Topp–Leone generated exponential distribution based on record values and its characterization. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1034.

Alzaatreh:2014:GLD

[AGS14] Ayman Alzaatreh, Indranil Ghosh, and Hassan Said. On the gamma-logistic distribution. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/691.

Al-Hadhrami:2009:BIV

[AHAO09] Said Ali Al-Hadhrami and Amer Ibrahim Al-Omari. Bayesian inference on the variance of normal distribution using moving extremes ranked set sampling. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/419.

Ahmed:2020:ESM

[Ahm20] Mohamed Ali Ahmed. Extending Singh–Maddala distribution. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1079.

Ajami:2017:PEW

[AJ17] M. Ajami and S. M. A. Jahanshahi. Parameter estimation in weighted Rayleigh distribution. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/937.

Algina:2004:ATE

[AK04a] James Algina and H. J. Keselman. Assessing treatment effects in randomized longitudinal two-group designs with missing observations. Journal of Modern Applied Statistical Methods, 3(2): 271–287, November 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/138.

Algina:2004:CML

[AK04b] James Algina and H. J. Keselman. A comparison of methods for longitudinal analysis with missing data. Journal of Modern Applied Statistical Methods, 3(1):13–26, May 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/112.

Alawadhi:2010:MCA

[AK10] Shafiqah Alawadhi and Mokhtar Konsowa. Markov chain analysis and student academic progress: an empirical comparative study. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/515.

Alzaatreh:2013:GHN

[AK13] Ayman Alzaatreh and Kristen Knight. On the gamma-half normal distribution and its applications. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/646.

Akhtar:2018:JBR

[AK18a] M. Tanwir Akhtar and Athar Ali Khan. JMASM 51: Bayesian reliability analysis of binomial model — application to success/ failure data. Journal of Modern Applied Statistical Methods, 17 (2):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/1009.

Arenson:2018:BBM

[AK18b] Ethan A. Arenson and George Karabatsos. A Bayesian betamixture model for nonparametric IRT (BBM-IRT). Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/972.

Abujarad:2020:JBS

[AK20] Mohammed H. A. Abujarad and Athar Ali Khan. JMASM 57: Bayesian survival analysis of Lomax family models with Stan (R). Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1099.

Anwar:2021:OSI

[AK21] Zaki Anwar and Abdul Nasir Khan. Order statistics from the ishita distribution and associated inference. Journal of Modern Applied Statistical Methods, 20(1):??, February 2021. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1113.

Ahmad:2011:BRA

[AKA11] Sheikh P. Ahmad, A. A. Khan, and A. Ahmed. Bayesian regression analysis with examples in S-PLUS and R. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/541.

Aldahlan:2020:NGF

[AKA20] Maha A. D. Aldahlan, Mohamed G. Khalil, and Ahmed Z. Afify. A new generalized family of distributions for lifetime data. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/1075.

Ashutosh:2024:IGS

[AKK24] Ashutosh Ashutosh, B. B. Khare, and S. Khare. Improved generalized synthetic estimator for domain mean. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1163.

Algina:2003:AGT

[AKO03] James Algina, H. J. Keselman, and Abdul R. Othman. Analyzing group by time effects in longitudinal two-group randomized trial designs with missing data. Journal of Modern Applied Statistical Methods, 2(1):50–72, May 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/66.

Algina:2006:CIE

[AKP06] James Algina, H. J. Keselman, and Randall D. Penfield. Confidence intervals for an effect size when variances are not equal. Journal of Modern Applied Statistical Methods, 5(1):??, ???? 2006. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/222.

Algina:2008:CIS

[AKP08] James Algina, H. J. Keselman, and Randall D. Penfield. Confidence intervals for the squared multiple semipartial correlation coefficient. Journal of Modern Applied Statistical Methods, 7(1): ??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/341.

Al-Kassab:2010:CBU

[AKQ10] Mowafaq M. Al-Kassab and Omar Q. Qwaider. A comparison between unbiased ridge and least squares regression methods using simulation technique. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/505.

Algamal:2015:APB

[AL15] Zakariya Yahya Algamal and Muhammad Hisyam Lee. Applying penalized binary logistic regression with correlation based elastic net for variables selection. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/760.

Almheidat:2016:GWD

[ALF16] Maalee Almheidat, Carl Lee, and Felix Famoye. A generalization of the Weibull distribution with applications. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/876.

Alkahtani:2013:RDV

[Alk13] Saad Saeed Alkahtani. Robustness of DEWMA versus EWMA control charts to non-normal processes. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/649.

Alonso:2004:JFR

[Alo04] Andrés M. Alonso. JMASM 10: a Fortran routine for sieve bootstrap prediction intervals. Journal of Modern Applied Statistical Methods, 3(1):239–249, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/135.

Alonso:2006:JCA

[Alo06] Andrés M. Alonso. JMASM 23: Cluster analysis in epidemiological data (Matlab). Journal of Modern Applied Statistical Methods, 5(1):273–280, May 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/243.

Alyass:2007:GMD

[Aly07] Kussiy K. Alyass. Global measure of the deviation of a wavelet density estimator. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/327.

Al-Masri:2010:CIT

[AM10] Abedel-Qader S. Al-Masri. Combining independent tests of conditional shifted exponential distribution. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/479.

Adewara:2014:SEU

[AM14] Johnson A. Adewara and Ugochukwu A. Mbata. Survival estimation using bootstrap, jackknife and K-repeated jackknife methods. Journal of Modern Applied Statistical Methods, 13(2): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/732.

Aboukhamseen:2017:GAC

[AM17] Suja M. Aboukhamseen and Rym A. M’Hallah. Genetic algorithms for cross-calibration of categorical data. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/921.

Ashour:2019:BEP

[AM19] Samir Kamel Ashour and Mohamed Salem Abdelwahab Muiftah. Bayesian estimation of the parameters of discrete Weibull Type (I) distribution. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1049.

Al-Nasser:2004:EML

[AN04] Amjad D. Al-Nasser. Estimation of multiple linear functional relationships. Journal of Modern Applied Statistical Methods, 3 (1):215–220, May 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/131.

Atem:2012:RMM

[ANA12] Folefac Atem, Julius S. Ngwa, and Abidemi Adeniji. Regression models for mixed over-dispersed Poisson and continuous clustered data: Modeling BMI and number of cigarettes smoked per day. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/601.

Ahmad:2016:JAM

[ANAS16] Wan Muhamad Amir W. Ahmad, Mohamad Arif Awang Nawi, Nor Azlida Aleng, and Mohamad Shafiq. JMASM 41: an alternative method for multiple linear model regression modeling, a technical combining of robust, bootstrap and fuzzy approach (SAS). Journal of Modern Applied Statistical Methods, 15(2): ??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/879.

Anderson:2013:CDBa

[And13a] Richard B. Anderson. Conceptual distinction between the critical p value and the Type I error rate in permutation testing. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/633. See [And13b] and

response [PB13].

Anderson:2013:CDBb

[And13b] Richard B. Anderson. Conceptual distinction between the critical p value and the Type I error rate in permutation testing: Author response to peer comments. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/

article/view/637. See [And13a] and response [PB13].

Ajibade:2015:DIS

[ANM15] Bright F. Ajibade, Chinwe R. Nwosu, and J. I. Mbegdu. The distribution of the inverse square root transformed error component of the multiplicative time series model. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/780.

Ade:2025:HEF

[ANO^{+}25] Ogunde Adebisi Ade, Oluwole Adegoke Nuga, Oseghale Innocent Osezuwa, Adekola Lanrewaju Olumide, and Adebayo Emmanuel. Harris extended Fréchet distribution: Properties, inference, and applications to failure and waiting time data. Journal of Modern Applied Statistical Methods, 24(1):31–53, ???? 2025. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/1339.

Ayinde:2008:RSE

[AO08] Kayode Ayinde and J. O. Olaomi. Robustness of some estimators of linear model with autocorrelated error terms when stochastic regressors are normally distributed. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/359.

Al-Omari:2015:NEE

[AO15] Amer Ibrahim Al-Omari. New entropy estimators with smaller root mean squared error. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/775.

Al-Omari:2011:MLE

[AOAH11] Amer Ibrahim Al-Omari and Said Ali Al-Hadhrami. On maximum likelihood estimators of the parameters of a modified Weibull distribution using extreme ranked set sampling. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/568.

Al-Omari:2016:NEP

[AOBCP16] Amer I. Al-Omari, Carlos N. Bouza, Dante Covarrubias, and Roma Pal. A new estimator of the population mean: an application to bioleaching studies. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/842.

Al-Omari:2019:EDA

[AOCAN19] Amer Ibrahim Al-Omari, Enrico Ciavolino, and Amjad D. Al-Nasser. Economic design of acceptance sampling plans for truncated life tests using three-parameter Lindley distribution. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/1052.

Ahad:2011:TER

[AOY11] Nor Aishah Ahad, Abdul Rahman Othman, and Sharipah Soaad Syed Yahaya. Type I error rates of the two-sample pseudomedian procedure. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/553.

Ahmad:2010:REL

[ARM10] Sanizah Ahmad, Norazan Mohamed Ramli, and Habshah Midi. Robust estimators in logistic regression: a comparative simulation study. Journal of Modern Applied Statistical Methods, 9(2): ??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/507.

Almasri:2005:TCR

[AS05] Abdullah Almasri and Ghazi Shukur. Testing the casual relation between sunspots and temperature using wavelets analysis. Journal of Modern Applied Statistical Methods, 4(1):??, ???? 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/175.

Abu-Shawiesh:2010:ACI

[AS10] Moustafa Omar Ahmed Abu-Shawiesh. Adjusted confidence interval for the population median of the exponential distribution. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/502.

Adhikari:2015:MLT

[AS15a] Achut Adhikari and Jay Schaffer. Modified Lilliefors test. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/754.

Aydin:2015:MCC

[AŞ15b] Demet Aydin and Birdal Şenoğlu. Monte Carlo comparison of the parameter estimation methods for the two-parameter Gumbel distribution. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/777.

Aryadoust:2016:VPB

[AS16] Vahid Aryadoust and Zahra Shahsavar. Validity of the Persian blog attitude questionnaire: an evidence-based approach. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/807.

Abreu:2017:DEC

[AS17] Rafael Perez Abreu and Jay R. Schaffer. A double EWMA control chart for the individuals based on a linear prediction. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/947.

Asar:2016:LTL

[Asa16] Yasin Asar. Liu-type logistic estimators with optimal shrinkage parameter. Journal of Modern Applied Statistical Methods, 15 (1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/822.

Al-Subh:2012:MEG

[ASAIJ12] S. A. Al-Subh, M. T. Alodat, Kamaruzaman Ibrahim, and Abdul Aziz Jemain. Modified EDF goodness of fit tests for logistic distribution under SRS and RSS. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/619.

Arunachalam:2016:CEB

[ASG16a] Rajarathinam Arunachalam, Mahalakshmi Sivasubramanian, and Dilip Kumar Ghosh. Construction of efficiency-balanced design using factorial design. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/798.

Arunachalam:2016:CPW

[ASG16b] Rajarathinam Arunachalam, Mahalakshmi Sivasubramanian, and Dilip Kumar Ghosh. Construction of pair-wise balanced design. Journal of Modern Applied Statistical Methods, 15(1):

??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/796.

Ahsanullah:2016:CCD

[ASK16] M. Ahsanullah, M. Shakil, and B. M. Golam Kibria. Characterizations of continuous distributions by truncated moment. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/802.

Amir:2016:JSR

[ASM^{+}16] Wan Muhamad Amir, Mohamad Shafiq, Kasypi Mokhtar, Nor Azlida Aleng, and Hanafi A. Rahim. JMASM 37: Simple response surface methodology using RSREG (SAS). Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/832.

Amir:2016:JAC

[ASR^{+}16] Wan Muhamad Amir, Mohamad Shafiq, Hanafi A. Rahim, Puspa Liza, and Azlida Aleng. JMASM 39: Algorithm for combining robust and bootstrap in multiple linear model regression (SAS). Journal of Modern Applied Statistical Methods, 15(1): ??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/834.

Al-Saleh:2004:EUB

[ASS04] Mohammad Fraiwan Al-Saleh and Hani M. Samawi. Estimation using bivariate extreme ranked set sampling with application to the bivariate normal distribution. Journal of Modern Applied Statistical Methods, 3(1):134–142, May 2004. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/122.

Abu-Shawiesh:2019:RCI

[ASS19] Moustafa Omar Ahmed Abu-Shawiesh and Aamir Saghir. Robust confidence intervals for the population mean alternatives to the Student-t confidence interval. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1025.

Aryal:2009:ATS

[AT09] Gokarna R. Aryal and Chris P. Tsokos. Application of the truncated skew Laplace probability distribution in maintenance system. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/433.

Abebe:2017:JCP

[ATvBB17] Haftom Temesgen Abebe, Frans E. S. Tan, Gerard J. P. van Breukelen, and Martijn P. F. Berger. JMASM 45: a computer program for Bayesian D-optimal binary repeated measurements designs (Matlab). Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/920.

Ali:2005:IPD

[AW05] M. Masoom Ali and Jungsoo Woo. Inference on (Y < X) in a Pareto distribution. Journal of Modern Applied Statistical Methods, 4(2):583–586, November 2005. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/215.

Arumairajan:2016:PCP

[AW16] Sivarajah Arumairajan and Pushpakanthie Wijekoon. Principal component preliminary test estimator in the linear regression model. Journal of Modern Applied Statistical Methods, 15(1): ??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/819.

Alkhalaf:2017:IPV

[AZ17] Arwa Alkhalaf and Bruno D. Zumbo. The impact of predictor variable(s) with skewed cell probabilities on Wald tests in binary logistic regression. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/927.

Amusa:2020:SSS

[AZN20] Lateef Babatunde Amusa, Temesgen Zewotir, and Delia North. A simulation study of some modern weighting methods for estimating treatment effects in observational studies. Journal of Modern Applied Statistical Methods, 19(2):??, July 2020. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1108.

Bridget:2011:SMC

[BA11] Ahani E. Bridget and O. Abass. A sequential Monte Carlo approach for online stock market prediction using hidden Markov models. Journal of Modern Applied Statistical Methods, 10(2): ??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/575.

Baklizi:2003:CIE

[Bak03] Ayman Baklizi. Confidence intervals for P (X < Y) in the exponential case with common location parameter. Journal of Modern Applied Statistical Methods, 2(2):341–349, November 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/91.

Baklizi:2007:SMF

[Bak07] Ayman Baklizi. A simple method for finding empirical likelihood type intervals for the ROC curve. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/331.

Baklizi:2006:LSB

[BAN06] Ayman Baklizi and Amjad Al-Nasser. Large sample and bootstrap intervals for the gamma scale parameter based on grouped data. Journal of Modern Applied Statistical Methods, 5(2):356– 366, November 2006. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/253.

Barron:2015:CCI

[Bar15] Dean S. Barron. Contrails: Causal inference using propensity scores. Journal of Modern Applied Statistical Methods, 14(2): ??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/783.

Bagui:2004:ACC

[BB04] Sikha Bagui and Subhash Bagui. An algorithm and code for computing exact critical values for the Kruskal–Wallis nonparametric one-way ANOVA. Journal of Modern Applied Statistical Methods, 3(2):498–503, November 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/160.

Bagui:2005:JAC

[BB05] Sikha Bagui and Sbuhash Bagui. JMASM 17: an algorithm and code for computing exact critical values for Friedman’s nonparametric ANOVA. Journal of Modern Applied Statistical Methods, 4(1):312–318, May 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/191.

Bagui:2006:JCP

[BB06] Sikha Bagui and Subhash Bagui. JMASM 25: Computing percentiles of skew-normal distributions. Journal of Modern Applied Statistical Methods, 5(2):575–588, November 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/274.

Balloun:2007:ILK

[BB07] Joseph L. Balloun and Hilton Barret. From information lost to knowledge gained: the benefits of analyzing all the research evidence. Journal of Modern Applied Statistical Methods, 6(2): ??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/326.

Bhuyan:2011:LMR

[BB11] Abhijit Bhuyan and Munindra Borah. LQ-moments for regional flood frequency analysis: a case study for the North-Bank Region of the Brahmaputra River, India. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/582.

Baird:2016:GDI

[BB16] Grayson L. Baird and Stephen L. Bieber. The Goldilocks Dilemma: Impacts of multicollinearity — a comparison of simple linear regression, multiple regression, and ordered variable regression models. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/803.

Baird:2019:SPC

[BB19] Grayson L. Baird and Stephen L. Bieber. Sampling the porridge: a comparison of ordered variable regression with F and R2 and multiple linear regression with corrected F and R2 in the presence of multicollinearity. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1023.

Begum:2012:GMH

[BBB12] Munni Begum, Jay Bagga, and C. Ann Blakey. Graphical modeling for high dimensional data. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/625.

Balloun:2011:OEN

[BBW11] Joseph L. Balloun, Hilton Barret, and Art Weinstein. One is not enough: the need for multiple respondents in survey research of organizations. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/543.

Blair:2002:TSE

[BC02] R. Clifford Blair and Stephen R. Cole. Two-sided equivalence testing of the difference between two means. Journal of Modern Applied Statistical Methods, 1(1):??, ???? 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/18.

Banicescu:2005:SPP

[BC05] Ioana Banicescu and Ricolindo L. Cariño. Simulation procedure in periodic cancer screening trials. Journal of Modern Applied Statistical Methods, 4(2):522–527, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/208.

Bhoj:2019:SUA

[BC19] Dinesh Bhoj and Girish Chandra. Simple unequal allocation procedure for ranked set sampling with skew distributions. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1055.

Baird:2016:RCR

[BD16] Grayson L. Baird and Sunny R. Duerr. Reflections concerning recent ban on NHST and confidence intervals. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/881.

Bal:2020:JMA

[BD20] Cagatay Bal and Serdar Demir. JMASM 55: MATLAB algorithms and source codes of ’cbnet’ function for univariate time series modeling with neural networks (MATLAB). Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1100.

Baklizi:2006:CSS

[BE06] Ayman Baklizi and Mohammed Al-Haj Ebrahem. Comparison of some simple estimators of the lognormal parameters based on censored samples. Journal of Modern Applied Statistical Methods, 5(1):??, ???? 2006. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/235.

Brown:2017:PSS

[BE17] Paul M. Brown and Justin A. Ezekowitz. Power and sample size estimation for nonparametric composite endpoints: Practical implementation using data simulations. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/935.

Beauducel:2011:IFS

[Bea11] André Beauducel. Indeterminacy of factor score estimates in slightly misspecified confirmatory factor models. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/566.

Beauducel:2017:SLB

[Bea17] André Beauducel. A Schmid–Leiman-based transformation resulting in perfect inter-correlations of three types of factor score predictors. Journal of Modern Applied Statistical Methods, 16 (1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/888.

Bendermacher:2010:BAL

[Ben10] Nol Bendermacher. Beyond alpha: Lower bounds for the reliability of tests. Journal of Modern Applied Statistical Methods, 9 (1):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/469.

Bendermacher:2017:UEG

[Ben17] Nol Bendermacher. An unbiased estimator of the greatest lower bound. Journal of Modern Applied Statistical Methods, 16(1): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/919.

Berger:2005:TSA

[Ber05] Vance W. Berger. Training statisticians to be alert to the dangers of misapplying statistical methods. Journal of Modern Applied Statistical Methods, 4(2):587–590, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/216.

Berger:2009:SD

[Ber09] Vance Berger. A Socratic dialogue. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/424.

Berger:2017:EDN

[Ber17] Vance W. Berger. An empirical demonstration of the need for exact tests. Journal of Modern Applied Statistical Methods, 16 (1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/884.

Bhandary:2009:CIE

[BF09] Madhusudan Bhandary and Koji Fujiwara. Confidence interval estimation for intraclass correlation coefficient under unequal family sizes. Journal of Modern Applied Statistical Methods, 8 (2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/442.

Bhandary:2011:ETE

[BF11] Madhusudan Bhandary and Koji Fujiwara. An exact test for the equality of intraclass correlation coefficients under unequal family sizes. Journal of Modern Applied Statistical Methods, 10 (1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/536.

Bhandary:2013:TIC

[BF13] Madhusudan Bhandary and Koji Fujiwara. Test for intraclass correlation coefficient under unequal family sizes. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/668.

Bhandary:2014:ATE

[BF14] Madhusudan Bhandary and Koji Fujiwara. An alternative test for the equality of intraclass correlation coefficients under unequal family sizes for several populations. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/690.

Bhandary:2015:TEP

[BG15] Madhusudan Bhandary and Arjun K. Gupta. Test for the equality of partial correlation coefficients for two populations. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/755.

Boik:2006:SOA

[BH06] Robert J. Boik and Ben Haaland. Second-order accurate inference on simple, partial, and multiple correlations. Journal of Modern Applied Statistical Methods, 5(2):283–308, November 2006. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/245.

Baird:2016:DOS

[BH16] Grayson L. Baird and Lisa L. Harlow. Does one size fit all? A case for context-driven null hypothesis statistical testing. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/792.

Bhandary:2008:ECM

[Bha08] Madhusudan Bhandary. Estimation of covariance matrix in signal processing when the noise covariance matrix is arbitrary. Journal of Modern Applied Statistical Methods, 7(1):??, ????

2008. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/355.

Bhandary:2010:ECI

[Bha10] Madhusudan Bhandary. On exact 100(1 − α)% confidence interval of autocorrelation coefficient in multivariate data when the errors are autocorrelated. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/468.

Bhattacharjee:2014:DCC

[Bha14] Atanu Bhattacharjee. Distance correlation coefficient: an application with Bayesian approach in clinical data analysis. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/709.

Brown:2012:WDI

[BHH12] Jennifer L. Brown, Gerald Halpin, and Glennelle Halpin. Weight: Does it really matter? Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/600.

Berger:2002:ATO

[BI02] Vance W. Berger and Anastasia Ivanova. Adaptive tests for ordered categorical data. Journal of Modern Applied Statistical Methods, 1(2):269–280, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/33.

Bichutskiy:2011:PTS

[Bic11] Vadim Y. Bichutskiy. A pooled two-sample median test based on density estimation. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/578.

Biswas:2011:HOM

[BII11] Soma Chowdhury Biswas, M. Ataharul Islam, and Jamal Nazrul Islam. Higher order Markov structure-based logistic model and likelihood inference for ordinal data. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/562.

Brooks:2011:SSC

[BJ11] Gordon P. Brooks and George A. Johanson. Sample size considerations for multiple comparison procedures in ANOVA. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/526.

Bashir:2022:SSR

[BJJ22] Nafeesa Bashir, J. P. S. Joorel, and T. R. Jan. Stress strength reliability in multicomponent model for Rayleigh-exponential (log logistic) distribution. Journal of Modern Applied Statistical Methods, 21(2):2–16, July 2022. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1127.

Bashir:2023:SSR

[BJJ23] Nafeesa Bashir, J. P. S. Joorel, and T. R. Jan. Stress strength reliability in multicomponent model for Rayleigh-exponential (log logistic) distribution. Journal of Modern Applied Statistical Methods, 22(2):??, July 2023. CODEN ???? ISSN 1538-9472.

Banik:2009:SDD

[BK09a] Shipra Banik and B. M. Golam Kibria. On some discrete distributions and their applications with real life data. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/434.

Bhandary:2009:TEN

[BK09b] Madhusudan Bhandary and Debasis Kundu. Test for the equality of the number of signals. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/447.

Biradar:2022:BES

[BK22] B. S. Biradar and Shivanna B. K. Bayesian estimation of the scale parameter of Weibull distribution using ranked set samples. Journal of Modern Applied Statistical Methods, 21(2):??, July 2022. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1215.

Banik:2012:TPC

[BKS12] Shipra Banik, B. M. Golam Kibria, and Dinesh Sharma. Testing the population coefficient of variation. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/613.

Blood:2013:EHI

[BKS13] Emily A. Blood, Leslie A. Kalish, and Lydia A. Shrier. Estimating heterogeneous intra-class correlation coefficients in dyadic ecological momentary assessment. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/654.

Blair:2002:CTN

[Bla02] R. Clifford Blair. Combining two nonparametric tests of location. Journal of Modern Applied Statistical Methods, 1(1):13–18, May 2002. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/3.

Berger:2002:PAR

[BLEL02] Vance W. Berger, Clifford E. Lunneborg, Michael D. Ernst, and Jonathan G. Levine. Parametric analyses in randomized clinical trials. Journal of Modern Applied Statistical Methods, 1(1):74– 82, May 2002. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/11.

Brobbey:2022:RTL

[BLNA^{+}22] Anita Brobbey, Lisa M. Lix, Alberto Nettel-Aguirre, Tyler Williamson, Samuel Wiebe, and Tolulope Sajobi. Robust trimmed likelihood discriminant analysis for multivariate repeated data. Journal of Modern Applied Statistical Methods, 21(1):??, February 2022. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1118.

Barnette:2005:TEF

[BM05] J. Jackson Barnette and James E. McLean. Type I error of four pairwise mean comparison procedures conducted as protected and unprotected tests. Journal of Modern Applied Statistical Methods, 4(2):446–459, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/201.

Bulte:2013:SCD

[BO13] Isis Bulté and Patrick Onghena. The single-case data analysis package: Analysing single-case experiments with R software. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/687.

Bashiru:2010:NTT

[BOO10] K. A. Bashiru, O. E. Olowofeso, and S. A. Owabumoye. Nonlinear trigonometric transformation time series modeling. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/503.

Booker:2019:RLM

[Boo19] Zsa-Zsa Booker. A revised logic model for educational program evaluation. Journal of Modern Applied Statistical Methods, 18 (2):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/1062.

Bose:2009:ECD

[Bos09] Chinmoy K. Bose. Efficiency of canonical discriminant function versus Mahalanobis distance in differentiating groups: Screening ovarian cancer in a multivariate system analysis using enzyme markers. Journal of Modern Applied Statistical Methods, 8(1): ??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/423.

Bouza:2010:RSS

[Bou10] Carlos N. Bouza. Ranked set sampling using auxiliary variables of a randomized response procedure for estimating the mean of a sensitive quantitative character. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/482.

Barot:2016:HBE

[BP16] D. R. Barot and M. N. Patel. Hierarchical Bayes estimation of reliability indexes of cold standby series system under general progressive Type II censoring scheme. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/872.

Bai:2010:ALR

[BPWR10] Haiyan Bai, Wei Pan, Leigh Lihshing Wang, and Phillip Neal Ritchey. Another look at resampling: Replenishing small samples with virtual data through S-SMART. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/476.

Bekrizadeh:2012:WCJ

[BPZ12] Hakim Bekrizadeh, Gholam Ali Parham, and Mohamd Reza Zadkarmi. Weighted Cook–Johnson copula and their characterizations: Application to probably modeling of the hot spring eruptions. Journal of Modern Applied Statistical Methods, 11(2): ??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/630.

Bond:2003:ISD

[BR03] Charles F. Bond and F. D. Richard. P ∗ index of segregation: Distribution under reassignment. Journal of Modern Applied Statistical Methods, 2(2):306–313, November 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/88.

Best:2014:CTA

[BR14] Donald J. Best and John Charles Rayner. Conover’s F test as an alternative to Durbin’s test. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/721.

Brainerd:2003:IFC

[Bra03] Walt Brainerd. The importance of Fortran in the 21st Century. Journal of Modern Applied Statistical Methods, 2(1):14–15, May 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/58.

Bai:2017:RAR

[BRT17] Z. D. Bai, C. R. Rao, and M. T. Tsai. A remark for the admissibility of Rao’s U-test. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/949.

Bunner:2002:ABI

[BS02] Jennifer Bunner and Shlomo S. Sawilowsky. Alternatives to S_{w} in the bracketed interval of the trimmed mean. Journal of Modern Applied Statistical Methods, 1(1):176–181, May 2002. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/21.

Barcelo:2004:MEA

[BS04] Maria Antònia Barceló and Marc Saez. A modification of the EM algorithm to estimate an Andersen–Gill gamma frailty model for multivariate failure time data. Journal of Modern Applied Statistical Methods, 3(2):504–517, November 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/161.

Bendermacher:2009:BKE

[BS09] Nol Bendermacher and Pierre Souren. Beyond kappa: Estimating inter-rater agreement with nominal classifications. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/404.

Beversdorf:2011:TCB

[BS11a] L. Beversdorf and Ping Sa. Tests for correlation on bivariate nonnormal data. Journal of Modern Applied Statistical Methods, 10 (2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/579.

Borysov:2011:ROS

[BS11b] P. Borysov and Ping Sa. A robust one-sided variability control chart. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/551.

Beauducel:2014:RCF

[BS14] André Beauducel and Frank Spohn. Retained-components factor transformation: Factor loadings and factor score predictors in the column space of retained components. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/723.

Broemeling:2005:PFB

[BW05] Lyle Broemeling and Dongfeng Wu. On the power function of Bayesian tests with application to design of clinical trials: the fixed-sample case. Journal of Modern Applied Statistical Methods, 4(1):163–171, May 2005. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/178.

Baklizi:2011:ISR

[BY11] Ayman Baklizi and Adil E. Yousif. Inference in simple regression for the intercept utilizing prior information on the slope. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/533.

Beasley:2009:ART

[BZ09] T. Mark Beasley and Bruno D. Zumbo. Aligned rank tests for interactions in split-plot designs: Distributional assumptions and stochastic heterogeneity. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/398.

Bhandary:2011:CST

[BZ11] Madhusudan Bhandary and Xuan Zhang. Comparison of several tests for combining several independent tests. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/555.

Bartolucci:2003:RRM

[BZBS03] Alfred A. Bartolucci, Shimin Zheng, Sejong Bae, and Karan P. Singh. Random regression models based on the elliptically contoured distribution assumptions with applications to longitudinal data. Journal of Modern Applied Statistical Methods, 2(2): 359–370, November 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/93.

bintiZafakali:2013:MHO

[bZbWA13] Nur Syabiha binti Zafakali and Wan Muhamad Amir bin W. Ahmad. Modeling and handling overdispersion health science data with zero-inflated Poisson model. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/659.

Cetin:2008:CIB

[CA08] Meral Cetin and Serpil Aktas. Confidence intervals based on robust estimators. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/360.

Camara:2003:ABC

[Cam03] Vincent A. R. Camara. Approximate Bayesian confidence intervals for the variance of a Gaussian distribution. Journal of Modern Applied Statistical Methods, 2(2):350–358, November 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/92.

Camara:2007:ABC

[Cam07] Vincent A. R. Camara. Approximate Bayesian confidence intervals for the mean of an exponential distribution versus Fisher matrix bounds models. Journal of Modern Applied Statistical Methods, 6(1):141–152, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/289.

Crimin:2008:RGL

[CAM08] Kimberly Crimin, Asheber Abebe, and Joseph W. McKean. Robust general linear models and graphics via a user interface (web RGLM). Journal of Modern Applied Statistical Methods, 7(1): ??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/367.

Camara:2009:ABC

[Cam09a] Vincent A. R. Camara. Approximate Bayesian confidence intervals for the mean of a Gaussian distribution versus Bayesian models. Journal of Modern Applied Statistical Methods, 8(2): ??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/443.

Camara:2009:NAB

[Cam09b] Vincent A. R. Camara. A new approximate Bayesian approach for decision making about the variance of a Gaussian distribution versus the classical approach. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/416.

Camara:2012:NAB

[Cam12] Vincent A. R. Camara. New approximate Bayesian confidence intervals for the coefficient of variation of a Gaussian distribution. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/595.

Cirillo:2012:RRE

[CB12] Marcelo Angelo Cirillo and Lúcia Pereira Barroso. Robust regression estimates in the prediction of latent variables in structural equation models. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/586.

Chandra:2018:SUR

[CBP18] Girish Chandra, Dinesh S. Bhoj, and Rajiv Pandey. Simple unbalanced ranked set sampling for mean estimation of response variable of developmental programs. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/980.

Chen:2009:CML

[CC09] Jinsong Chen and Jaehwa Choi. A comparison of maximum likelihood and expected a posteriori estimation for polychoric correlation using Monte Carlo simulation. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/427.

Choi:2019:CSI

[CC19] Youn-Jeng Choi and Allan S. Cohen. Comparison of scale identification methods in mixture IRT models. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1038.

Counsell:2019:CMU

[CCC19] Alyssa Counsell, Robert Philip Chalmers, and Robert A. Cribbie. Comparing means under heteroscedasticity and nonnormality: Further exploring robust means modeling. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1037.

Choi:2019:LGM

[CCH19] Jaehwa Choi, Jinsong Chen, and Jeffrey R. Harring. Logistic growth modeling with Markov chain Monte Carlo estimation. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1032.

Calik:2009:EVD

[CCT09] Sinan Calik, Cemil Colak, and Ayse Turan. On the expected values of distribution of the sample range of order statistics from the geometric distribution. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/410.

Chowhan:2009:GCA

[CD09] James Chowhan and Laura Duncan. Generating and comparing aggregate variables for use across datasets in multilevel analysis. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/455.

Carrion:2022:EVR

[CdlTE22] Rafael Perez Abreu Carrion, Hector de la Torre, and Samantha Estrada. An extreme value regression approach to the prediction of shrimp size distribution. Journal of Modern Applied Statistical Methods, 21(1):??, February 2022. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1120.

Chatzipantsiou:2019:JEE

[CDPT19] Christina Chatzipantsiou, Marios Dimitriadis, Manos Papadakis, and Michail Tsagris. JMASM 52: Extremely efficient permutation and bootstrap hypothesis tests using R. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1069.

Cilingturk:2014:HCS

[ÇE14] A. Mete Çilingtürk and Özlem Ergüt. Hierarchical clustering with simple matching and joint entropy dissimilarity measure. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/707.

Cangur:2015:CMF

[CE15] Sengul Cangur and Ilker Ercan. Comparison of model fit indices used in structural equation modeling under multivariate normality. Journal of Modern Applied Statistical Methods, 14 (1):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/759.

Clement:2024:MCS

[CEI24] Etebong P. Clement, Emmanuel J. Ekpenyong, and Matthew J. Iseh. Modified classes of stratified exponential ratio estimators of population mean with equal optimal efficiency. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1170.

Cengiz:2004:ABD

[Cen04] Mehmet Ali Cengiz. Accurate binary decisions for assessing coronary artery disease. Journal of Modern Applied Statistical Methods, 3(1):158–164, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/125.

Clement:2024:ERE

[CERA24] Etebong P. Clement, Idorenyin O. Etukudoh, Khalid Ul Islam Rather, and Victoria M. Akpan. On efficient ratio estimation of population mean by multivariate calibration weightings in double sampling for stratification. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1169.

Chiu:2003:ISW

[CF03] Christopher W. T. Chiu and Ronald S. Fesco. Incorporating sampling weights into the generalizability theory for large-scale analyses. Journal of Modern Applied Statistical Methods, 2(1): 108–127, May 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/70.

Cirillo:2010:GVR

[CFSF10] Marcelo Angelo Cirillo, Daniel Furtado Ferreira, Thelma Sáfadi, and Eric Batista Ferreira. Generalized variances ratio test for comparing k covariance matrices from dependent normal populations. Journal of Modern Applied Statistical Methods, 9(2): ??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/495.

Chowdhury:2015:CSP

[CGH15] Farzana Chowdhury, Jahida Gulshan, and Syed Shahadat Hossain. A comparison of semi-parametric and nonparametric methods for estimating mean time to event for randomly left censored data. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/762.

Chakrabartty:2022:ETS

[Cha22] Satyendra Nath Chakrabartty. Estimation of true scores, true score variance, reliability and tests of parallelism. Journal of Modern Applied Statistical Methods, 21(2):??, July 2022. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1129.

Chakrabartty:2024:RAQ

[Cha24] Satyendra Nath Chakrabartty. Relationships among quality parameters of tests. Journal of Modern Applied Statistical Methods, 23(1):??, November 2024. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1180.

Cheema:2014:SGG

[Che14] Jehanzeb R. Cheema. Some general guidelines for choosing missing data handling methods in educational research. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/720.

Coman:2014:SPA

[CID^{+}14] Emil N. Coman, Eugen Iordache, Lisa Dierker, Judith Fifield, and Jean J. Schensul. Statistical power of alternative structural models for comparative effectiveness research: Advantages of modeling unreliability. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/692.

Choi:2018:IED

[CKCL18] Youn-Jeng Choi, Seohyun Kim, Allan S. Cohen, and Zhenqiu Lu. The impact of equating on detection of treatment effects. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1004.

Chung:2012:JST

[CKP12] Hyewon Chung, Jiseon Kim, and Ryoungsun Park. JMASM 32: SAS template for single-subject experimental designs. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/609.

Chung:2018:ISS

[CKPJ18] Hyewon Chung, Jiseon Kim, Ryoungsun Park, and Hyeonjeong Jean. The impact of sample size in cross-classified multiple membership multilevel models. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/978.

Clif:2011:TCF

[Cli11] Norman Clif. A test that combines frequency and quantitative information. Journal of Modern Applied Statistical Methods, 10 (1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/518.

Corcoran:2002:ELP

[CM02] Christopher D. Corcoran and Cyrus R. Mehta. Exact level and power of permutation, bootstrap, and asymptotic tests of trend. Journal of Modern Applied Statistical Methods, 1(1):32–41, May 2002. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/7.

Chiodo:2002:ESL

[CMR02] Ernest P. Chiodo, Joseph L. Musial, and J. Sia Robinson. An error in statistical logic in the application of genetic paternity testing. Journal of Modern Applied Statistical Methods, 1(1):??, ???? 2002. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/16.

Conover:2002:SLM

[Con02] W. J. Conover. Some locally most powerful rank tests for correlation. Journal of Modern Applied Statistical Methods, 1(1): 19–23, May 2002. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/4.

Consoli:2018:LER

[Con18] Maurizio Consoli. Letter to the editor: Regarding a possible nonnull interpretation of the Michelson–Morley experiment. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/986.

Cox:2013:AAE

[Cox13] M. A. A. Cox. An approximate approach to the economic design of X̄ charts by considering the cost of quality. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/651.

Chen:2013:CCI

[CP13] Li-Ting Chen and Chao-Ying Joanne Peng. Constructing confidence intervals for effect sizes in ANOVA designs. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/664.

Cavicchioli:2018:RFA

[CPDP18] Maddalena Cavicchioli, Angeliki Papana, Ariadni Papana Dagiasis, and Barbara Pistoresi. A random forests approach to assess determinants of central bank independence. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/997.

Chu:2004:CEA

[CPT04] LiPing Chu, S. James Press, and Judith M. Tanur. Confidence elicitation and anchoring in the respondent-generated intervals (RGI) protocol. Journal of Modern Applied Statistical Methods, 3(2):417–431, November 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/154.

Corain:2004:MMN

[CS04] Livio Corain and Luigi Salmaso. Multivariate and multistrata nonparametric tests: the NonParametric combination method. Journal of Modern Applied Statistical Methods, 3(2):443–461, November 2004. CODEN ???? ISSN 1538-9472. URL https:/

/jmasm.com/index.php/jmasm/article/view/156.

Cuzzocrea:2006:PPI

[CS06] John L. Cuzzocrea and Shlomo S. Sawilowsky. Pietro Paoli, Italian algebraist. Journal of Modern Applied Statistical Methods, 5(2):593–595, November 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/276.

Chen:2008:TST

[CS08] Haiyan Chen and Howard H. Stratton. Test for spatio-temporal counts being Poisson. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/361.

Cuzzocrea:2009:RNI

[CS09] John L. Cuzzocrea and Shlomo Sawilowsky. Robustness to nonindependence and power of the I test for trend in construct validity. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/413.

Chen:2012:EMP

[CS12] Cheng C. Chen and Sarjinder Singh. Estimation of multinomial proportions using higher order moments of scrambling variables in randomized response sampling. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/591.

Camara:2005:BRM

[CT05] Vincent A. R. Camara and Chris P. Tsokos. Bayesian reliability modeling using Monte Carlo integration. Journal of Modern Applied Statistical Methods, 4(1):172–186, May 2005. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/180.

Cong:2009:MMB

[CT09] Chunling Cong and Chris P. Tsokos. Markov modeling of breast cancer. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/456.

Chen:2005:MIM

[CTDVD05] Ling Chen, Mariana Toma-Drane, Robert F. Valois, and J. Wanzer Drane. Multiple imputation for missing ordinal data. Journal of Modern Applied Statistical Methods, 4(1):??, ???? 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/188.

Cribbie:2007:TTG

[CWBK07] Robert A. Cribbie, Rand R. Wilcox, Carmen Bewell, and H. J. Keselman. Tests for treatment group equality when data are nonnormal and heteroscedastic. Journal of Modern Applied Statistical Methods, 6(1):117–132, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/287.

Dyer:2002:SSI

[DAC02] John N. Dyer, B. Michael Adams, and Michael D. Conerly. A simulation study of the impact of forecast recovery for control charts applied to ARMA processes. Journal of Modern Applied Statistical Methods, 1(2):343–353, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/40.

Dansawad:2025:EUM

[Dan25] Napattchan Dansawad. Efficient utilization of multiple auxiliary variables for nonresponse problem in estimating the population mean under sub-sampling technique. Journal of Modern Applied Statistical Methods, 24(1):15–22, ???? 2025. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1332.

Dayton:2003:MCU

[Day03] C. Mitchell Dayton. Model comparisons using information measures. Journal of Modern Applied Statistical Methods, 2(2):281– 292, November 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/86.

Dayton:2017:REM

[Day17] Chauncey M. Dayton. A reinterpretation and extension of Mc-Nemar’s test. Journal of Modern Applied Statistical Methods, 16 (1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/883.

Dubnicka:2002:RBP

[DBH02] Suzanne R. Dubnicka, R. Clifford Blair, and Thomas P. Hettmansperger. Rank-based procedures for mixed paired and two- sample designs. Journal of Modern Applied Statistical Methods, 1(1):??, ???? 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/6.

Dutta:2007:OCC

[DDM07] Ganesh Dutta, Premadhis Das, and Nripes Kumar Mandal. Optimum choice of covariates for a series of SBIBDS obtained through projective geometry. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/337.

Dakhn:2017:SPM

[DEE17] Laila Naji Ba Dakhn, Mohammed Al-Haj Ebrahem, and Omar Eidous. Semi-parametric method to estimate the time-to-failure distribution and its percentiles for simple linear degradation model. Journal of Modern Applied Statistical Methods, 16(2): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/940.

Deichmann:2006:UEC

[DEH^{+}06] Joel I. Deichmann, Abdolreza Eshghi, Dominique Haughton, Selin Sayek, Nicholas Teebagy, and Heikki Topi. Understanding Eurasian convergence: Application of Kohonen self-organizing maps. Journal of Modern Applied Statistical Methods, 5(1):73– 94, May 2006. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/228.

Demirtas:2004:MIL

[Dem04a] Hakan Demirtas. Modeling incomplete longitudinal data. Journal of Modern Applied Statistical Methods, 3(2):305–321, November 2004. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/140.

Demirtas:2004:PRN

[Dem04b] Hakan Demirtas. Pseudo-random number generation in R for commonly used multivariate distributions. Journal of Modern Applied Statistical Methods, 3(2):485–497, November 2004. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/159.

Demirtas:2005:JPR

[Dem05] Hakan Demirtas. JMASM 16: Pseudo-random number generation in R for some univariate distributions. Journal of Modern Applied Statistical Methods, 4(1):300–311, May 2005. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/190.

Devika:2016:AET

[DGJ16] Shanmugasundaram Devika, Sebastian George, and Lakshmanan Jeyaseelan. Application of Esscher transformed Laplace distribution in microarray gene expression data. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/816.

Dielman:2006:CSP

[Die06] Terry Dielman. Choosing smoothing parameters for exponential smoothing: Minimizing sums of squared versus sums of absolute errors. Journal of Modern Applied Statistical Methods, 5(1):??, ???? 2006. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/231.

Dielman:2009:ENH

[Die09a] Terry E. Dielman. Email: a note on hypothesis tests after correction for autocorrelation: Solace for the Cochrane–Orcutt method? Journal of Modern Applied Statistical Methods, 8(1): ??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/403.

Dielman:2009:LAV

[Die09b] Terry E. Dielman. Least absolute value vs. least squares estimation and inference procedures in regression models with asymmetric error distributions. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/407.

Dielman:2011:EHT

[Die11] Terry E. Dielman. Estimation and hypothesis testing in LAV regression with autocorrelated errors: Is correction for autocorrelation helpful? Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/563.

DiSario:2002:PGA

[DiS02] Robert DiSario. A program for generating all permutations of 1, 2, …, n. Journal of Modern Applied Statistical Methods, 1(2): 518–522, November 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/56.

Desai:2019:SVM

[DK19] Shivaji Shripati Desai and D. N. Kashid. Support vector machine-based modified sp statistic for subset selection with non-normal error terms. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1033.

Desai:2016:MIW

[DMBR16] Manisha Desai, Aya A. Mitani, Susan W. Bryson, and Thomas Robinson. Multiple imputation when rate of change is the outcome of interest. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/795.

Dewi:2022:INN

[DNUH22] Wardhani Utami Dewi, Khoirin Nisa, Mustofa Usman, and Netti Herawati. Implementation of K-nearest neighbors algorithm on regional food security classification in Indonesia. Journal of Modern Applied Statistical Methods, 21(2):??, July 2022. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1196.

Dorugade:2016:IRE

[Dor16] Ashok Vithoba Dorugade. Improved ridge estimator in linear regression with multicollinearity, heteroscedastic errors and outliers. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/856.

Doyle:2009:ECP

[Doy09] Suzanne R. Doyle. Examples of computing power for zeroinflated and overdispersed count data. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/429.

Dayton:2005:JPB

[DP05] C. Mitchell Dayton and Xuemei Pan. JMASM 21: PCIC SAS: Best subsets using information criteria. Journal of Modern Applied Statistical Methods, 4(2):621–626, November 2005. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/221.

DelosReyes:2021:ECC

[DP21] John Mart V. DelosReyes and Miguel A. Padilla. Estimation of correlation confidence intervals via the bootstrap: Non normal distributions. Journal of Modern Applied Statistical Methods, 20(2):??, August 2021. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1189.

DCunha:2014:BIV

[DR14] Juliet G. D’Cunha and K. A. Rao. Bayesian inference for volatility of stock prices. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/746.

Danish:2018:OSB

[DR18] Faizan Danish and S. E. H. Rizvi. Optimum stratification in bivariate auxiliary variables under Neyman allocation. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/967.

Derrick:2017:TSC

[DRTW17] Ben Derrick, Bethan Russ, Deirdre Toher, and Paul White. Test statistics for the comparison of means for two samples that include both paired and independent observations. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/890.

DAgostino:2002:CDD

[DS02] Ralph D’Agostino, Sr. and Lisa M. Sullivan. Chronic disease data and analysis: Current state of the field. Journal of Modern Applied Statistical Methods, 1(2):228–239, November 2002. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/29.

Duzan:2016:SMP

[DS16] Hanan Duzan and Nurul Sima Binti Mohamaed Sharif. Solution to the multicollinearity problem by adding some constant to the diagonal. Journal of Modern Applied Statistical Methods, 15(1): ??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/823.

Devika:2025:MAT

[DS25] M. Devika and R. Sivaraman. Modulo arithmetic triangle of numbers. Journal of Modern Applied Statistical Methods, 24(1): ??, ???? 2025. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1399.

Drucker:2008:JSH

[DSCS08] Daniel S. Drucker, Claude Schochet, John Cuzzocrea, and Shlomo Sawilowsky. Jacques Salomon Hadamard and the use of symbols in teaching differential calculus. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/370.

Daraz:2018:EFP

[DSK18] Umer Daraz, Javid Shabbir, and Hina Khan. Estimation of finite population mean by using minimum and maximum values in stratified random sampling. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/973.

Drane:2004:TRT

[DT04] J. Wanzer Drane and W. Gregory Thatcher. A test–retest transition matrix: a modification of McNemar’s test. Journal of Modern Applied Statistical Methods, 3(1):213–214, May 2004. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/130.

Dubnicka:2004:RBE

[Dub04] Suzanne R. Dubnicka. A rank-based estimation procedure for linear models with clustered data. Journal of Modern Applied Statistical Methods, 3(1):39–48, May 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/114.

Dardick:2019:ICS

[DW19] William R. Dardick and Brandi A. Weiss. An investigation of chi-square and entropy based methods of item-fit using item level contamination in item response theory. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1059.

Derrick:2019:PNP

[DWT19] Ben Derrick, Paul White, and Deirdre Toher. Parametric and non-parametric tests for the comparison of two samples which both include paired and unpaired observations. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1040.

Dikheel:2020:RLW

[DY20] Tahir R. Dikheel and Alaa Q. Yaseen. Robust lag weighted lasso for time series model. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1081.

Dyer:2011:SSR

[Dye11] John N. Dyer. A simulation study of the relative efficiency of the minimized integrated square error estimator (L2E) for phase I control charting. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/544.

Dyer:2016:MCS

[Dye16] John N. Dyer. Monte Carlo simulation design for evaluating normal-based control chart properties. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/868.

Ebrahem:2009:EPR

[EAM09] Mohammed Al-Haj Ebrahem and Abedel-Qader Al-Masri. Estimating the parameters of Rayleigh cumulative exposure model in simple step-stress testing. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/438.

Ebrahem:2005:ESS

[EAN05] Mohammed Al-Haj Ebrahem and Amjad D. Al-Nasser. Estimating the slope of simple linear regression in the presence of outliers. Journal of Modern Applied Statistical Methods, 4(2): ??, ???? 2005. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/206.

Eidous:2013:AIC

[EAS13] Omar Eidous and Samar Al-Salman. Akaike information criterion to select the parametric detection function for kernel estimator using line transect data. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/680.

Eidous:2004:KBE

[EB04] Omar M. Eidous and Ayman Baklizi. Kernel-based estimation of P (X < Y) with paired data. Journal of Modern Applied Statistical Methods, 3(1):143–148, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/123.

Ebrahem:2006:BIP

[Ebr06] Mohammed Al-Haj Ebrahem. Bootstrap intervals of the parameters of lognormal distribution using power rule model and accelerated life tests. Journal of Modern Applied Statistical Methods, 5(2):381–389, November 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/255.

Elam:2008:TSS

[EC08] Matthew E. Elam and Kenneth E. Case. Two-stage short-run (X, MR) control charts. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/362.

Egecioglu:2011:LCB

[EG11] Ömer Eğecioğlu and Ayça Ebru Giritligil. The likelihood of choosing the Borda-winner with partial preference rankings of the electorate. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/549.

Eickhoff:2005:QML

[Eic05] Jens C. Eickhoff. Quasi-maximum likelihood estimation for latent variable models with mixed continuous and polytomous data. Journal of Modern Applied Statistical Methods, 4(2):473– 481, November 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/203.

Eidous:2004:SIK

[Eid04] Omar M. Eidous. Some improvements in kernel estimation using line transect sampling. Journal of Modern Applied Statistical Methods, 3(1):149–157, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/124.

Evangelaras:2003:SPD

[EK03] H. Evangelaras and Christos Koukouvinos. Screening properties and design selection of certain two-level designs. Journal of Modern Applied Statistical Methods, 2(1):87–107, May 2003. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/69.

ElMontasser:2011:OTS

[El 11] Ghassen El Montasser. The overall F -tests for seasonal unit roots under nonstationary alternatives: Some theoretical results and a Monte Carlo investigation. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/529.

Elam:2006:TNU

[Ela06] Matthew E. Elam. Two new unbiased point estimates of a population variance. Journal of Modern Applied Statistical Methods, 5(1):194–216, May 2006. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/237.

Elam:2008:DRP

[Ela08] Matthew E. Elam. Delete and revise procedures for two-stage short-run control charts. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/389.

Eledum:2017:MCS

[Ele17] Hussein Yousif Eledum. A Monte Carlo study of the effects of variability and outliers on the linear correlation coefficient. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/936.

Edionwe:2014:LBI

[EM14] Efosa Edionwe and Julian L. Mbegbu. Local bandwidths for improving performance statistics of model-robust regression 2. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/747.

Eidous:2010:CSB

[EME10] Omar M. Eidous, Mohammad Abd Alrahem Shafeq Marie, and Mohammed H. Baker Al-Haj Ebrahem. A comparative study for bandwidth selection in kernel density estimation. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/484.

Espinosa:2023:LEF

[EN23] Oscar Espinosa and Fabio H. Nieto. Leverage effect on financial series using a Bayesian TAR model. Journal of Modern Applied Statistical Methods, 22(2):??, November 2023. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1193.

Eni:2009:PRA

[Eni09] Daniel Eni. Performance ratings of an autocovariance base estimator (ABE) in the estimation of GARCH model parameters when the normality assumption is invalid. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/446.

Erdogdu:2009:IAA

[EO09] Oya S. Erdogdu and Levent Ozbek. Industrialization in animal agriculture: a Kalman filter analysis. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/421.

Ebuh:2013:ITA

[EO13] G. U. Ebuh and I. C. A. Oyeka. Intrinsically ties adjusted nonparametric method for the analysis of two sampled data. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/667.

El-Sherpieny:2020:NGF

[ESAH20] El-Sayed A. El-Sherpieny, Salwa Assar, and Tamer Helal. A new generating family of distributions: Properties and applications to the Weibull exponential model. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1086.

Ercan:2007:ECA

[EYS^{+}07] Ilker Ercan, Berna Yazici, Deniz Sigirli, Bulent Ediz, and Ismet Kan. Examining Cronbach alpha, theta, omega reliability coefficients according to sample size. Journal of Modern Applied Statistical Methods, 6(1):291–303, May 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/302.

Famoye:2011:EAG

[FA11] Felix Famoye and Oluwakemi Aremu. Error analysis on the generalized negative binomial distribution. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/560.

Feroze:2013:BEP

[FA13] Navid Feroze and Muhammad Aslam. On Bayesian estimation and predictions for two-component mixture of the Gompertz distribution. Journal of Modern Applied Statistical Methods, 12(2): ??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/675.

Fahoome:2002:JRF

[Fah02a] Gail F. Fahoome. JMASM 1: RANGEN 2.0 (Fortran 90/95). Journal of Modern Applied Statistical Methods, 1(1):182–190, May 2002. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/25.

Fahoome:2002:TNS

[Fah02b] Gail F. Fahoome. Twenty nonparametric statistics and their large sample approximations. Journal of Modern Applied Statistical Methods, 1(2):248–268, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/32.

Fan:2010:IMM

[Fan10] Weihua Fan. Impact of measurement model modification on structural parameter integrity when measurement model is misspecified. Journal of Modern Applied Statistical Methods, 9(1): ??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/470.

Fay:2002:JCV

[Fay02] Bruce R. Fay. JMASM 4: Critical values for four nonparametric and/or distribution-free tests of location for two independent samples. Journal of Modern Applied Statistical Methods, 1(2): 489–517, November 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/55.

Fay:2006:ETE

[Fay06] Bruce R. Fay. The effect on Type I error and power of various methods of resolving ties for six distribution-free tests of location. Journal of Modern Applied Statistical Methods, 5(1):

22–40, May 2006. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/224.

Fuchs:2004:QPC

[FB04] Camil Fuchs and Vance W. Berger. Quantifying the proportion of cases attributable to an exposure. Journal of Modern Applied Statistical Methods, 3(1):54–64, May 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/116.

Furlow:2010:EMI

[FB10] Carolyn F. Furlow and S. Natasha Beretvas. An evaluation of multiple imputation for meta-analytic structural equation modeling. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/472.

Fang:2008:MCP

[FBR^{+}08] Hua Fang, Gordon P. Brooks, Maria L. Rizzo, Kimberly A. Espy, and Robert S. Barcikowski. A Monte Carlo power analysis of traditional repeated measures and hierarchical multivariate linear models in longitudinal data analysis. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/348.

Feng:2004:MCE

[FC04] Du Feng and Norman Cliff. Monte Carlo evaluation of ordinal d with improved confidence interval. Journal of Modern Applied Statistical Methods, 3(2):322–332, November 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/141.

Feng:2009:JDA

[FC09] Du Feng and Normal Clif. JMASM 29: Dominance analysis of independent data (Fortran). Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/459.

Feroze:2013:PEB

[FEB13] Navid Feroze and Ibrahim El-Batal. Parameter estimations based on Kumaraswamy progressive Type II censored data with random removals. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/678.

Ferrieri:2010:MO

[Fer10] Gaetano Ferrieri. Measuring openness. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/475.

Ferrieri:2012:FIO

[Fer12] Gaetano Ferrieri. Ferrieri’s index of openness applied to remittances to developing countries. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/629.

Ferreira:2017:NEP

[Fer17] Marta Ferreira. A new estimator for the Pickands dependence function. Journal of Modern Applied Statistical Methods, 16(1): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/901.

Finch:2008:CFL

[FF08a] W. Holmes Finch and Brian F. French. Comparing factor loadings in exploratory factor analysis: a new randomization test. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/372.

Finch:2008:UEF

[FF08b] W. Holmes Finch and Brian F. French. Using exploratory factor analysis for locating invariant referents in factor invariance studies. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/357.

Finch:2012:PEM

[FF12] W. Holmes Finch and Brian F. French. Parameter estimation with mixture item response theory models: a Monte Carlo comparison of maximum likelihood and Bayesian methods. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/596.

Finch:2013:MCC

[FF13] Holmes Finch and Brian French. A Monte Carlo comparison of robust MANOVA test statistics. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/663.

French:2016:FIT

[FF16] Brian F. French and Holmes Finch. Factorial invariance testing under different levels of partial loading invariance within a multiple group confirmatory factor analysis model. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/811.

Finch:2011:CFR

[Fin11] W. Holmes Finch. A comparison of factor rotation methods for dichotomous data. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/564.

Finch:2014:CMG

[Fin14] Holmes Finch. A comparison of methods for group prediction with high dimensional data. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/722.

Finney:2019:CEN

[Fin19] Tia L. Finney. Confirmative evaluation: New CIPP evaluation model. Journal of Modern Applied Statistical Methods, 18(2): ??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1065.

Fradette:2003:CRP

[FKL^{+}03] Katherine Fradette, H. J. Keselman, Lisa Lix, James Algina, and Rand R. Wilcox. Conventional and robust paired and independent-samples t tests: Type I error and power rates. Journal of Modern Applied Statistical Methods, 2(2):481–496, November 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/106.

Fadhilah:2025:CRR

[FKP^{+}25] Rahmi Fadhilah, Heri Kuswanto, Dedy Dwi Prastyo, Dinda Ayu Safira, and M. Y. Matdoan. Comparison of RACOG and RACOG-RUS for classifying imbalanced data on gradient boosting and naïve Bayes performance. Journal of Modern Applied Statistical Methods, 24(1):89–104, October 2025. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1336.

Famoye:2004:BND

[FLE04] Felix Famoye, Carl Lee, and Nicholas Eugene. Beta–normal distribution: Bimodality properties and application. Journal of Modern Applied Statistical Methods, 3(1):85–103, May 2004. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/119.

Fog:2015:PRN

[Fog15] Agner Fog. Pseudo-random number generators for vector processors and multicore processors. Journal of Modern Applied Statistical Methods, 14(1):308–334, May 5, 2015. CODEN ???? ISSN

1538-9472. URL http://digitalcommons.wayne.edu/jmasm/
vol14/iss1/23;   https://jmasm.com/index.php/jmasm/article/
view/768;   https://www.agner.org/random/theory/randomvector.
pdf.

Famoye:2003:VSP

[FR03] Felix Famoye and Daniel E. Rothe. Variable selection for Poisson regression model. Journal of Modern Applied Statistical Methods, 2(2):380–388, November 2003. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/95.

Fu:2004:SMA

[FR04] Pingfu Fu and J. Sunil Rao. On a simple method for analyzing multivariate survival data using sample survey methods. Journal of Modern Applied Statistical Methods, 3(2):345–356, November 2004. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/143.

Frane:2015:PFT

[Fra15] Andrew V. Frane. Are per-family Type I error rates relevant in social and behavioral science? Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/750.

Frane:2017:EPA

[Fra17] Andrew V. Frane. Errors in a program for approximating confidence intervals. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/923.

Frane:2019:MOM

[Fra19] Andrew V. Frane. Misguided opposition to multiplicity adjustment remains a problem. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1068.

Fu:2002:REM

[FRJ02] Pingfu Fu, J. Sunil Rao, and Jiming Jiang. Robust estimation of multivariate failure data with time-modulated frailty. Journal of Modern Applied Statistical Methods, 1(2):367–378, November 2002. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/42.

Felder:2015:SED

[FS15] Valerie Felder and Shlomo S. Sawilowsky. Special education distributions and analysis. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/765.

Fukuda:2005:MSB

[Fuk05] Kosei Fukuda. Model-selection-based monitoring of structural change. Journal of Modern Applied Statistical Methods, 4(1): 155–162, May 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/177.

Fukuda:2007:PUR

[Fuk07] Kosei Fukuda. Practical unit-root analysis using information criteria: Simulation evidence. Journal of Modern Applied Statistical Methods, 6(1):258–264, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/299.

Frankland:2002:QBP

[FZ02] B. W. Frankland and Bruno D. Zumbo. Quantifying bimodality Part I: an easily implemented method using SPSS. Journal of Modern Applied Statistical Methods, 1(1):??, ???? 2002. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/23.

Frankland:2009:QBP

[FZ09] B. W. Frankland and Bruno D. Zumbo. Quantifying bimodality Part 2: a likelihood ratio test for the comparison of a unimodal normal distribution and a bimodal mixture of two normal distributions. Journal of Modern Applied Statistical Methods, 8(1): ??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/399.

Ghosh:2017:VBD

[GA17] Dilip Kumar Ghosh and Sangeeta Ahuja. On variance balanced designs. Journal of Modern Applied Statistical Methods, 16(2): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/930.

Ganguly:2012:ELR

[Gan12] Shyam S. Ganguly. Exact logistic regression for a matched pairs case-control design with polytomous exposure variables. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/624.

Ganguly:2014:RRA

[Gan14] S. S. Ganguly. Robust regression analysis for non-normal situations under symmetric distributions arising in medical research. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/715.

Gao:2004:DBP

[Gao04] Yonghong Gao. Depth based permutation test for general differences in two multivariate populations. Journal of Modern Applied Statistical Methods, 3(1):49–53, May 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/115.

Gauraha:2017:GLL

[Gau17] Niharika Gauraha. Graphical log-linear models: Fundamental concepts and applications. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/912.

Gruman:2007:EHT

[GCAC07] Jamie A. Gruman, Robert A. Cribbie, and Chantal A. Arpin-Cribbie. The effects of heteroscedasticity on tests of equivalence. Journal of Modern Applied Statistical Methods, 6(1):133–140, May 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/288.

Gierl:2008:UCM

[GCH08] Mark J. Gierl, Ying Cui, and Steve Hunka. Using connectionist models to evaluate examinees’ response patterns to achievement tests. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/358.

Garg:2024:PAR

[GCP24] Neha Garg, Ambika C, and Menakshi Pachori. Proportionality adjusted ratio-type calibration estimators of population mean under stratified sampling. Journal of Modern Applied Statistical Methods, 23(2):??, January 2024. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1194.

Gagne:2002:BRM

[GD02] Phill Gagné and C. Mitchell Dayton. Best regression model using information criteria. Journal of Modern Applied Statistical Methods, 1(2):479–488, November 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/54.

Ghosh:2020:PBD

[GDG20] D. K. Ghosh, N. R. Desai, and Shreya Ghosh. Pairwise balanced designs from cyclic PBIB designs. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1073.

George:2014:CSS

[Geo14] Florence George. A comparison of shape and scale estimators of the two-parameter Weibull distribution. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/689.

Gosky:2009:CSB

[GG09] Ross M. Gosky and Sujit K. Ghosh. A comparative study of Bayesian model selection criteria for capture–recapture models for closed populations. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/400.

Ghahramani:2014:IC

[Gha14] Masume Ghahramani. The information criterion. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/742.

Gray:2015:CBC

[GHSF^{+}15] Kathy Gray, Brittany Hampton, Tony Silveti-Falls, Allison Mc-Connell, and Casey Bausell. Comparison of Bayesian credible intervals to frequentist confidence intervals. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/753.

Gideon:2007:CC

[Gid07] Rudy A. Gideon. The correlation coefficients. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/324.

Giles:2024:PML

[Gil24] David E. Giles. The performance of the maximum likelihood estimator for the bell distribution for count data. Journal of Modern Applied Statistical Methods, 23(2):??, January 2024. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1218.

Gould:2017:PUP

[GJ17] Anthony M. Gould and Jean-Etienne Joullié. ’Parallel universe’ or ’proven future’? The language of dependent means t-test interpretations. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/934.

Gercsik:2004:CHT

[GK04] Vladimir Gercsik and Mark Kelbert. On comparison of hypothesis tests in the Bayesian framework without loss function. Journal of Modern Applied Statistical Methods, 3(2):399–405, November 2004. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/147.

Goyal:2020:NPT

[GK20] Manish Goyal and Narinder Kumar. Non-parametric tests for testing of scale parameters. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1089.

Good:2006:LAV

[GL06] Phillip I. Good and Clifford E. Lunneborg. Limitations of the analysis of variance. Journal of Modern Applied Statistical Methods, 5(1):41–43, May 2006. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/225.

G:2005:NPT

[GMS05] Rosa Arboretti G., Marco Marozzi, and Luigi Salmaso. Nonparametric pooling and testing of preference ratings for full-profile conjoint analysis experiments. Journal of Modern Applied Statistical Methods, 4(2):545–552, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/211.

Good:2002:ECE

[Goo02] Phillip I. Good. Extensions of the concept of exchangeability and their applications. Journal of Modern Applied Statistical Methods, 1(2):243–247, November 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/31.

Good:2009:AME

[Goo09] Phillip I. Good. Analysis of MultiFactor experimental designs. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/428.

Good:2013:VVS

[Goo13] Phillip I. Good. p-values versus significance levels. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/634.

Guerrero:2004:VSP

[GP04] Victor M. Guerrero and Rafael Perera. Variance stabilizing power transformation for time series. Journal of Modern Applied Statistical Methods, 3(2):357–369, November 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/144.

George:2011:EPJ

[GR11] Florence George and K. M. Ramachandran. Estimation of parameters of Johnson’s system of distributions. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/559.

Goyal:2020:BIL

[GRM20] Teena Goyal, Piyush Kant Rai, and Sandeep K. Maurya. Bayesian inference for logarithmic transformed exponential distribution in presence of adaptive progressive Type-II censoring. Journal of Modern Applied Statistical Methods, 19(2):??, July 2020. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1106.

Gupta:2006:AWR

[GS06] Sat Gupta and Javid Shabbir. An alternative to Warner’s randomized response model. Journal of Modern Applied Statistical Methods, 5(2):328–331, November 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/249.

Gosky:2019:SSI

[GS19] Ross M. Gosky and Joel Sanqui. A simulation study on increasing capture periods in Bayesian closed population capture– recapture models with heterogeneity. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1031.

Garg:2024:MER

[GSP24] Neha Garg, Housila P. Singh, and Menakshi Pachori. Modified exponential ratio-type estimator of population mean in stratified sampling using calibration approach. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1145.

Green:2010:CSS

[GT10] Samuel B. Green and Marilyn S. Thompson. Can specification searches be useful for hypothesis generation? Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/474.

Gullu:2018:SSN

[GT18] Özlem Güllü and Mustafa Agah Tekindal. Sample size for noninferiority tests for one proportion: a simulation study. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/964.

Gemperli:2003:FGL

[GV03] Armin Gemperli and Penelope Vounatsou. Fitting generalized linear mixed models for point-referenced spatial data. Journal of Modern Applied Statistical Methods, 2(2):497–511, November 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/107.

Giarola:2018:EMS

[GVCM18] Luciane Teixeira Passos Giarola, Mario Javier Ferrua Vivanco, Marcelo Angelo Cirillo, and Fortunato Silva Menezes. Extended method for several dichotomous covariates to estimate the instantaneous risk function of the Aalen additive model. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/979.

Guan:2009:BSS

[GW09] Liming Guan and John P. Wendell. Bias in stabilized sieve sampling. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/417.

Gelin:2007:OCD

[GZ07] Michaela N. Gelin and Bruno D. Zumbo. Operating characteristics of the DIF MIMIC approach using Jöreskog’s covariance matrix with ML and WLS estimation for short scales. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/330.

Grine:2017:PQL

[GZ17] Razika Grine and Halim Zeghdoudi. On Poisson quasi-Lindley distribution and its applications. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/944.

Hassan:2007:NGN

[HA07] Anwar Hassan and Sheikh Nilal Ahmad. A new generalization of negative Ploya–Eggenberger distribution and its applications. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/338.

Hassan:2014:LTA

[HA14] Anwar Hassan and Mehraj Ahmad. Life testing analysis of failure censored generalized exponentiated data. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/733.

Hall:2004:MDS

[Hal04] Bruce W. Hall. Mentoring doctoral students: a personal perspective. Journal of Modern Applied Statistical Methods, 3(2):

576–577, November 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/150.

Hancock:2003:FCM

[Han03] Gregory R. Hancock. Fortune cookies, measurement error, and experimental design. Journal of Modern Applied Statistical Methods, 2(2):293–305, November 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/87.

Harlow:2010:SRR

[Har10] Lisa L. Harlow. On scientific research: the role of statistical modeling and hypothesis testing. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/493.

Harwell:2018:SUB

[Har18] Michael Harwell. A strategy for using bias and RMSE as outcomes in Monte Carlo studies in statistics. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/991.

Harwell:2019:ITE

[Har19] Michael Harwell. The importance of Type I error rates when studying bias in Monte Carlo studies in statistics. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1012.

Hancock:2008:SOL

[HB08a] Gregory R. Hancock and Michelle M. Buehl. Second-order latent growth models with shifting indicators. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/344.

Hassan:2008:SPQ

[HB08b] Anwar Hassan and Sheikh Bilal. On some properties of quasinegative-binomial distribution and its applications. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/395.

Hassan:2012:SNI

[HB12] Anwar Hassan and Sheikh Bilal. On some negative integer moments of quasi-negative-binomial distribution. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/617.

Hussain:2020:DSR

[HBA20] Jaffer Hussain, S. Balamurali, and Muhammad Aslam. Design of SkSP-r plan for popular statistical distributions. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1091.

Headrick:2002:JMS

[Hea02] Todd C. Headrick. JMASM 3: a method for simulating systems of correlated binary data. Journal of Modern Applied Statistical Methods, 1(1):195–201, May 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/27.

Headrick:2003:JAG

[Hea03] Todd C. Headrick. JMASM 6: an algorithm for generating exact critical values for the Kruskal–Wallis one-way ANOVA. Journal of Modern Applied Statistical Methods, 2(1):268–271, May 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/83.

Headrick:2004:PTS

[Hea04] Todd C. Headrick. On polynomial transformations for simulating multivariate non-normal distributions. Journal of Modern Applied Statistical Methods, 3(1):65–71, May 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/117.

Headrick:2006:JNC

[Hea06] Todd C. Headrick. JMASM 24: Numerical computing for thirdorder power method polynomials (Excel). Journal of Modern Applied Statistical Methods, 5(2):567–574, November 2006. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/273.

Heiser:2006:STT

[Hei06] David A. Heiser. Statistical tests, tests of significance, and tests of a hypothesis using Excel. Journal of Modern Applied Statistical Methods, 5(2):551–566, November 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/272.

Heo:2002:EDR

[Heo02] Moonseong Heo. Exploration of distributions of ratio of partial sum of sample eigenvalues when all population eigenvalues are the same. Journal of Modern Applied Statistical Methods, 1(2): 420–427, November 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/48.

Heo:2008:UWW

[Heo08] Moonseong Heo. Utility of weights for weighted kappa as a measure of interrater agreement on ordinal scale. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/356.

Holland:2016:TMI

[HH16] Christopher H. Holland and Tom Holland. A two and more independent factor form of the Wilcoxon–Mann–Witney test, extendable to other permutation-based tests. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/793.

Hassan:2020:EEP

[HHN20] Amal S. Hassan, Saeed Elsayed Hemeda, and Said G. Nassr. On the extension of exponentiated Pareto distribution. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1087.

Higgins:2004:STR

[Hig04] James J. Higgins. Statistics and technology: Reflections on 35 years of change. Journal of Modern Applied Statistical Methods, 3(2):567–575, November 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/149.

Hittner:2009:EPD

[Hit09] James B. Hittner. Effects of population distribution, sample size and correlation structure on Huberty’s effect size R. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/402.

Holgersson:2010:MBV

[HK10] H. E. T. Holgersson and Peter S. Karlsson. Model based vs. model independent tests for cross-correlation. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/467.

Harwell:2017:EDD

[HKP17] Michael Harwell, Nidhi Kohli, and Yadira Peralta. Experimental design and data analysis in computer simulation studies in the behavioral sciences. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/925.

Hyslop:2003:NFT

[HL03] Terry Hyslop and Paul J. Lupinacci. A nonparametric fitted test for the Behrens–Fisher problem. Journal of Modern Applied Statistical Methods, 2(2):414–424, November 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/98.

Hasselblad:2007:TTC

[HL07] Vic Hasselblad and Yulia Lokhnygina. Tests for 2 × 2 tables in clinical trials. Journal of Modern Applied Statistical Methods, 6 (2):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/318.

Harring:2016:CEM

[HL16] Jeffrey R. Harring and Junhui Liu. A comparison of estimation methods for nonlinear mixed-effects models under model misspecification and data sparseness: a simulation study. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/812.

Hall:2008:BCI

[HM08] Matthew Hall and Matthew S. Mayo. Bootstrap confidence intervals and coverage probabilities of regression parameter estimates using trimmed elemental estimation. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/386.

Heiny:2010:ESD

[HM10] Erik L. Heiny and Daniel J. Mundform. The effectiveness of stepwise discriminant analysis as a post hoc procedure to a significant MANOVA. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/464.

Huitema:2007:TSI

[HML07] Bradley E. Huitema, Joseph W. McKean, and Sean Laraway. Time-series intervention analysis using ITSACORR: Fatal flaws. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/312.

Hall:2007:GLM

[HMNW07] Sandra Hall, Matthew S. Mayo, Xu-Feng Niu, and James C. Walker. Generalized linear mixed-effects models for the analysis of odor detection data. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/333.

Hosseini:2019:MLE

[HNG19] Saman Hosseini, Parviz Nasiri, and Sharad Damodar Gore. Maximum likelihood estimations based on upper record values for probability density function and cumulative distribution function in exponential family and investigating some of their properties. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1050.

Hamdan:2005:USM

[HNWD05] Hasan Hamdan, John Nolan, Melanie Wilson, and Kristen Dardia. Using scale mixtures of normals to model continuously compounded returns. Journal of Modern Applied Statistical Methods, 4(1):214–226, May 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/182.

Holgersson:2006:SNN

[Hol06] H. E. T. Holgersson. Simulation of non-normal autocorrelated variables. Journal of Modern Applied Statistical Methods, 5(2): 408–416, November 2006. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/258.

Haughton:2003:BAP

[HP03] Dominique Haughton and Nguyen Phong. Bayesian analysis of poverty rates: the case of Vietnamese provinces. Journal of Modern Applied Statistical Methods, 2(1):189–194, May 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/77.

Haganawiga:2023:GDI

[HPSR23] Kumur John Haganawiga, Surya Kant Pal, Anu Sirohi, and Varsha Rathi. Gender discrepancy in infant mortality in Nigeria: Evidence from NDHS Data 2018. Journal of Modern Applied Statistical Methods, 22(1):??, February 2023. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1131.

Hashemizadeh:2016:ALM

[HR16] Elham Hashemizadeh and Sohrab Rahbar. The application of Legendre multiwavelet functions in image compression. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/864.

Hussain:2008:LEU

[HS08] Zawar Hussain and Javid Shabbir. Logit estimation using Warner’s randomized response model. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/350.

Hill:2011:BMC

[HS11] Jack C. Hill and Shlomo S. Sawilowsky. Bias in Monte Carlo simulations due to pseudo-random number generator initial seed selection. Journal of Modern Applied Statistical Methods, 10(1):

??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/521.

Herrera:2018:DSQ

[HS18] Carlos Narciso Bouza Herrera and Prayas Sharma. Dealing with sensitive quantitative variables: a comparison of sampling designs for the procedure of Gupta and Thornton. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/976.

Habshah:2018:RHC

[HSA18] M. Habshah, Muhammad Sani, and Jayanthi Arasan. Robust heteroscedasticity consistent covariance matrix estimator based on robust Mahalanobis distance and diagnostic robust generalized potential weighting methods in linear regression. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/970.

Hurley:2003:NMN

[Hur03] W. J. Hurley. A note on MLEs for normal distribution parameters based on disjoint partial sums of a random sample. Journal of Modern Applied Statistical Methods, 2(2):475–477, November 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/104.

Hurley:2007:NPT

[Hur07] W. J. Hurley. A note on probability trees. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/336.

Hurley:2008:LNS

[Hur08] W. J. Hurley. On the length of NHL shootouts. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/366.

Hurley:2009:IAC

[Hur09] W. J. Hurley. An inductive approach to calculate the MLE for the double exponential distribution. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/451.

Huson:2007:PSC

[Hus07] L. W. Huson. Performance of some correlation coefficients when applied to zero-clustered data. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/325.

Harring:2011:PIS

[HW11] Jeffrey R. Harring and John A. Wasko. Probabilistic inferences for the sample Pearson product moment correlation. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/558.

Herawati:2024:PRR

[HWS^{+}24] Netti Herawati, Ameliana Wijayanti, Agus Sutrisno, Nusyirwan, and Misgiyati. The performance of ridge regression, LASSO, and elastic-net in controlling multicollinearity: a simulation and application. Journal of Modern Applied Statistical Methods, 23 (2):??, January 2024. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1258.

Hertzberg:2006:RQI

[HXH06] Vicki Stover Hertzberg, Frank Xu, and Michael Haber. Restricted quasi-independent model resolves paradoxical behaviors of Cohen’s kappa. Journal of Modern Applied Statistical Methods, 5(2):417–431, November 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/259.

Hu:2003:SRM

[HY03] Jianhua Hu and Guosheng Yin. A semiparametric regression model for oligonucleotide arrays. Journal of Modern Applied Statistical Methods, 2(1):256–267, May 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/82.

Hossianzadeh:2016:BAD

[HZ16] A. Hossianzadeh and K. Zare. Bayesian analysis of discrete skewed Laplace distribution. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/873.

Iseh:2024:SEP

[IB24] Matthew Iseh and Mbuotidem Bassey. Smoothing of estimators of population mean using calibration technique with sample errors. Journal of Modern Applied Statistical Methods, 23(1):??, October 2024. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1142.

Islam:2007:CDM

[ICS07] M. A. Islam, R. I. Chowdhury, and K. P. Singh. Covariate dependent Markov models for analysis of repeated binary outcomes. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/329.

Ismail:2006:CRC

[IJ06] Noriszura Ismail and Abdul Aziz Jemain. A comparison of risk classification methods for claim severity data. Journal of Modern Applied Statistical Methods, 5(2):513–528, November 2006. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/268.

Ismail:2008:CIS

[IJ08] Noriszura Ismail and Abdul Aziz Jemain. Construction of insurance scoring system using regression models. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/394.

Im:2020:PBB

[Im20] Seongah Im. Performance of the beta-binomial model for clustered binary responses: Comparison with generalized estimating equations. Journal of Modern Applied Statistical Methods, 19 (1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/1084.

Irfan:2024:BEP

[IS24] Mohd Irfan and A. K. Sharma. Bayesian estimation and prediction for inverse power Maxwell distribution with applications to tax revenue and health care data. Journal of Modern Applied Statistical Methods, 23(1):??, October 2024. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1141.

Iki:2018:MDM

[IT18] Kiyotaka Iki and Sadao Tomizawa. Measure of departure from marginal average point-symmetry for two-way contingency tables. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1003.

Jan:2018:BAT

[JA18] Uzma Jan and S. P. Ahmad. Bayesian approximation techniques for scale parameter of Laplace distribution. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/996.

Jafari:2010:FMT

[JADB10] Peyman Jafari, Noori Akhtar-Danesh, and Zahra Bagheri. A flexible method for testing independence in two-way contingency tables. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/500.

Jan:2024:MOA

[JAJ24] Rameesa Jan, Peer Bilal Ahmad, and T. R. Jan. Marshall–Olkin alpha power inverse Lindley distribution and its applications. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1151.

Jayabalan:2017:SMC

[Jay17] Jayabharathiraj Jayabalan. Stochastic model for cancer cell growth through single forward mutation. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/913.

Jinn:2004:PP

[Jin04] Jann-Huei Jinn. The President’s problem. Journal of Modern Applied Statistical Methods, 3(2):476–484, November 2004. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/158.

Jadhav:2014:RWS

[JK14] Nileshkumar H. Jadhav and D. N. Kashid. Robust Winsorized shrinkage estimators for linear regression model. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/724.

Jia:2023:INC

[JPH23] Hao Jia, Yadira Peralta, and Michael Harwell. The influence of number of clusters and level-2 residual distribution on multilevel parameter estimates: a latin hypercube design perspective. Journal of Modern Applied Statistical Methods, 22(1):??, February 2023. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1197.

Jacyna:2016:DBB

[JR16] Garry M. Jacyna and Scott L. Rosen. Developing Bayesianbased confidence bounds for non-identically distributed observations using the Lyapunov condition. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/866.

Jaggi:2010:NBB

[JVA10] Seema Jaggi, Cini Varghese, and N. R. Abeynayake. Neighbor balanced block designs for two factors. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/501.

Kadilar:2016:NET

[Kad16] Gamze Özel Kadilar. A new exponential type estimator for the population mean in simple random sampling. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/848.

Elaal:2021:TIK

[kAEAE21] Mervat khalifa Abd Ellah Abd Elaal. Truncated inverted Kumaraswamy: Estimation and predication based on maximum likelihood method. Journal of Modern Applied Statistical Methods, 20(1):??, February 2021. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1109.

Keselman:2005:RCI

[KAF05] H. J. Keselman, James Algina, and Katherine Fradette. Robust confidence intervals for effect size in the two-group case. Journal of Modern Applied Statistical Methods, 4(2):353–371, November 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/194.

Kanyongo:2005:DCN

[Kan05] Gibbs Y. Kanyongo. Determining the correct number of components to extract from a principal components analysis: a Monte Carlo study of the accuracy of the scree plot. Journal of Modern Applied Statistical Methods, 4(1):120–133, May 2005. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/174.

Kanyongo:2006:IRF

[Kan06] Gibbs Y. Kanyongo. The influence of reliability on four rules for determining the number of components to retain. Journal of Modern Applied Statistical Methods, 5(2):332–343, November 2006. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/250.

Kibria:2016:SRR

[KB16] B. M. Golam Kibria and Shipra Banik. Some ridge regression estimators and their performances. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/797.

Kanyongo:2007:RSP

[KBKBG07] Gibbs Y. Kanyongo, Gordon P. Brook, Lydia Kyei-Blankson, and Gulsah Gocmen. Reliability and statistical power: How measurement fallibility affects power and required sample sizes for several parametric and nonparametric statistics. Journal of Modern Applied Statistical Methods, 6(1):81–90, May 2007. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/284.

Karioti:2006:PGM

[KC06] Vasiliki Karioti and Chrys Caroni. Properties of the GAR(1) model for time series of counts. Journal of Modern Applied Statistical Methods, 5(1):140–151, May 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/233.

Kumar:2020:SRS

[KC20] Sanjay Kumar and Priyanka Chhaparwal. A simple random sampling modified dual to product estimator for estimating population mean using order statistics. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1078.

Kumar:2017:URT

[KCA^{+}17] Jitendra Kumar, Anoop Chaturvedi, Umme Afifa, Shafat Yousuf, and Saurabh Kumar. Unit root test for panel data AR(1) time series model with linear time trend and augmentation term: a Bayesian approach. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/931.

Kroopnick:2010:ACB

[KCCD10] Marc H. Kroopnick, Jinsong Chen, Jaehwa Choi, and C. Mitchell Dayton. Assessing classification bias in latent class analysis: Comparing resubstitution and leave-one-out methods. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/465.

Kumar:2018:URV

[KD18] Devendra Kumar and Sanku Dey. Upper record values from extended exponential distribution. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1002.

Kumar:2018:MGF

[KDMAA18] Devendra Kumar, Sanku Dey, Mansoor Rashid Malik, and Fahad M. Al-Aboud. Moment generating functions of complementary exponential-geometric distribution based on k-th lower record values. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/963.

Khedhiri:2012:ESK

[KE12] Sami Khedhiri and Ghassen El Montasser. An extension of the seasonal KPSS test. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/588.

Kellow:2005:EFA

[Kel05] J. Thomas Kellow. Exploratory factor analysis in two measurement journals: Hegemony by default. Journal of Modern Applied Statistical Methods, 4(1):283–287, May 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/187.

Keselman:2015:PFF

[Kes15] H. J. Keselman. Per family or familywise Type I error control: “Eether, eyether, neether, nyther, let’s call the whole thing off!”. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/751.

Koran:2016:JPB

[KH16] Jennifer Koran and Todd C. Headrick. JMASM 35: a percentilebased power method: Simulating multivariate non-normal continuous distributions (SAS). Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/830.

Khalaf:2012:IEP

[Kha12a] Ghadban Khalaf. Improved estimator in the presence of multicollinearity. Journal of Modern Applied Statistical Methods, 11 (1):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/594.

Khalaf:2012:PRP

[Kha12b] Ghadban Khalaf. A proposed ridge parameter to improve the least square estimator. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/623.

Khalaf:2013:CBB

[Kha13] Ghadban Khalaf. A comparison between biased and unbiased estimators in ordinary least squares regression. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/676.

Kuncheva:2005:SIB

[KHC05] Ludmila I. Kuncheva, Zoë S. J. Hoare, and Peter D. Cockcroft. Selection of independent binary features using probabilities: an example from veterinary medicine. Journal of Modern Applied Statistical Methods, 4(2):528–537, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/209.

Khoo:2004:AQC

[Kho04] Michael B. C. Khoo. An alternative Q chart incorporating a robust estimator of scale. Journal of Modern Applied Statistical Methods, 3(1):72–84, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/118.

Khoo:2005:NAI

[Kho05] Michael B. C. Khoo. A nonrigorous approach of incorporating sensitizing rules into multivariate control charts. Journal of Modern Applied Statistical Methods, 4(2):567–582, November 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/213.

Khoo:2007:MCC

[Kho07] Michael B. C. Khoo. A modified X̄ control chart for samples drawn from finite populations. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/332.

Kamil:2010:MDS

[KI10] Anton Abdulbasah Kamil and Khlipah Ibrahim. Maximum downside semi deviation stochastic programming for portfolio optimization problem. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/510.

Khalaf:2014:RRI

[KI14] Ghadban Khalaf and Mohamed Iguernane. Ridge regression and ill-conditioning. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/735.

Khalaf:2016:MRP

[KI16] Ghadban Khalaf and Mohammed Iguernane. Multicollinearity and a ridge parameter estimation approach. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/858.

Kibria:2006:ASI

[Kib06] B. M. Golam Kibria. Applications of some improved estimators in linear regression. Journal of Modern Applied Statistical Methods, 5(2):367–380, November 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/254.

Kibria:2008:RPI

[Kib08] B. M. Golam Kibria. Robust predictive inference for multivariate linear models with elliptically contoured distribution using Bayesian, classical and structural approaches. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/388.

King:2003:BCI

[Kin03] Jason E. King. Bootstrapping confidence intervals for robust measures of association. Journal of Modern Applied Statistical Methods, 2(2):512–519, November 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/108.

Kumar:2012:SMS

[KIR12] R. Sampath Kumar, M. Indra, and R. Radhakrishnan. Selection of mixed sampling plan with QSS-1(n; cN, cT) plan as attribute plan indexed through MAPD and LQL. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/616.

Kittani:2010:IPT

[Kit10] Hilmi F. Kittani. Incidence and prevalence for a triply censored data. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/504.

Kundu:2006:ATI

[KJ06] Debasis Kundu and Avijit Joarder. Analysis of Type-II progressively hybrid censored competing risks data. Journal of Modern Applied Statistical Methods, 5(1):??, ???? 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/234.

Khan:2012:BEE

[KJ12] Adil H. Khan and T. R. Jan. Bayesian estimation of Erlang distribution under different generalized truncated distributions as priors. Journal of Modern Applied Statistical Methods, 11(2): ??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/622.

Khan:2014:EMC

[KJ14a] Adil H. Khan and T. R. Jan. Estimation of multi component systems reliability in stress-strength models. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/738.

Khan:2014:REG

[KJ14b] Adil H. Khan and T. R. Jan. Reliability estimates of generalized Poisson distribution and generalized geometric series distribution. Journal of Modern Applied Statistical Methods, 13(2): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/737.

Khamis:2010:CBT

[KJI10] Faisal G. Khamis, Abdul Aziz Jemain, and Kamarulzaman Ibrahim. On a comparison between two measures of spatial association. Journal of Modern Applied Statistical Methods, 9(1):

??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/471.

Kundu:2009:TIP

[KJK09] Debasis Kundu, Avijit Joarder, and Hare Krishna. On Type-II progressively hybrid censoring. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/444.

Kamal:2018:DBM

[KJMA18] Murshid Kamal, Syed Aqib Jalil, Syed Mohd Muneeb, and Irfan Ali. A distance based method for solving multi-objective optimization problems. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/985.

Khoolenjani:2013:DRN

[KK13] Nayereh B. Khoolenjani and Kavoos Khorshidian. Distribution of the ratio of normal and Rice random variables. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/686.

Kamble:2016:VSR

[KK16] Tejaswi S. Kamble and Dattatraya N. Kashid. Variable selection in regression using multilayer feedforward network. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/818.

Kumar:2020:BIC

[KK20a] Ashok Kumar and Jitendra Kumar. Bayesian inference for C-AR model with spherically symmetric error. Journal of Modern Applied Statistical Methods, 19(2):??, July 2020. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1107.

Kumar:2020:TSS

[KK20b] Narinder Kumar and Arun Kumar. Two sample statistical test for location parameters. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1094.

Kumar:2019:COS

[KKK19] Sumit Kumar, M. J. S. Khan, and Surinder Kumar. Concomitant of order statistics from new bivariate Gompertz distribution. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1056.

Katsaragakis:2005:CST

[KKST05] Stylianos Katsaragakis, Christos Koukouvinos, Stella Stylianou, and Eleni-Maria Theodoraki. Comparison of statistical tests in logistic regression: the case of hypernatreamia. Journal of Modern Applied Statistical Methods, 4(2):514–521, November 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/207.

Kafle:2013:BJR

[KKT13] Ram C. Kafle, Netra Khanal, and Chris P. Tsokos. Bayesian joinpoint regression model for childhood brain cancer mortality. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/681.

Kowalchuk:2006:MCP

[KKW^{+}06] Rhonda K. Kowalchuk, H. J. Keselman, Rand R. Wilcox, James Algina, and James Algina. Multiple comparison procedures, trimmed means and transformed statistics. Journal of Modern Applied Statistical Methods, 5(1):??, ???? 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/226.

Krstic:2016:WME

[KKZB16] Goran Krstic, Nikolas S. Krstić, and Mauricio Zambrano-Bigiarini. The br^{2}-weighting method for estimating the effects of air pollution on population health. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/875.

Kleinhans:2019:LMP

[Kle19] Kimberly A. Kleinhans. The logic model, participatory evaluation and out of school art programs. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1063.

Kong:2007:PBG

[KLS07] Lingji Kong, Carl Lee, and J. H. Sepanski. On the properties of beta–gamma distribution. Journal of Modern Applied Statistical Methods, 6(1):187–211, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/293.

Klockars:2002:TER

[KM02] Alan J. Klockars and Tim P. Moses. Type I error rates for rankbased tests of homogeneity of slopes. Journal of Modern Applied Statistical Methods, 1(2):452–460, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/51.

Kawasaki:2014:CTC

[KM14] Yohei Kawasaki and Etsuo Miyaoka. Comparison of three calculation methods for a Bayesian inference of two Poisson parameters. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/712.

Kaur:2018:BSB

[KMA18] Kamaljit Kaur, Kalpana K. Mahajan, and Sangeeta Arora. Bayesian and semi-Bayesian estimation of the parameters of generalized inverse Weibull distribution. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/974.

Kumari:2021:BEP

[KMA21] Ritu Kumari, Kalpana K. Mahajan, and Sangeeta Arora. Bayesian estimation of the parameters and reliability measures of Nakagami distribution. Journal of Modern Applied Statistical Methods, 20(2):??, July 2021. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1114.

Kundu:2006:JCW

[KMG06] Debasis Kundu, Anubhav Manglick, and Rameshwar D. Gupta. JMASM 22: a convenient way of generating normal random variables using generalized exponential distribution. Journal of Modern Applied Statistical Methods, 5(1):266–272, May 2006. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/242.

Khoo:2005:EPS

[KN05] Michael B. C. Khoo and T. F. Ng. Enhancing the performance of a short run multivariate control chart for the process mean. Journal of Modern Applied Statistical Methods, 4(1):227–239, May 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/183.

Knapp:2002:SRS

[Kna02] Thomas R. Knapp. Some reflections on significance testing. Journal of Modern Applied Statistical Methods, 1(2):??, ???? 2002. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/30.

Knapp:2007:BR

[Kna07] Thomas R. Knapp. Bimodality revisited. Journal of Modern Applied Statistical Methods, 6(1):8–20, May 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/278.

Knapp:2015:PD

[Kna15] Thomas R. Knapp. In (partial) defense of 0.05. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/771.

Knapp:2018:SWG

[Kna18] Tom Knapp. Should we give up on causality? Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/990.

Knapp:2019:AAM

[Kna19a] Thomas R. Knapp. Almost all missing data are MNAR. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1044.

Knapp:2019:PSI

[Kna19b] Thomas R. Knapp. A primer on statistical inferences for finite populations. Journal of Modern Applied Statistical Methods, 18 (2):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/1045.

Kock:2016:NNP

[Koc16] Ned Kock. Non-normality propagation among latent variables and indicators in PLS-SEM simulations. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/801.

Kock:2018:SMD

[Koc18] Ned Kock. Single missing data imputation in PLS-based structural equation modeling. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/959.

Konstantopoulos:2013:CMP

[Kon13] Spyros Konstantopoulos. Constructing a more powerful test in two-level block randomized designs. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/639.

Korter:2016:SAF

[Kor16] Grace O. Korter. A spatial analytical framework for examining road traffic crashes. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/813.

Keselman:2013:PTN

[KOW13] H. J. Keselman, Abdul R. Othman, and Rand R. Wilcox. Preliminary testing for normality: Is this a good practice? Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/661.

Kesleman:2016:GLM

[KOW16] Harvey J. Kesleman, Abdul R. Othman, and Rand R. Wilcox. Generalized linear model analyses for treatment group equality when data are non-normal. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/789.

Kantam:2014:LRT

[KPR14] R. R. L. Kantam, M. C. Priya, and M. S. Ravikumar. Likelihood ratio type test for linear failure rate distribution vs. exponential distribution. Journal of Modern Applied Statistical Methods, 13 (1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/697.

Khoo:2006:CIM

[KQC06] Michael B. C. Khoo, S. H. Quah, and C. K. Ch’ng. A combined individuals and moving range control chart. Journal of Modern Applied Statistical Methods, 5(1):248–257, May 2006. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/240.

Kantam:2016:LFC

[KR16] R. R. L. Kantam and M. S. Ravikumar. Limited failure censored life test sampling plan in Burr Type X distribution. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/860.

Krammer:2018:ALR

[Kra18] Georg Krammer. The Andersen likelihood ratio test with a random split criterion lacks power. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1001.

Kalish:2002:ANI

[KRP02] Leslie A. Kalish, Katherine A. Riester, and Stuart J. Pocock. Accounting for non-independent observations in 2 × 2 tables, with application to correcting for family clustering in exposure-risk relationship studies. Journal of Modern Applied Statistical Methods, 1(2):379–386, November 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/43.

Kantam:2013:ETT

[KRR13] R. R. L. Kantam, V. Ramakrishna, and M. S. Ravikumar. Estimation and testing in Type I generalized half logistic distribution. Journal of Modern Applied Statistical Methods, 12(1): ??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/653.

Kantam:2014:ETT

[KRR14] R. R. L. Kantam, V. Ramakrishna, and M. S. Ravikumar. Estimation and testing in Type-II generalized half logistic distribution. Journal of Modern Applied Statistical Methods, 13(1): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/703.

Khoo:2006:REW

[KS06] Michael B. C. Khoo and S. Y. Sim. A robust exponentially weighted moving average control chart for the process mean. Journal of Modern Applied Statistical Methods, 5(2):464–474, November 2006. CODEN ???? ISSN 1538-9472. URL https:/

/jmasm.com/index.php/jmasm/article/view/264.

Kanyongo:2009:RBI

[KS09] Gibbs Y. Kanyongo and James B. Schreiber. Relationship between internal consistency and goodness of fit maximum likelihood factor analysis with varimax rotation. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/436.

Kaihani:2010:OMP

[KS10] Reza Kaihani and Ali Reza Seif. Optimal meter placement by reconciliation conventional measurements and phasor measurement units (PMUs). Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/487.

Kantam:2014:SME

[KS14] R. R. L. Kantam and B. Sriram. Some methods of estimation from censored samples in exponential and gamma models. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/734.

Khatun:2025:EEQ

[KS25a] Mahfuza Khatun and Sikandar Siddiqui. Estimating extreme quantiles of unknown distributions using the peak over thresholds method. Journal of Modern Applied Statistical Methods, 24 (1):23–30, ???? 2025. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1333.

Krishnan:2025:ODC

[KS25b] R. Harihara Krishnan and Ananthi Sheshasaayee. Optimizing diabetes classification: BOA-enhanced ML with EDA and SMOTE. Journal of Modern Applied Statistical Methods, 24(1): ??, ???? 2025. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1397.

Kumar:2010:GCC

[KSB10] Sunil Kumar, Housila P. Singh, and Sandeep Bhougal. A general class of chain-type estimators in the presence of non-response under double sampling scheme. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/508.

Kawasaki:2013:CTC

[KSM13] Yohei Kawasaki, Asanao Shimokawa, and Etsuo Miyaoka. Comparison of three calculation methods for a Bayesian inference of P (π1 > π2). Journal of Modern Applied Statistical Methods, 12 (2):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/674.

Kamali:2010:ENE

[KSS10] Hamid Reza Kamali and Parisa Shahnazari-Shahrezaei. Estimating the non-existent mean and variance of the F -distribution by simulation. Journal of Modern Applied Statistical Methods, 9 (2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/517.

Kumar:2013:IEF

[Kum13] Sunil Kumar. Improved estimators in finite population surveys: Theory and applications. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/647.

Kumar:2014:PME

[KV14] Sunil Kumar and M. Viswanathaiah. Population mean estimation with sub sampling the non-respondents using two phase sampling. Journal of Modern Applied Statistical Methods, 13(1): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/698.

Keselman:2004:PCR

[KWAO04] H. J. Keselman, Rand R. Wilcox, James Algina, and Abdul R. Othman. A power comparison of robust test statistics based on adaptive estimators. Journal of Modern Applied Statistical Methods, 3(1):27–38, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/113.

Keselman:2002:TTS

[KWOF02] H. J. Keselman, Rand R. Wilcox, Abdul R. Othman, and Katherine Fradette. Trimming, transforming statistics, and bootstrapping: Circumventing the biasing effects of heteroscedasticity and nonnormality. Journal of Modern Applied Statistical Methods, 1(2):288–309, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/35.

Kan:2006:ICC

[KY06] Betül Kan and Berna Yazici. The individuals control chart in case of non-normality. Journal of Modern Applied Statistical Methods, 5(2):542–550, November 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/271.

Ke:2008:SNR

[KY08] Weiming Ke and Rui Yao. Selection of non-regular fractional factorial designs when some two-factor interactions are important. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/347.

Koh:2008:MGC

[KZ08] Kim H. Koh and Bruno D. Zumbo. Multi-group confirmatory factor analysis for testing measurement invariance in mixed item format data. Journal of Modern Applied Statistical Methods, 7 (2):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/381.

Kroc:2018:CM

[KZ18] Edward Kroc and Bruno D. Zumbo. Calibration of measurements. Journal of Modern Applied Statistical Methods, 17(2): ??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1000.

Lesser:2014:EGM

[LA14] Martin L. Lesser and Meredith B. Akerman. An exploratory graphical method for identifying associations in r×c contingency tables. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/693.

Leinwander:2018:MIC

[LA18] Aaron J. Leinwander and Mohammad A. Aziz. Modeling insurance claims using flexible skewed and mixture probability distributions. Journal of Modern Applied Statistical Methods, 17(1): ??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/982.

Lukman:2017:MCS

[LAA17] Adewale Folaranmi Lukman, Kayode Ayinde, and Adegoke S. Ajiboye. Monte Carlo study of some classification-based ridge parameter estimators. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/905.

Lantz:2016:PTN

[LAM16] Björn Lantz, Roy Andersson, and Peter Manfredsson. Preliminary tests of normality when comparing three independent samples. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/844.

Loh:2017:IPL

[LAMB17] Yue Fang Loh, Jayanthi Arasan, Habshah Midi, and M. R. Abu Bakar. Inferential procedures for log logistic distribution with doubly interval censored data. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/950.

Landsittel:2009:EMC

[Lan09] Douglas Landsittel. Estimating model complexity of feedforward neural networks. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/439.

Lantz:2013:ICV

[Lan13] Björn Lantz. The impact of continuity violation on ANOVA and alternative methods. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/665.

Lance:2019:JM

[Lan19] Michael Lance. JMASM 53: MicceriRD. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1070.

Leite:2010:PMI

[LB10] Walter Leite and S. Natasha Beretvas. The performance of multiple imputation for likert-type items with missing data. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/466.

Liao:2016:OIA

[LB16] Hongjing Liao and Gordon Brooks. Outlier impact and accommodation methods: Multiple comparisons of Type I error rates. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/808.

Liu:2019:FBC

[LB19] Xing Liu and Haiyan Bai. Forward and backward continuation ratio models for ordinal response variables. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1048.

Lloyd:2019:GBL

[LBC^{+}19] Jennifer E. V. Lloyd, Jacqui Boonstra, Lisa Chen, Barry Forer, and Ruth Hershler. A glossary on building longitudinal, population-based data linkages to explore children’s developmental trajectories. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1027.

Lipovetsky:2005:RDS

[LC05] Stan Lipovetsky and Michael Conklin. Regression by data segments via discriminant analysis. Journal of Modern Applied Statistical Methods, 4(1):63–74, May 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/168.

Lipovetsky:2020:BSS

[LC20] Stan Lipovetsky and Michael W. Conklin. Bayesian sensitivityspecificity and ROC analysis for finding key drivers. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1085.

Li:2017:JAS

[LCW^{+}17] Isaac Li, Yi-Hsin Chen, Yan Wang, Patricia Rodríguez de Gil, Thanh Pham, Diep Nguyen, Eun Sook Kim, and Jeffrey D. Kromrey. JMASM 47: ANOVA HOV: a SAS macro for testing homogeneity of variance in one-factor ANOVA models (SAS). Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/954.

LeBeau:2016:ISC

[LeB16] Brandon LeBeau. Impact of serial correlation misspecification with the linear mixed model. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/806.

Lecoutre:2007:ALC

[Lec07] Bruno Lecoutre. Another look at the confidence intervals for the noncentral T distribution. Journal of Modern Applied Statistical Methods, 6(1):107–116, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/286.

Lee:2011:FET

[Lee11a] Tze-San Lee. Fisher’s exact test for misclassified data. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/530.

Lee:2011:MPS

[Lee11b] Tze-San Lee. Matched-pair studies with misclassified ordinal data. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/524.

Lee:2012:PRM

[Lee12] Tze-San Lee. A Poisson regression model for female radium dial workers. Journal of Modern Applied Statistical Methods, 11(1): ??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/602.

Lee:2013:CCS

[Lee13] Tze-San Lee. Case-control studies with jointly misclassified exposure and confounding variables. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/670.

Lee:2015:BFC

[Lee15] Tzesan Lee. The Bayes factor for case-control studies with misclassified data. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/781.

Lehner:2015:EAA

[Leh15] Paul E. Lehner. Estimating the accuracy of automated classification systems using only expert ratings that are less accurate than the system. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/758.

Liu:2010:RSB

[LEKG10] Xian Liu, Charles C. Engel, Han Kang, and Kristie L. Gore. Reducing selection bias in analyzing longitudinal health data with high mortality rates. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/498.

Lesik:2011:NPA

[Les11] Sally A. Lesik. New perspectives in applying the regressiondiscontinuity design for program evaluation: a simulation analysis. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/534.

Lesik:2011:SES

[LF11] Sally A. Lesik and Carolyn R. Fallahi. Salary equity studies: an analysis of using the Blinder–Oaxaca decomposition to estimate differences in faculty salaries by gender. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/574.

Levin:2014:IRT

[LFG14] Joel R. Levin, John M. Ferron, and Boris S. Gafurov. Improved randomization tests for a class of single-case intervention designs. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/719.

Levin:2019:ITI

[LFG19] Joel R. Levin, John M. Ferron, and Boris S. Gafurov. An improved two independent-samples randomization test for singlecase AB-type intervention designs: a 20-year journey. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1015.

Lee:2007:BWD

[LFO07] Carl Lee, Felix Famoye, and Olugbenga Olumolade. Beta-Weibull distribution: Some properties and applications to censored data. Journal of Modern Applied Statistical Methods, 6(1):

??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/292.

Lix:2004:MCR

[LH04] Lisa M. Lix and Aynslie M. Hinds. Multivariate contrasts for repeated measures designs under assumption violations. Journal of Modern Applied Statistical Methods, 3(2):333–344, November 2004. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/142.

Li:2009:LBS

[LH09] Chin-Shang Li and Daniel L. Hunt. A linear B-spline threshold dose–response model with dose-specific response variation applied to developmental toxicity studies. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/448.

Lee:2013:TAN

[LH13] Tze-San Lee and Qin Hui. Testing the assumption of nondifferential misclassification in case-control studies. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/671.

Leach:2014:BPS

[LH14] Lesley F. Leach and Robin K. Henson. Bias and precision of the squared canonical correlation coefficient under nonnormal data condition. Journal of Modern Applied Statistical Methods, 13(1): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/694.

Lei:2004:ART

[LHB04] Xiaosheng Lei, Janet K. Holt, and T. Mark Beasley. Aligned rank tests as robust alternatives for testing interactions in multiple group repeated measures designs with heterogeneous covariances. Journal of Modern Applied Statistical Methods, 3(2): 462–475, November 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/157.

Liu:2011:UFM

[LHH11] Min Liu, Gregory R. Hancock, and Jeffrey R. Harring. Using finite mixture modeling to deal with systematic measurement error: a case study. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/539.

Li:2014:IFU

[LHM14] Ming Li, Jeffrey R. Harring, and George B. Macready. Investigating the feasibility of using Mplus in the estimation of growth mixture models. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/717.

Liu:2018:CMI

[LHO^{+}18] Benmei Liu, Erin Hennessy, April Oh, Laura A. Dwyer, and Linda Nebeling. Comparison of multiple imputation methods for categorical survey items with high missing rates: Application to the family life, activity, sun, health and eating (FLASHE) study. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/975.

Lipovetsky:2006:ECL

[Lip06] Stan Lipovetsky. Entropy criterion in logistic regression and Shapley value of predictors. Journal of Modern Applied Statistical Methods, 5(1):95–106, May 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/229.

Lipovetsky:2007:OLM

[Lip07] Stan Lipovetsky. Optimal Lp-metric for minimizing powered deviations in regression. Journal of Modern Applied Statistical Methods, 6(1):219–227, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/295.

Lipovetsky:2009:MRP

[Lip09] Stan Lipovetsky. Multiple regression in pair correlation solution. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/405.

Lipovetsky:2010:NPB

[Lip10] Stan Lipovetsky. Nonlinear parameterization in bi-criteria sample balancing. Journal of Modern Applied Statistical Methods, 9 (1):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/477.

Lipovetsky:2012:RSL

[Lip12] Stan Lipovetsky. Regression split by levels of the dependent variable. Journal of Modern Applied Statistical Methods, 11(2): ??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/612.

Lipovetsky:2013:HGB

[Lip13a] Stan Lipovetsky. How good is best? Multivariate case of Ehrenberg–Weisberg analysis of residual errors in competing regressions. Journal of Modern Applied Statistical Methods, 12(2): ??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/673.

Lipovetsky:2013:PTS

[Lip13b] Stan Lipovetsky. Priorities in Thurstone scaling and steady-state probabilities in Markov stochastic modeling. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/643.

Lipovetsky:2016:GSV

[Lip16] Stan Lipovetsky. Generalized singular value decomposition with additive components. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/814.

Lipovetsky:2017:FAL

[Lip17a] Stan Lipovetsky. Factor analysis by limited scales: Which factors to analyze? Journal of Modern Applied Statistical Methods, 16 (1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/894.

Lipovetsky:2017:PPC

[Lip17b] Stan Lipovetsky. Prediction of percent change in linear regression by correlated variables. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/941.

Lipovetsky:2018:RRC

[Lip18] Stan Lipovetsky. Regressions regularized by correlations. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/968.

Lipovetsky:2019:RMP

[Lip19] Stan Lipovetsky. Regression modeling and prediction by individual observations versus frequency. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1018.

Lestari:2024:FTE

[LIT24] Suci Lestari, Anna Islamiyati, and Georgina Maria Tinungki. Forecasting with triple exponential smoothing on data of the number of passengers departing domestic flights at sultan hasanuddin international airport. Journal of Modern Applied Statistical Methods, 23(2):??, January 2024. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1217.

Liu:2009:ORA

[Liu09] Xing Liu. Ordinal regression analysis: Fitting the proportional odds model using Stata, SAS and SPSS. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/457.

Liu:2014:FSL

[Liu14] Xing Liu. Fitting stereotype logistic regression models for ordinal response variables in educational research (Stata). Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/748.

Liu:2012:ORA

[LK12] Xing Liu and Hari Koirala. Ordinal regression analysis: Using generalized ordinal logistic regression models to estimate educational data. Journal of Modern Applied Statistical Methods, 11 (1):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/603.

Liu:2013:FPO

[LK13] Xing Liu and Hari Koirala. Fitting proportional odds models to educational data with complex sampling designs in ordinal logistic regression. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/657.

Liu:2019:IPP

[LKW^{+}19] Yan Liu, Chanmin Kim, Amrey D. Wu, Paul Gustafson, and Edward Kroc. Investigating the performance of propensity score approaches for differential item functioning analysis. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1028.

Lix:2007:CPA

[LL07] Lisa M. Lix and Anita M. Lloyd. A comparison of procedures for the analysis of multivariate repeated measurements. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/313.

Lim:2016:CUT

[LL16] Wooi K. Lim and Alice W. Lim. A comparison of usual t-test statistic and modified t-test statistics on skewed distribution functions. Journal of Modern Applied Statistical Methods, 15 (2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/841.

Liao:2017:OIA

[LLB17] Hongjing Liao, Yanju Li, and Gordon P. Brooks. Outlier impact and accommodation on power. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/896.

Li:2003:CVT

[LLHT03] Chin-Shang Li, Hua Liang, Ying-Hen Hsieh, and Shiing-Jer Twu. Comparison of viral trajectories in aids studies by using nonparametric mixed-effects models. Journal of Modern Applied Statistical Methods, 2(2):443–450, November 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/101.

Lipovetsky:2015:MPC

[LM15] Stan Lipovetsky and Igor Mandel. Modeling probability of causal and random impacts. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/761.

Lakshmi:2017:PLI

[LM17] B. Vijaya Lakshmi and V. Mohan. Plant leaf image detection method using a midpoint circle algorithm for shape-based feature extraction. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/907.

Lipovetsky:2019:ACF

[LM19] Stan Lipovetsky and Vladimir Manewitsch. Analytical closedform solution for general factor with many variables. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1019.

Lyles:2009:CAC

[LMME09] Robert H. Lyles, Reneé H. Moore, Amita K. Manatunga, and Kirk A. Easley. Covariate-adjusted constrained Bayes predictions of random intercepts and slopes. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/401.

Ling:2020:HAM

[LMRM^{+}20] Albee Ling, Maria Montez-Rath, Maya Mathur, Kris Kapphahn, and Manisha Desai. How to apply multiple imputation in propensity score matching with partially observed confounders: a simulation study and practical recommendations. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1072.

Lorenz:2015:CSU

[LMS15] Akiva J. Lorenz, Barry S. Markman, and Shlomo Sawilowsky. Caution for software use of new statistical methods (r). Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/785.

Lukman:2015:TSR

[LOA15] Adewale Folaranmi Lukman, Oyedeji Isola Osowole, and Kayode Ayinde. Two stage robust ridge method in a linear regression model. Journal of Modern Applied Statistical Methods, 14(2): ??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/773.

Lloyd:2013:JMI

[LOCMS13] Jennifer E. V. Lloyd, Jelena Obradović, Richard M. Carpiano, and Frosso Motti-Stefanidi. JMASM 32: Multiple imputation of missing multilevel, longitudinal data: a case when practical considerations trump best practices? Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/660.

Liu:2011:ORA

[LOK11] Xing Liu, Ann A. O’Connell, and Hari Koirala. Ordinal regression analysis: Predicting mathematics proficiency using the continuation ratio model. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/561.

Lorah:2019:EMM

[Lor19] Julie Lorah. Estimating a multilevel model with complex survey data: Demonstration using TIMSS. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1057.

Lovric:2019:CBS

[Lov19a] Miodrag M. Lovric. Conflicts in Bayesian statistics between inference based on credible intervals and Bayes factors. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1014.

Lovric:2019:ANR

[Lov19b] Miodrag M. Lovric. On the authentic notion, relevance, and solution of the Jeffreys–Lindley paradox in the zettabyte era. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1013.

Liou:2006:UHA

[LP06] Show-Mann Liou and Chao-Ying Joanne Peng. The use of hierarchical ANCOVA in curriculum studies. Journal of Modern Applied Statistical Methods, 5(1):230–247, May 2006. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/239.

Levin:2003:TIS

[LR03] Joel R. Levin and Daniel H. Robinson. The trouble with interpreting statistically nonsignificant effect sizes in single-study investigations. Journal of Modern Applied Statistical Methods, 2 (1):??, ???? 2003. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/62.

Lone:2020:ISS

[LRT20] Showkat Ahmad Lone, Ahmadur Rahman, and Tanveer A. Tarray. Inference for step-stress partially accelerated life test model with an adaptive Type-I progressively hybrid censored data. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1080.

Long:2005:RTT

[LS05] Michael C. Long and Ping Sa. Right-tailed testing of variance for non-normal distributions. Journal of Modern Applied Statistical Methods, 4(1):187–213, May 2005. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/181.

Lachlan:2006:CTE

[LS06] Kenneth Lachlan and Patric R. Spence. Corrections for Type I error in social science research: a disconnect between theory and practice. Journal of Modern Applied Statistical Methods, 5(2): 490–494, November 2006. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/266.

Laloe:2013:XAA

[LS13] Thomas Laloë and Rémi Servien. The X-Alter algorithm: a parameter-free method of unsupervised clustering. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/645.

Leavline:2016:AMS

[LS16a] E. Jebamalar Leavline and Sutha Shunmugam. Analysis and modeling of statistical properties of FMDFB subband coefficients. Journal of Modern Applied Statistical Methods, 15(1): ??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/826.

Levin:2016:CHC

[LS16b] Joel R. Levin and Ronald C. Serlin. The composite hypothesis contrast procedure: a novel sequential multiple-comparison approach. Journal of Modern Applied Statistical Methods, 15(1): ??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/787.

Lowenstein:2017:RPC

[LS17] Linda C. Lowenstein and Shlomo S. Sawilowsky. Robustness and power comparison of the Mood–Westenberg and Siegel–Tukey tests. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/893.

Landsittel:2002:NDL

[LSAA02] Douglas Landsittel, Harshinder Singh, Vincent C. Arena, and Stewart J. Anderson. Null distribution of the likelihood ratio statistic for feed-forward neural networks. Journal of Modern Applied Statistical Methods, 1(2):333–342, November 2002. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/39.

Li:2007:SBL

[LT07] Chin-Shang Li and Wanzhu Tu. A spline-based lack-of-fit test for independent variable effect. Journal of Modern Applied Statistical Methods, 6(1):239–247, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/297.

Lu:2016:ADS

[Lu16] Ying Lu. Analyzing different sampling designs (SAS). Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/829.

Lui:2006:IER

[Lui06] Kung-Jong Lui. Interval estimation of risk difference in simple compliance randomized trials. Journal of Modern Applied Statistical Methods, 5(2):395–407, November 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/257.

Lui:2007:RIR

[Lui07] Kung-Jong Lui. Reply (to Ian R. White). Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/340. See [Whi07].

Lui:2010:NHT

[Lui10] Kung-Jong Lui. Notes on hypothesis testing under a single-stage design in phase II trial. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/496.

Liu:2005:SSC

[LW05] Honghu Liu and Tongtong Wu. Sample size calculation and power analysis of time-averaged difference. Journal of Modern Applied Statistical Methods, 4(2):434–445, November 2005. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/200.

Liu:2016:CIU

[LWZK16] Ruiqi Liu, Dongfeng Wu, Xiang Zhang, and Seongho Kim. Compound identification using penalized linear regression on metabolomics. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/805.

Liang:2006:PSL

[LX06] Hua Liang and Yuanhui Xiao. Penalized splines for longitudinal data with an application in AIDS studies. Journal of Modern Applied Statistical Methods, 5(1):130–139, May 2006. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/232.

Lloyd:2007:NPD

[LZ07] Jennifer E. V. Lloyd and Bruno D. Zumbo. The non-parametric difference score: a workable solution for analyzing two-wave change when the measures themselves change across waves. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/315.

Liu:2014:RIP

[LZW14] Yan Liu, Bruno D. Zumbo, and Amery D. Wu. Relative importance of predictors in multilevel modeling. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/688.

Mascha:2006:PBE

[MA06] Edward J. Mascha and Jeffrey M. Albert. Properties of bound estimators on treatment effect heterogeneity for binary outcomes. Journal of Modern Applied Statistical Methods, 5(1):181–193, May 2006. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/236.

Makinde:2017:CDF

[MA17] Olusola Samuel Makinde and Adeyinka Damilare Adewumi. A comparison of depth functions in maximal depth classification rules. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/903.

Mudiasir:2018:WVG

[MA18] Sofi Mudiasir and S. P. Ahmad. Weighted version of generalized inverse Weibull distribution. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1008.

Malik:2024:GIK

[MA24] Aliya Syed Malik and S. P. Ahmad. Generalized inverted Kumaraswamy–Rayleigh distribution: Properties and application. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1182.

Madan:2003:SSA

[MADT03] Kailash C. Madan, Walid Abu-Dayyeh, and Firas Tayyan. Steady state analysis of an M/D/2 queue with Bernoulli schedule server vacations. Journal of Modern Applied Statistical Methods, 2(1):202–209, May 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/79.

Makinde:2017:MRO

[Mak17] Olusola Samuel Makinde. Multivariate rank outlyingness and correlation effects. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/895.

Malkin:2002:UET

[Mal02] Robert A. Malkin. An unconditional exact test for small samples matched binary pairs. Journal of Modern Applied Statistical Methods, 1(1):69–73, May 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/10.

Mantalos:2005:TSS

[Man05] Panagiotis Mantalos. Two sides of the same coin: Bootstrapping the restricted vs. unrestricted model. Journal of Modern Applied Statistical Methods, 4(1):35–42, May 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/164.

Manly:2013:RTV

[Man13] Bryan Manly. Randomization test p-values versus significance levels. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/636.

Muhammad:2008:AEH

[MAP08] Faqir Muhammad, Muhammad Aslam, and G. R. Pasha. Adaptive estimation of heteroscedastic linear regression model using probability weighted moments. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/384.

Marsaglia:2003:RNG

[Mar03] George Marsaglia. Random number generators. Journal of Modern Applied Statistical Methods, 2(1):2–13, May 2003. CODEN ???? ISSN 1538-9472. URL http://stat.fsu.

edu/pub/diehard/;   http://www.csis.hku.hk/~diehard/
;   https://jmasm.com/index.php/jmasm/article/view/57.

Mardaneh:2012:SME

[Mar12] Karim K. Mardaneh. Small-to-medium enterprises and economic growth: a comparative study of clustering techniques. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/626.

Mondragon:2005:CNR

[MB05] Paul Fredrick Mondragon and Brian Borchers. A comparison of nonlinear regression codes. Journal of Modern Applied Statistical Methods, 4(1):343–351, May 2005. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/189.

Musial:2002:LFD

[MBS02] Joseph L. Musial, Patrick D. Bridge, and Nicol R. Shamey. A longitudinal follow-up of discrete mass at zero with gap. Journal of Modern Applied Statistical Methods, 1(2):411–419, November 2002. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/47.

Maji:2016:EUE

[MBS16] Reba Maji, Arnab Bandyopadhyay, and G. N. Singh. Efficient and unbiased estimation procedure of population mean in twophase sampling. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/846.

Maggio:2017:JWB

[MBS17] Saverpierre Maggio, Gokul Bhandari, and Shlomo S. Sawilowsky. JMASM 50: a Web-based shiny application for conducting a two dependent samples maximum test (R). Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/957.

Mulvenon:2007:ANP

[MBWZ07] Sean W. Mulvenon, M. Austin Betz, Kening Wang, and Bruno D. Zumbo. Application of a new procedure for power analysis and comparison of the adjusted univariate and multivariate tests in repeated measures designs. Journal of Modern Applied Statistical Methods, 6(1):36–52, May 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/281.

Mends-Cole:2007:PCI

[MC07] S. Jonathan Mends-Cole. Probability coverage and interval length for Welch’s and Yuen’s techniques: Shift in location, change in scale, and (un)Equal sizes. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/316.

Mends-Cole:2008:PCI

[MC08] S. Jonathan Mends-Cole. Probability of coverage and interval length for two-group techniques assessing the median and trimmed mean. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/352.

Mann:2011:WLD

[MC11] Sara Mann and James Chowhan. Weighting large datasets with complex sampling designs: Choosing the appropriate variance estimation method. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/527.

Mills:2009:HRB

[MCL09] Laura Mills, Robert A. Cribbie, and Wei-Ming Luh. A heteroscedastic, rank-based approach for analyzing 2 × 2 independent groups designs. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/426.

McCarthy:2003:RAF

[MDS03] Joseph McCarthy, Robert DiSario, and Hakan Saraoglu. A recursive algorithm for fractionally differencing long data series.

Journal of Modern Applied Statistical Methods, 2(1):272–278, May 2003. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/84.

Mecklin:2003:CET

[Mec03] Christopher J. Mecklin. A comparison of equivalence testing in combination with hypothesis testing and effect sizes. Journal of Modern Applied Statistical Methods, 2(2):329–340, November 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/90.

Martinez-Espineira:2008:ERB

[MEH08] Roberto Martínez-Espiñeira and Joseph M. Hilbe. Effect on recreation benefit estimates from correcting for on-site sampling biases and heterogeneous trip overdispersion in count data recreation demand models (STATA). Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/368.

McCombs:2009:ETD

[MEP09] Edward L. McCombs, Matthew E. Elam, and David B. Prat. Estimating task duration in PERT using the Weibull probability distribution. Journal of Modern Applied Statistical Methods, 8 (1):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/420.

Mahmoud:2016:BEB

[MESSE16] Mohamed A. W. Mahmoud, Rashad M. El-Sagheer, Ahmed A. Soliman, and Ahmed H. Abd Ellah. Bayesian estimation of P [Y < X] based on record values from the Lomax distribution and MCMC technique. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/810.

Meyvisch:2016:EAF

[Mey16] Paul Meyvisch. Evaluation of the addition of Firth’s penalty term to the Bradley–Terry likelihood. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/850.

Makinde:2011:IOA

[MF11] Olusola Samuel Makinde and Olusoga Akin Fasoranbaku. Identification of optimal autoregressive integrated moving average model on temperature data. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/581.

Michaeloudis:2011:MLS

[Mic11] Androulla Michaeloudis. Maximum likelihood solution for the linear structural relationship with three parameters known. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/545.

Mickelson:2013:MCS

[Mic13] William T. Mickelson. A Monte Carlo simulation of the robust rank-order test under various population symmetry conditions. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/638.

Min:2013:OLR

[Min13] Hosik Min. Ordered logit regression modeling of the self-rated health in Hawai‘i, with comparisons to the OLS model. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/682.

Mostajeran:2018:ESN

[MIN18] Abdolrasoul Mostajeran, Nasrolah Iranpanah, and Rassoul Noorossana. An explanatory study on the non-parametric multivariate T2 control chart. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/983.

Mir:2008:SBG

[Mir08] Khurshid Ahmad Mir. Size-biased generalized negative binomial distribution. Journal of Modern Applied Statistical Methods, 7 (2):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/378.

Mahdavi:2017:EWE

[MJ17] Abbas Mahdavi and Leila Jabari. An extended weighted exponential distribution. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/898.

Mirzaei:2020:NGF

[MJ20] Shahryar Mirzaei and S. M. A. Jahanshahi. A new goodness of fit measure based on income inequality curves. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1090.

Moses:2005:TAT

[MK05] Tim Moses and Alan Klockars. Testing for aptitude-treatment interactions in analysis of covariance and randomized block designs under assumption violations. Journal of Modern Applied Statistical Methods, 4(2):460–472, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/202.

Moses:2008:ESA

[MK08] Tim Moses and Alan Klockars. An evaluation of standard, alternative, and robust slope test strategies. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/346.

Moses:2010:IDG

[MK10] Tim Moses and Alan Klockars. The influence of data generation on simulation study results: Tests of mean differences. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/462.

Mustafa:2023:LBE

[MK23] Abdelfattah Mustafa and M. I. Khan. Length biased extended Rayleigh distribution and its applications. Journal of Modern Applied Statistical Methods, 22(2):??, November 2023. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1199.

Malik:2023:IRE

[MKA23a] Sachin Malik, Kanika, and Atul. An improvement in regression estimator through exponential estimator using two auxiliary variables. Journal of Modern Applied Statistical Methods, 22(1):??, February 2023. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1130.

Malik:2023:MRE

[MKA23b] Sachin Malik, Kanika, and Atul. A modified ratio estimator using two auxiliary variables. Journal of Modern Applied Statistical Methods, 22(2):??, ???? 2023. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1140.

Maansson:2019:NLT

[MKS19] Kristofer Månsson, B. M. Golam Kibria, and Ghazi Shukur. A new Liu type of estimator for the restricted SUR estimator. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1024.

Montes:2003:CPB

[ML03] Guillermo Montes and Bohdan S. Lotyczewski. Correcting publication bias in meta-analysis: a truncation approach. Journal of Modern Applied Statistical Methods, 2(2):433–442, November 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/100.

Manikandan:2024:MLB

[ML24] Manikandan and Wang Hong Ling. Machine learning-based book recommendation systems: a comparative study of CFNN and KNN algorithms. Journal of Modern Applied Statistical Methods, 23(2):63–80, December 2024. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1382.

Munoz:2010:AMI

[MLS10] Breda Munoz, Virginia M. Lesser, and Ruben A. Smith. Applying multiple imputation with geostatistical models to account for item nonresponse in environmental data. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/485.

Mahmood:2019:CBM

[MMH19] Ehab A. Mahmood, Habshah Midi, and Abdul Ghapor Hussin. The comparison between maximum weighted and trimmed likelihood estimator of the simple circular regression model. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1067.

Mukherjee:2007:ERD

[MMM07] Ajit Mukherjee, Ajit Mathur, and Rakesh Mittal. Estimation of risk for developing cardiac problem in patients of type 2 diabetes as obtained by the technique of density estimation. Journal of Modern Applied Statistical Methods, 6(1):320–323, May 2007. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/304.

Midorikawa:2008:ADP

[MMS08] Shuichi Midorikawa, Etsuo Miyaoka, and Bruce Smith. Application of dynamic Poisson models to Japanese cancer mortality data. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/391.

Mak:2005:EPV

[MN05] T. K. Mak and Fassil Nebebe. Estimation of process variances in robust parameter designs. Journal of Modern Applied Statistical Methods, 4(2):394–401, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/196.

Mak:2017:ARP

[MN17] Tak K. Mak and Fassil Nebebe. Analysis of robust parameter designs. Journal of Modern Applied Statistical Methods, 16(1): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/892.

Mbaeyi:2021:DAS

[MN21] George C. Mbaeyi and Chijioke J. Nweke. On discriminant analysis with some bivariate exponential distributions. Journal of Modern Applied Statistical Methods, 20(2):??, July 2021. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1117.

Meek:2007:CVW

[MOD07] Gary E. Meek, Ceyhun Ozgur, and Kenneth Dunning. Comparison of the t vs. Wilcoxon signed-rank test for Likert scale data and small samples. Journal of Modern Applied Statistical Methods, 6(1):91–106, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/285.

Moses:2012:UDL

[Mos12] Tim Moses. Underlying distributions in loglinear models of discrete data. Journal of Modern Applied Statistical Methods, 11 (1):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/584.

McCandless:2016:BES

[MPC^{+}16] Lawrence C. McCandless, Michelle L. Patterson, Lauren B. Currie, Akm Moniruzzaman, and Julian M. Somers. Bayesian estimation of the size of a street-dwelling homeless population. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/800.

Morris:2011:PTC

[MPS11] Tracy L. Morris, Mark E. Payton, and Stephanie A. Santorico. A permutation test for compound symmetry with application to gene expression data. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/556.

Montez-Rath:2017:GGR

[MRKM^{+}17] Maria E. Montez-Rath, Kristopher Kapphahn, Maya B. Mathur, Aya A. Mitani, David J. Hendry, and Manisha Desai. Guidelines for generating right-censored outcomes from a Cox model extended to accommodate time-varying covariates. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/887.

Mahmood:2017:DOU

[MRMH17] Ehab A. Mahmood, Sohel Rana, Habshah Midi, and Abdul Ghapor Hussin. Detection of outliers in univariate circular data using robust circular distance. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/945.

Miller:2003:IMC

[MS03] Laura K. Miller and Ping Sa. Improved multiple comparisons with the best in response surface methodology. Journal of Modern Applied Statistical Methods, 2(1):247–255, May 2003. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/81.

Mutan:2009:MCC

[MŞ09] Oya Can Mutan and Birdal Şenoğlu. A Monte Carlo comparison of regression estimators when the error distribution is long-tailed symmetric. Journal of Modern Applied Statistical Methods, 8(1): ??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/408.

Maggio:2014:JTD

[MS14] Saverpierre Maggio and Shlomo Sawilowsky. JMASM 33: a two dependent samples maximum test calculator: Excel. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/718.

Mundform:2011:NRR

[MSK^{+}11] Daniel J. Mundform, Jay Schaffer, Myoung-Jin Kim, Dale Shaw, and Ampai Thongteeraparp. Number of replications required in Monte Carlo simulation studies: a synthesis of four studies. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/520.

Mantalos:2007:EGG

[MSS07] Panagiotis Mantalos, Ghazi Shukur, and Pär Sjölander. The effect of GARCH (1,1) on the Granger causality test in stable VAR models. Journal of Modern Applied Statistical Methods, 6 (2):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/320.

Mishra:2019:RES

[MSS19] Uttama Mishra, S. Siddiqui, and J. R. Singh. Robustness of the EWMA sampling plan to non-normality. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1061.

Maurya:2020:NRS

[MSS20a] Sandeep Kumar Maurya, Sanjay K. Singh, and Umesh Singh. A new right-skewed upside down bathtub shaped heavy-tailed distribution and its applications. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1095.

Maurya:2020:CSN

[MSS20b] Sandeep Kumar Maurya, Sanjay Kumar Singh, and Umesh Singh. Comparative study of new and traditional estimators of a new lifetime model. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1088.

Mehta:2024:RPT

[MT24] Priya Mehta and Rajesh Tailor. Ratio and product type estimators for ratio of two population means in case of poststratification. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1168.

Mugdadi:2012:WHD

[Mug12] Abdel-Razzaq Mugdadi. The weighted Hellinger distance for kernel distribution estimator of function of observations. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/606.

Menon:2007:COH

[MW07] Mahesh Menon and Todd S. Woodward. A comparison of onehigh-threshold and two-high-threshold multinomial models of source monitoring. Journal of Modern Applied Statistical Methods, 6(1):279–290, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/301.

Mammadov:2006:SMA

[MYY^{+}06] Mammadagha Mammadov, Berna Yazici, Šenay Yolaçan, Atilla Aslanargun, Ali Fuat Yüzer, and Embiya Ağaoğlu. Statistical methods and artificial neural networks. Journal of Modern Applied Statistical Methods, 5(2):495–512, November 2006. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/267.

Messaadia:2017:AGL

[MZ17] Hamouda Messaadia and Halim Zeghdoudi. Around gamma Lindley distribution. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/946.

Naqash:2016:BAG

[NAA16] Saima Naqash, S. P. Ahmad, and Aquil Ahmed. Bayesian analysis of generalized exponential distribution. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/871.

Nkemnole:2013:PEC

[NAK13] E. B. Nkemnole, O. Abass, and R. A. Kasumu. Parameter estimation of a class of hidden Markov model with diagnostics. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/652.

Nanna:2002:HTV

[Nan02] Michael J. Nanna. Hotelling’s T2 vs. the rank transform with real Likert data. Journal of Modern Applied Statistical Methods, 1(1):83–99, May 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/12.

Nandram:2005:BSA

[Nan05] Balgobin Nandram. A Bayesian subset analysis of sensory evaluation data. Journal of Modern Applied Statistical Methods, 4(2): 482–499, November 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/204.

Nandakishore:2010:BAC

[Nan10] L. V. Nandakishore. Bayesian analysis for component manufacturing processes. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/516.

Navaei:2007:LDT

[Nav07] Leader Navaei. Large deviations techniques for error exponents to multiple hypothesis LAO testing. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/321.

Nofal:2018:TEA

[NAY^{+}18] Zohdy M. Nofal, Ahmed Z. Afify, Haitham M. Yousof, Daniele Cristina Tita Granzotto, and Francisco Louzada. The transmuted exponentiated additive Weibull distribution: Properties and applications. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/961.

Nasiri:2016:BEP

[NE16] Parviz Nasiri and Hassan Esfandyarifar. E-Bayesian estimation of the parameter of the logarithmic series distribution. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/870.

Neel:2004:NGF

[Nee04] John H. Neel. A new goodness-of-fit test for item response theory. Journal of Modern Applied Statistical Methods, 3(2):581– 593, November 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/152.

Neuhaeuser:2009:MTA

[Neu09] Markus Neuhäeuser. A maximum test for the analysis of ordered categorical data. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/450.

Neuhaeuser:2010:ETB

[Neu10] Markus Neuhäeuser. An equivalence test based on n and p. Journal of Modern Applied Statistical Methods, 9(1):??, ????

2010. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/488.

Ng:2006:IER

[Ng06] Vee Ming Ng. Inference for P (Y < X) for exponential and related distributions. Journal of Modern Applied Statistical Methods, 5(2):323–327, November 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/248.

Neuhauser:2006:MTA

[NH06] Markus Neuhäuser and Ludwig A. Hothorn. Maximum tests are adaptive permutation tests. Journal of Modern Applied Statistical Methods, 5(2):317–322, November 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/247.

Nasiri:2012:SIL

[NH12] Parviz Nasiri and Saman Hosseini. Statistical inferences for Lomax distribution based on record values (Bayesian and classical). Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/597.

Nwogu:2016:STS

[NIN16] Eleazar Chukwunenye Nwogu, Iheanyi Sylvester Iwueze, and Valentine Uchenna Nlebedim. Some tests for seasonality in time series data. Journal of Modern Applied Statistical Methods, 15 (2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/857.

Nguyen:2018:COT

[NKC18] Diep Nguyen, Eun Sook Kim, and Yi-Hsin Chen. Can one test fit all? Responses to the article “Striving for Simple but Effective Advice for Comparing the Central Tendency of Two Populations” (Ruxton & Neuhäuser, 2018). Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/993. See [RN18].

Nguyen:2016:PTT

[NKdG^{+}16] Diep T. Nguyen, Eun Sook Kim, Patricia Rodriguez de Gil, Anh Kellermann, and Yi-Hsin Chen. Parametric tests for two population means under normal and non-normal distributions. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/794.

Nguyen:2019:ECT

[NKW^{+}19] Diep Nguyen, Eunsook Kim, Yan Wang, Thanh Vinh Pham, and Yi-Hsin Chen. Empirical comparison of tests for one-factor ANOVA under heterogeneity and non-normality: a Monte Carlo study. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1046.

Ng:2010:NBE

[NLQ10] Set Foong Ng, Heng Chin Low, and Soon Hoe Quah. A new biased estimator derived from principal component regression estimator. Journal of Modern Applied Statistical Methods, 9(1): ??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/480.

Novak:2017:SEA

[NM17] Rosilei S. Novak and Jair M. Marques. Study evaluating the alterations caused in an exploratory factor analysis when multivariate normal data is dichotomized. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/951.

Novak:2019:SVD

[NM19] Rosilei S. Novak and Jair M. Marques. A study verifying the dimensioning of a multivariate dichotomized sample in exploratory factor analysis. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1041.

Nakonezny:2005:EER

[NR05] Paul A. Nakonezny and Joseph Lee Rodgers. An empirical evaluation of the retrospective pretest: Are there advantages to looking back? Journal of Modern Applied Statistical Methods, 4(1): ??, ???? 2005. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/184.

Nahm:2002:QSM

[NRSGRN02] Abraham Y. Nahm, S. Subba Rao, Luis E. Solis-Galvan, and T. S. Ragu-Nathan. The Q-sort method: Assessing reliability and construct validity of questionnaire items at a pre-testing stage. Journal of Modern Applied Statistical Methods, 1(1):114– 125, May 2002. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/15.

Nakonezny:2007:JHM

[NS07] Paul A. Nakonezny and Robert D. Shull. JMASM 26: Hettmansperger and Mckean linear model aligned rank test for the single covariate and one-way ANCOVA case (SAS). Journal of Modern Applied Statistical Methods, 6(1):??, ???? 2007. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/307.

Nagineni:2024:LNV

[NSJ^{+}24] Srinivas Nagineni, Nandhinidevi S, Balaraju J, G. R. Sakthidharan, M. Saravana Karthikeyan, and Vinayak Musale. Leveraging nanosensor and vision transformer for robust anomaly detection in autonomous vehicles. Journal of Modern Applied Statistical Methods, 23(2):15–32, December 2024. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1380.

Nurwiani:2014:RRC

[NSSO14] Nurwiani, S. Sunaryo, Setiawan, and B. W. Otok. Ridge regression in calibration models with symmetric padding extension — Daubechies wavelet transform preprocessing. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/702.

Nurwiani:2025:IST

[NT25] Nurwiani and Lia Budi Tristanti. Independent samples t test and the Mann–Whitney–Wilcoxon test to know the effect of the drill method on mathematics learning outcomes. Journal of Modern Applied Statistical Methods, 24(1):??, ???? 2025. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1331.

Ng:2009:LRM

[NW09] Marie Ng and Rand R. Wilcox. Level robust methods based on the least squares regression estimator. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/431.

Ng:2010:SSE

[NW10] Marie Ng and Rand R. Wilcox. The small-sample efficiency of some recently proposed multivariate measures of location. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/463.

Nascimento:2019:DCC

[NXFN19] Diego Nascimento, Cleber Xavier, Israel Felipe, and Francisco Louzada Neto. Dynamic conditional correlation GARCH: a multivariate time series novel using a Bayesian approach. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/1021.

Ou:2011:RIR

[OA11] Juchi Ou and Jeffrey M. Albert. Robust inference for regression with spatially correlated errors. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/557.

Ozdemir:2010:AFM

[OAA10] Ozer Ozdemir, Atilla Aslanargun, and Senay Asma. ANN forecasting models for ISE national-100 index. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/514.

Otieno:2003:MEW

[OAC03] B. Sango Otieno and Christine M. Anderson-Cook. A more efficient way of obtaining a unique median estimate for circular data. Journal of Modern Applied Statistical Methods, 2(1):168– 176, May 2003. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/75.

Ozgul:2017:NEB

[ÖÇ17] Nilgün Özgül and Hülya Çıngı. A new estimator based on auxiliary information through quantitative randomized response techniques. Journal of Modern Applied Statistical Methods, 16(1): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/902.

Oyamakin:2013:CEH

[OCB13] S. O. Oyamakin, A. U. Chukwu, and T. A. Bamiduro. On comparison of exponential and hyperbolic exponential growth models in height/diameter increment of PINES (Pinus caribaea). Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/683.

OConnell:2007:MLO

[OD07] Ann A. O’Connell and Heather Levitt Doucette. Modeling longitudinal ordinal response variables for educational data. Journal of Modern Applied Statistical Methods, 6(1):304–319, May 2007. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/303.

Odiase:2012:ESP

[Odi12] Justice I. Odiase. Empirical sampling from permutation space with unique patterns. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/592.

Ozgur:2017:SSS

[ODLR17] Ceyhun Ozgur, Min Dou, Yang Li, and Grace Rogers. Selection of statistical software for data scientists and teachers. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/916.

Ozkaya:2012:EMC

[OE12] Guven Ozkaya and Ilker Ercan. Examining multiple comparison procedures according to error rate, power type and false discovery rate. Journal of Modern Applied Statistical Methods, 11(2): ??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/615.

Oxenyuk:2017:DFV

[OGKH17] Victoria Oxenyuk, Sneh Gulati, B. M. Golam Kibria, and Shahid Hamid. Distribution fits for various parameters in the Florida Public Hurricane Loss Model. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/908.

Omer:2020:TDC

[OHQ^{+}20] Talha Omer, Zawar Hussain, Muhammad Qasim, Said Farooq Shah, and Akbar Ali Khan. Two different classes of shrinkage estimators for the scale parameter of the Rayleigh distribution. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1097.

Oner:2017:JCS

[ÖK17] Metin Öner and İpek Deveci Kocakoç. JMASM 49: a compilation of some popular goodness of fit tests for normal distribution: Their algorithms and MATLAB codes (MATLAB). Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/956.

Othman:2002:TS

[OKW^{+}02] Abdul R. Othman, H. J. Keselman, Rand R. Wilcox, Katherine Fradette, and A. R. Padmanabhan. A test of symmetry. Journal of Modern Applied Statistical Methods, 1(2):310–315, November 2002. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/36.

Onwuegbuzie:2003:SSE

[OL03] Anthony J. Onwuegbuzie and Joel R. Levin. Without supporting statistical evidence, where would reported measures of substantive importance lead? To no good effect. Journal of Modern Applied Statistical Methods, 2(1):133–151, May 2003. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/72.

OConnell:2011:MDP

[OL11] Ann A. O’Connell and Xing Liu. Model diagnostics for proportional and partial proportional odds models. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/531.

Olanrewaju:2020:PLE

[Ola20] Rasaki Olawale Olanrewaju. Penalized likelihood estimation of gamma distributed response variable via corrected solution of regression coefficients. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1096.

Ogundele:2016:JAA

[OMN16] Suraju Olaniyi Ogundele, J. I. Mbegbu, and C. R. Nwosu. JMASM 42: an alternative algorithm and programming implementation for least absolute deviation estimator of the linear regression models (r). Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/880.

Oyeka:2014:PCR

[ON14] I. C. A. Oyeka and C. C. Nnanatu. Pairwise comparison in repeated measures. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/725.

Odiase:2005:JAG

[OO05] Justice I. Odiase and Sunday M. Ogbonmwan. JMASM 18: an algorithm for generating unconditional exact permutation distribution for a two-sample experiment. Journal of Modern Applied Statistical Methods, 4(1):319–332, May 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/192.

Ossai:2018:EMV

[OO18] Emmanuel Ogochukwu Ossai and Abimibola Victoria Oladugba. The estimation of missing values in rectangular lattice designs. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1005.

Ogunde:2024:TIT

[OON24] Adebisi A. Ogunde, Innocent O. Oseghale, and Okechukwu B. Nwanyibuife. The Type II Topp–Leone inverse power Lomax distribution with simulation and applications. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1176.

Opdyke:2003:FPT

[Opd03] J. D. Opdyke. Fast permutation tests that maximize power under conventional Monte Carlo sampling for pairwise and multiple comparisons. Journal of Modern Applied Statistical Methods, 2 (1):27–49, May 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/65.

Opdyke:2005:SPN

[Opd05] J. D. Opdyke. A single, powerful, nonparametric statistic for continuous-data telecommunications parity testing. Journal of Modern Applied Statistical Methods, 4(2):372–393, November 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/195.

Ocana-Rilola:2002:TME

[OR02] Ricardo Ocana-Rilola. Two methods to estimate homogeneous Markov processes. Journal of Modern Applied Statistical Methods, 1(1):??, ???? 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/17.

Ocana-Riola:2003:HMP

[ORSCMG03] Ricardo Ocaña-Riola, Emilio Sanchez-Cantalejo, and Carmen Martinez-Garcia. Homogeneous Markov processes for breast cancer analysis. Journal of Modern Applied Statistical Methods, 2 (1):210–217, May 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/80.

Odumade:2010:UTV

[OS10] Oluseun Odumade and Sarjinder Singh. Use of two variables having common mean to improve the Bar-Lev, Bobovitch and Boukai randomized response model. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/499.

Otsu:2005:TNA

[Ots05] Taisuke Otsu. Testing normality against the Laplace distribution. Journal of Modern Applied Statistical Methods, 4(2):402– 407, November 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/197.

Olaniran:2017:BHT

[OY17] Oyebayo Ridwan Olaniran and Waheed Babatunde Yahya. Bayesian hypothesis testing of two normal samples using bootstrap prior technique. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/952.

Oyamakin:2011:HDR

[Oya11] S. O. Oyamakin. Height-diameter relationship in tree modeling using simultaneous equation techniques in correlated normal deviates. Journal of Modern Applied Statistical Methods, 10(2): ??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/576.

Othman:2012:RML

[OYK^{+}12] Abdul R. Othman, The Sin Yan, H. J. Keselman, Rand R. Wilcox, and James Algina. Robust modifications of the Levene and O’Brien tests for spread. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/587.

Olubusoye:2016:MVA

[OYO16] Olusanya E. Olubusoye, Olaoluwa S. Yaya, and Oluwadare O. Ojo. Misspecification of variants of autoregressive GARCH models and effect on in-sample forecasting. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/855.

Ozgur:2016:ZDH

[Ozg16] Ceyhun Ozgur. Z and t distributions in hypothesis testing: Unequal division of Type I risk. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/809.

Padilla:2004:TER

[PA04] Miguel A. Padilla and James Algina. Type I error rates for a one factor within-subjects design with missing values. Journal of Modern Applied Statistical Methods, 3(2):406–416, November 2004. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/153.

Padilla:2007:TER

[PA07] Miguel A. Padilla and James Algina. Type I error rates of the Kenward–Roger adjusted degree of freedom F -test for a splitplot design with missing values. Journal of Modern Applied Statistical Methods, 6(1):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/283.

Pundir:2014:EAU

[PA14] Sudesh Pundir and R. Amala. Evaluation of area under the constant shape bi-Weibull ROC curve. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/706.

Pan:2018:FRM

[Pan18] Tianshu Pan. Fitting the Rasch model under the logistic regression framework to reduce estimation bias. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/969.

Papageorgiou:2016:OES

[Pap16] Ioulia Papageorgiou. Optimal estimation and sampling allocation in survey sampling under a general correlated superpopulation model. Journal of Modern Applied Statistical Methods, 15 (2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/853.

Parker:2010:DMI

[Par10] David Parker. Derivation of mass independent quantum treatment of phenomenon. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/489.

Pastushenko:2009:LES

[Pas09] Vassili F. Pastushenko. Least error sample distribution function. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/432.

Paul:2005:TGF

[Pau05] Sudhir R. Paul. Testing goodness of fit of the geometric distribution: an application to human fecundability data. Journal of Modern Applied Statistical Methods, 4(2):425–433, November 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/199.

Pesarin:2013:RAC

[PB13] Fortunato Pesarin and Stefano Bonnini. A response to Anderson’s (2013) Conceptual Distinction between the Critical p value and Type I Error Rate in Permutation Testing. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/635. See [And13a, And13b].

Poston:2014:MDS

[PC14] Dudley L. Poston and Eugenia Conde. Missing data and the statistical modeling of adolescent pregnancy. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/744.

Peng:2015:AAI

[PC15] Chao-Ying Joanne Peng and Li-Ting Chen. Algorithms for assessing intervention effects in single-case studies. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/767.

Peng:2018:HMD

[PC18] Chao-Ying Joanne Peng and Li-Ting Chen. Handling missing data in single-case studies. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/962.

Pan:2005:SSS

[PD05] Xuemei Pan and C. Mitchell Dayton. Sample size selection for pair-wise comparisons using information criteria. Journal of Modern Applied Statistical Methods, 4(2):601–608, November 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/219.

Pan:2011:FIM

[PD11] Xuemei Pan and C. Mitchell Dayton. Factors influencing the mixture index of model fit in contingency tables showing independence. Journal of Modern Applied Statistical Methods, 10(1): ??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/532.

Padilla:2013:BIE

[PD13] Miguel A. Padilla and Jasmin Divers. Bootstrap interval estimation of reliability via coefficient omega. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/644.

Parham:2013:AMD

[PDC13] G. Parham, A. Daneshkhah, and O. Chatrabgoun. Approximation multivariate distribution of main indices of Tehran Stock Exchange with pair-copula. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/684.

Phadnis:2008:MIP

[PE08] Milind A. Phadnis and Matthew E. Elam. A methodology to improve PCI use in industry. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/392.

Perry:2003:CEU

[Per03] Kimberly T. Perry. A critical examination of the use of preliminary tests in two-sample tests of location. Journal of Modern Applied Statistical Methods, 2(2):314–328, November 2003. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/89.

Perrett:2006:MAU

[Per06] Jamis J. Perrett. A method for analyzing unreplicated experiments using information on the intraclass correlation coefficient. Journal of Modern Applied Statistical Methods, 5(2):432–442, November 2006. CODEN ???? ISSN 1538-9472. URL https:/

/jmasm.com/index.php/jmasm/article/view/260.

Perez:2007:LEE

[Pér07a] Adriana Pérez. Longitudinal evaluation of estimates in an establishment survey after ration imputation. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/335.

Perez:2007:UFI

[Pér07b] Adriana Pérez. Using the fractional imputation methodology to evaluate variance due to hot deck imputation in survey data. Journal of Modern Applied Statistical Methods, 6(1):248–257, May 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/298.

Peterson:2002:SMA

[Pet02] Kathleen Peterson. Six modifications of the aligned rank transform test for interaction. Journal of Modern Applied Statistical Methods, 1(1):100–109, May 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/13.

Paidipati:2025:DMS

[PH25] Kiran Kumar Paidipati and Komaragiri Hyndhavi. Diet management study on Indian population through optimization models — the way towards reaching Blue Zone’s lifestyle. Journal of Modern Applied Statistical Methods, 24(1):54–74, ???? 2025. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/1334.

Ping:2008:AIC

[PI08] Hoo Ling Ping and M. Ataharul Islam. Analyzing incomplete categorical data: Revisiting maximum likelihood estimation (Mle) procedure. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/383.

Pathiyil:2007:SPE

[PJ07] Mathachan Pathiyil and E. S. Jeevanand. Semi parametric estimation of some reliability measures of geometric distribution. Journal of Modern Applied Statistical Methods, 6(2):??, ????

2007. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/322.

Para:2014:DGB

[PJ14] B. A. Para and T. R. Jan. Discrete generalized Burr-type XII distribution. Journal of Modern Applied Statistical Methods, 13 (2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/730.

Paraguas:2006:MSM

[PK06] Maria Divina S. Paraguas and Anton Abdulbasah Kamil. Model selection of meat demand system using the Rotterdam Model and the Almost Ideal Demand System (AIDS). Journal of Modern Applied Statistical Methods, 5(2):529–537, November 2006. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/269.

Prakash:2025:SAA

[PKRK25] Gyan Prakash, Manish Kumar, Shiv Kumar Rana, and Sukhi Gowda KE. A statistical approach for assessment of growth rate and instability of wheat in selected states of India. Journal of Modern Applied Statistical Methods, 24(1):75– 88, ???? 2025. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1335.

Poston:2003:UZI

[PM03] Dudley L. Poston, Jr. and Sherry L. McKibben. Using zeroinflated count regression models to estimate the fertility of U. S. women. Journal of Modern Applied Statistical Methods, 2(2): 371–379, November 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/94.

Poston:2008:MRM

[PM08] Dudley L. Poston, Jr. and Hosik Min. The multinomial regression modeling of the cause-of-death mortality of the oldest old in the U.S. Journal of Modern Applied Statistical Methods, 7(2): ??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/393.

Pearson:2010:RSS

[PM10] Robert H. Pearson and Daniel J. Mundform. Recommended sample size for conducting exploratory factor analysis on dichotomous data. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/494.

Pan:2014:FMC

[PM14] Yi Pan and Matthew T. McBee. A flexible method for conducting power analysis for two- and three-level hierarchical linear models in R. Journal of Modern Applied Statistical Methods, 13 (1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/716.

Padilla:2008:TER

[PMZ08] Miguel A. Padilla, YoungKyoung Min, and Guili Zhang. Type I error rates of the Kenward–Roger F -test for a split-plot design with missing values and non-normal data. Journal of Modern Applied Statistical Methods, 7(2):385–397, November 2008. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/373.

Peng:2003:UML

[PN03] Chao-Ying Joanne Peng and Rebecca Naegle Nichols. Using multinomial logistic models to predict adolescent behavioral risk. Journal of Modern Applied Statistical Methods, 2(1):177–188, May 2003. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/76.

Pazira:2010:EGE

[PN10] Hassan Pazira and Parviz Nasiri. Estimations on the generalized exponential distribution using grouped data. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/481.

Pastore:2008:CDM

[PNG08] Massimiliano Pastore, Massimo Nucci, and Giovanni Galfano. Comparing different methods for multiple testing in reaction time data. Journal of Modern Applied Statistical Methods, 7 (1):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/349.

Pratama:2024:FRU

[PNZM24] Muhammad Deo Pratama, Khoirin Nisa, La Zakaria, and Mona Arif Muda. Face recognition using principal component analysis with Euclidean distance and artificial neural network classification methods. Journal of Modern Applied Statistical Methods, 23(2):??, January 2024. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1259.

Posch:2002:AET

[Pos02a] Margaret Posch. Asymptotic and exact tests in 2 × C ordered categorical contingency tables with StatXact 2.0–4.0. Journal of Modern Applied Statistical Methods, 1(1):167–175, May 2002. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/24.

Poston:2002:SMF

[Pos02b] Dudley L. Poston, Jr. The statistical modeling of the fertility of Chinese women. Journal of Modern Applied Statistical Methods, 1(2):387–396, November 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/44.

Patel:2017:DPS

[PP17] Ronak M. Patel and Achyut C. Patel. The double prior selection for the parameter of exponential life time model under Type II censoring. Journal of Modern Applied Statistical Methods, 16(1): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/904.

Paneru:2018:EZI

[PPC18] Khyam Paneru, R. Noah Padgett, and Hanfeng Chen. Estimation of zero-inflated population mean: a bootstrapping approach. Journal of Modern Applied Statistical Methods, 17(1): ??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/966.

Pak:2014:IRD

[PPS14] Abbas Pak, Gholam Ali Parham, and Mansour Saraj. Inference for the Rayleigh distribution based on progressive Type-II fuzzy censored data. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/705.

Prakash:2010:BSS

[Pra10a] Gyan Prakash. On Bayesian shrinkage setup for item failure data under a family of life testing distribution. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/511.

Prakash:2010:SEI

[Pra10b] Gyan Prakash. Shrinkage estimation in the inverse Rayleigh distribution. Journal of Modern Applied Statistical Methods, 9 (1):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/478.

Prakash:2012:IED

[Pra12] Gyan Prakash. Inverted exponential distribution under a Bayesian viewpoint. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/598.

Prakash:2014:CPE

[Pra14] Gyan Prakash. Change point estimation for Pareto Type-II model. Journal of Modern Applied Statistical Methods, 13(1): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/708.

Prakash:2015:BAU

[Pra15] Gyan Prakash. Bayesian analysis under progressively censored Rayleigh data. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/776.

Prakash:2020:PDU

[Pra20] Gyan Prakash. Pareto distribution under hybrid censoring: Some estimation. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1083.

Press:2004:RGI

[Pre04] S. James Press. Respondent-generated intervals (RGI) for recall in sample surveys. Journal of Modern Applied Statistical Methods, 3(1):104–116, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/120.

Peng:2002:MSL

[PS02a] Chao-Ying Joanne Peng and Tak-Shing Harry So. Modeling strategies in logistic regression with SAS, SPSS, Systat, BMDP, Minitab, and STATA. Journal of Modern Applied Statistical Methods, 1(1):??, ???? 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/22.

Perry:2002:CDT

[PS02b] Kimberly T. Perry and Michael R. Stoline. A comparison of the D’Agostino S_{u} test to the triples test for testing of symmetry versus asymmetry as a preliminary test to testing the equality of means. Journal of Modern Applied Statistical Methods, 1(2): 316–325, November 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/37.

Potdar:2017:CIS

[PS17] Kiran Ganpati Potdar and D. T. Shirke. Confidence intervals for the scaled half-logistic distribution under progressive Type-II censoring. Journal of Modern Applied Statistical Methods, 16 (1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/900.

Pal:2018:EMT

[PS18] Surya K. Pal and Housila P. Singh. Estimation of mean with two-parameter ratio-product-ratio estimator in double sampling using ancillary information under non-response. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/999.

Patidar:2022:PRR

[PS22] Preeti Patidar and Housila P. Singh. Partial randomized response model for estimating a rare sensitive attribute in probability proportional to size measures using Poisson distribution. Journal of Modern Applied Statistical Methods, 21(1):??, February 2022. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1121.

Pham:2018:MLE

[PTC18] Minh H. Pham, Chris Tsokos, and Bong-Jin Choi. Maximum likelihood estimation for the generalized Pareto distribution and goodness-of-fit test with censored data. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1006.

Perkins:2004:CBF

[PW04] Jan Perkins and Daniel Wang. A comparison of Bayesian and frequentist statistics as applied in a simple repeated measures example. Journal of Modern Applied Statistical Methods, 3(1): 227–233, May 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/133.

Pushpa:2024:OML

[PW24] Pushpa and Zhang Wei. Optimal machine learning based forecasting model to analyze the impact of trade liberalization in China’s economy on export performance. Journal of Modern Applied Statistical Methods, 23(2):33–62, December 2024. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1381.

Padilla:2011:EIC

[PZ11] Miguel A. Padilla and Guili Zhang. Estimating internal consistency using Bayesian methods. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/542.

Qumsiyeh:2012:CRS

[QS12] Maher Qumsiyeh and Gerald Shaughnessy. Comparison of resampling methods to generalized linear models and transformations in factorial and fractional factorial designs. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/590.

Qu:2006:CPE

[QT06] Leming Qu and Yi-Cheng Tu. Change point estimation of bilevel functions. Journal of Modern Applied Statistical Methods, 5(2): 347–355, November 2006. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/252.

Qu:2005:BWE

[Qu05] Leming Qu. Bayesian wavelet estimation of long memory parameter. Journal of Modern Applied Statistical Methods, 4(1): 140–154, May 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/176.

Qumsiyeh:2013:UBE

[Qum13] Maher Qumsiyeh. Using the bootstrap for estimating the sample size in statistical experiments. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/640.

Qin:2016:LVM

[QWL^{+}16] Li Qin, Lisa Weissfeld, Michele Levine, Marsha Marcus, and Feng Dai. Latent variable model for weight gain prevention data with informative intermittent missingness. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/869.

Rahardja:2014:CIM

[Rah14] Dewi Rahardja. Comparison of individual and moving range chart combinations to individual charts in terms of ARL after designing for a common “All OK” ARL. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/736.

Rahardja:2017:RMS

[Rah17] Dewi Rahardja. A review of the multiple-sample tests for the continuous-data type. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/889.

Rashid:2018:NLD

[RAJ18] Adil Rashid, Zahooor Ahmad, and T. R. Jan. A new lifetime distribution for series system: Model, properties and application. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/960.

Rajaguru:2021:RSA

[Raj21] Gulasekaran Rajaguru. A response surface analysis of critical values for IPS panel unit root tests. Journal of Modern Applied Statistical Methods, 20(1):??, February 2021. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1110.

Rao:2011:DAS

[Rao11] G. Srinivasa Rao. Double acceptance sampling plans based on truncated life tests for Marshall–Olkin extended Lomax distribution. Journal of Modern Applied Statistical Methods, 10(1): ??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/528.

Rao:2014:ERM

[Rao14] Gadde Srinivasa Rao. Estimation of reliability in multicomponent stress-strength based on generalized Rayleigh distribution. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/710.

Rao:2017:BRM

[Rao17] C. R. Rao. Book review: Multivariate statistical methods, a primer. Journal of Modern Applied Statistical Methods, 16(1): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/922.

Ravichandran:2017:NDS

[Rav17] Joghee Ravichandran. A note on determination of sample size from the perspective of six sigma quality. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/897.

Razzaghi:2002:EBS

[Raz02] Mehdi Razzaghi. On the estimation of binomial success probability with zero occurrence in sample. Journal of Modern Applied Statistical Methods, 1(2):326–332, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/38.

Razzaghi:2013:PLF

[Raz13] Mehdi Razzaghi. The probit link function in generalized linear models for data mining applications. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/650.

Riviello:2009:DLO

[RB09] Clare Riviello and S. Natasha Beretvas. Detecting lag-one autocorrelation in interrupted time series experiments with small datasets. Journal of Modern Applied Statistical Methods, 8(2): ??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/437.

Radhakrishnan:2011:CCC

[RB11] R. Radhakrishnan and P. Balamurugan. Construction of control charts based on six sigma initiatives for the number of defects and average number of defects per unit. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/572.

Rosopa:2018:CUT

[RBAR18] Patrick J. Rosopa, Alice M. Brawley, Theresa P. Atkinson, and Stephen A. Robertson. On the conditional and unconditional Type I error rates and power of tests in linear models with heteroscedastic errors. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/995.

Rekabdar:2017:ADI

[RC17] Ghasem Rekabdar and Rahim Chinipardaz. Approximating the distribution of indefinite quadratic forms in normal variables by maximum entropy density estimation. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/942.

Rao:2016:BIM

[RD16] K. Aruna Rao and Juliet Gratia D’Cunha. Bayesian inference for median of the lognormal distribution. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/865.

Reed:2004:JCT

[Ree04] James F. Reed III. JMASM 11: Comparing two small binomial proportions. Journal of Modern Applied Statistical Methods, 3 (1):250–258, May 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/136.

Reed:2006:ABC

[Ree06] James F. Reed III. AB/BA crossover trials — binary outcome. Journal of Modern Applied Statistical Methods, 5(2):452–457, November 2006. CODEN ???? ISSN 1538-9472. URL https:/

/jmasm.com/index.php/jmasm/article/view/262.

Reed:2007:BBC

[Ree07] James F. Reed III. Better binomial confidence intervals. Journal of Modern Applied Statistical Methods, 6(1):153–161, May 2007. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/290.

Reed:2009:CFC

[Ree09a] James F. Reed III. Closed form confidence intervals for small sample matched proportions. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/441.

Reed:2009:ICI

[Ree09b] James F. Reed III. Improved confidence intervals for the difference between two proportions. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/412.

Reed:2011:EGC

[Ree11a] James F. Reed III. Extension of Grizzle’s classic crossover design. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/546.

Reed:2011:HOC

[Ree11b] James F. Reed III. Higher order C(t, p, s) crossover designs. Journal of Modern Applied Statistical Methods, 10(2):??, ????

2011. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/577.

Reed:2012:FPC

[Ree12] James F. Reed III. Four period crossover designs. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/607.

Reschenhofer:2008:FDM

[Res08] Erhard Reschenhofer. Frequency domain modeling with piecewise constant spectra. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/380.

Roberts:2003:AEC

[RH03] J. Kyle Roberts and Robin K. Henson. Not all effects are created equal: a rejoinder to sawilowsky. Journal of Modern Applied Statistical Methods, 2(1):226–230, May 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/61.

Rashid:2014:CGD

[RJ14] Adil Rashid and T. R. Jan. A compound of Geeta distribution with generalized beta distribution. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/704.

Rasekhi:2014:OLD

[RJR14] Mahdi Rasekhi, B. Jamshidi, and F. Rivaz. Optimal location design for prediction of spatial correlated environmental functional data. Journal of Modern Applied Statistical Methods, 13 (2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/743.

Rao:2013:DBG

[RK13] B. Srinivasa Rao and R. R. L. Kantam. Discriminating between generalized exponential distribution and some life test models based on population quantiles. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/679.

Ryan:2002:MCI

[RL02] George W. Ryan and Steven D. Leadbetter. On the misuse of confidence intervals for two means in testing for the significance of the difference between the means. Journal of Modern Applied Statistical Methods, 1(2):473–478, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/53.

Robinson:2010:QRC

[RL10] Daniel H. Robinson and Joel R. Levin. The not-so-quiet revolution: Cautionary comments on the rejection of hypothesis testing in favor of a “Causal” modeling alternative. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/491.

Rao:2016:TPN

[RL16] Calyampudi Radhakrishna Rao and Miodrag M. Lovric. Testing point null hypothesis of a normal mean and the truth: 21st Century perspective. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/

view/836. See remarks [ZK16].

Richter:2007:MCM

[RM07] Scott J. Richter and Melinda H. McCann. Multiple comparison of medians using permutation tests. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/314.

Rose:2016:JMC

[RM16] Sarah A. Rose and Barry Markman. JMASM 40: Monte Carlo simulations for structural equation modelling (Revolution R). Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/878.

Ramos:2014:OPE

[RMA14] Pedro L. Ramos, Fernando A. Moala, and Jorge A. Achcar. Objective priors for estimation of extended exponential geometric distribution. Journal of Modern Applied Statistical Methods, 13 (2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/729.

Rose:2017:LSA

[RMS17] Sarah A. Rose, Barry Markman, and Shlomo Sawilowsky. Limitations in the systematic analysis of structural equation model fit indices. Journal of Modern Applied Statistical Methods, 16 (1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/886.

Ruxton:2018:SSE

[RN18] Graeme Ruxton and Markus Neuhäuser. Striving for simple but effective advice for comparing the central tendency of two populations. Journal of Modern Applied Statistical Methods, 17(2): ??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/992. See re-

sponses [NKC18].

Ratnaparkhi:2012:LBL

[RNN12] Makarand V. Ratnaparkhi and Uttara V. Naik-Nimbalkar. The length-biased lognormal distribution and its application in the analysis of data from oil field exploration studies. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/604.

Ratnaparkhi:2013:LBV

[RNN13a] Makarand V. Ratnaparkhi and Uttara V. Naik-Nimbalkar. The length-biased versus random sampling for the binomial and Poisson events. Journal of Modern Applied Statistical Methods, 12 (1):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/641.

Ratnaparkhi:2013:NCM

[RNN13b] Makarand V. Ratnaparkhi and Uttara V. Naik-Nimbalkar. A note on α-curvature of the manifolds of the length-biased lognormal and gamma distributions in view of related applications in data analysis. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/658.

Rodgers:2010:SMM

[Rod10] Joseph Lee Rodgers. Statistical and mathematical modeling versus NHST? There’s no competition! Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/492.

Richter:2003:JUS

[RP03a] Scott J. Richter and Mark E. Payton. JMASM 8: Using SAS to perform two-way analysis of variance under variance heterogeneity. Journal of Modern Applied Statistical Methods, 2(2): 520–524, November 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/109.

Richter:2003:PTW

[RP03b] Scott J. Richter and Mark E. Payton. Performing two-way analysis of variance under variance heterogeneity. Journal of Modern Applied Statistical Methods, 2(1):??, ???? 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/73.

Richards:2003:WAQ

[RR03] Tom Richards and Lyn Richards. The way ahead in qualitative computing. Journal of Modern Applied Statistical Methods, 2 (1):16–26, May 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/59.

Ramsey:2007:OTO

[RR07] Philip H. Ramsey and Patricia P. Ramsey. Optimal trimming and outlier elimination. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/310.

Ren:2024:NTB

[RR24] Luh Yu Ren and Peter Ren. Normality of the T -tests for buy and sell days from moving average trading rules on the NASDAQ. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1181.

Rajendran:2008:LLM

[RRB08] Krishnan Rajendran, Thandavarayan Ramamurthy, and Sujit Kumar Bhattacharya. Log-linear model to assess socioeconomic and environmental factors with childhood diarrhea using hospital based surveillance. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/365.

Rao:2012:EVC

[RRB12] B. Srinivasa Rao, J. Pratapa Reddy, and G. Sarath Babu. Extreme value charts and analysis of means (ANOM) based on the log logistic distribution. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/628.

Ramsey:2011:RPI

[RRHA11] Philip H. Ramsey, Patricia P. Ramsey, Priscila Hachimine, and Nancy Andiloro. Robustness, power and interpretability of pairwise tests of discriminant functions in MANOVA. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/552.

Rajendran:2007:MLR

[RRS07] Krishnan Rajendran, Thandavarayan Ramamurthy, and Dipika Sur. Multinomial logistic regression model for the inferential risk age groups for infection caused by Vibrio cholerae in Kolkata, India. Journal of Modern Applied Statistical Methods, 6(1):324– 330, May 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/305.

Rao:2023:AMB

[RRS23] B. Srinivasa Rao, B. Vara Prasad Rao, and V. R. Bala Suseela. Analysis of means based on exponentiated inverted Weibull distribution. Journal of Modern Applied Statistical Methods, 22(2): ??, July 2023. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1136.

Ranathunga:2017:MMM

[RS17] Kapuruge N. O. Ranathunga and Roshini Sooriyarachchi. Multivariate multilevel modeling of age related diseases. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/909.

RamaShanker:2020:GGP

[RS20] RamaShanker and Kamlesh Kumar Shukla. A generalization of generalized Poisson–Lindley distribution and its applications. Journal of Modern Applied Statistical Methods, 19(2):??, July 2020. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1105.

Rai:2020:CBE

[RSQ20] Piyush Kant Rai, Alka Singh, and Muhammad Qasim. Calibration-based estimators using different distance measures under two auxiliary variables: a comparative study. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1082.

Rai:2022:CEU

[RTA22] Piyush Kant Rai, Shiwani Tiwari, and Alka. Calibration estimators under two auxiliary variables using linear and nonlinear constraints in sample surveys. Journal of Modern Applied Statistical Methods, 21(1):??, February 2022. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1219.

Rodrigo:2016:RNN

[RTS16] Hansapani S. Rodrigo, Chris P. Tsokos, and Taysseer Sharaf. Regularized neural network to identify potential breast cancer: a Bayesian approach. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/867.

Ruscio:2008:CCI

[Rus08] John Ruscio. Constructing confidence intervals for Spearman’s rank correlation with ordinal data: a simulation study comparing analytic and bootstrap methods. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/376.

Ratnaparkhi:2009:BMS

[RWS09] Makarand V. Ratnaparkhi, Vasant B. Waikar, and Fredrick J. Schuurmann. The bootstrap method for the selection of a shrinkage factor in two-stage estimation of the reliability function of an exponential distribution. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/415.

Rahardja:2016:CRT

[RYZ16] Dewi Rahardja, Ying Yang, and Zhiwei Zhang. A comprehensive review of the two-sample independent or paired binary data, with or without stratum effects. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/849.

Sen:2002:SSC

[SA02] Pali Sen and Mary Anderson. Simulation study of chemical inhibition modeling. Journal of Modern Applied Statistical Methods, 1(2):397–404, November 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/45.

Shamilov:2007:FHB

[SA07] Aladdin Shamilov and Senay Asma. A Fano–Huffman based statistical coding method. Journal of Modern Applied Statistical Methods, 6(1):265–278, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/300.

Samkar:2010:RRB

[SA10] Hatice Samkar and Ozlem Alpu. Ridge regression based on some robust estimators. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/506.

Smadi:2011:BTM

[SA11] Mahmoud M. Smadi and M. T. Alodat. Bayesian threshold moving average models. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/540.

Sultan:2012:PEP

[SA12] Raja Sultan and S. P. Ahmad. Posterior estimates of Poisson distribution using R software. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/632.

Sultan:2013:CPL

[SA13] Raja Sultan and S. P. Ahmad. Comparison of parameters of lognormal distribution based on the classical and posterior estimates. Journal of Modern Applied Statistical Methods, 12(2): ??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/677.

Shittu:2014:EWD

[SA14] Olanrewaju I. Shittu and K. A. Adepoju. On the exponentiated Weibull distribution for modeling wind speed in South Western Nigeria. Journal of Modern Applied Statistical Methods, 13(1): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/714.

Smadi:2012:ECO

[SAA12] Abdullah A. Smadi and Nour H. Abu-Afouna. An extension of Cochran–Orcutt procedure for generalized linear regression models with periodically correlated errors. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/621.

Saalman:2019:ODP

[Saa19a] Dustin R. Saalman. Objectives driven participatory evaluation model. Journal of Modern Applied Statistical Methods, 18(2): ??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1064.

Sadik:2019:SAE

[SAA19b] Kusman Sadik, Rahma Anisa, and Euis Aqmaliyah. Small area estimation on zero-inflated data using frequentist and Bayesian approach. Journal of Modern Applied Statistical Methods, 18(1): ??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1022.

Safi:2008:VET

[Saf08] Samir Safi. Variance estimation in time series regression models. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/385.

Safi:2011:EEA

[Saf11] Samir Safi. Explicit equations for ACF in autoregressive processes in the presence of heteroscedasticity disturbances. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/570.

Saha:2005:KWM

[Sah05] Amitava Saha. Kim and Warde’s mixed randomized response technique for complex surveys. Journal of Modern Applied Statistical Methods, 4(2):538–544, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/210.

Saha:2009:QTT

[Sah09] Krishna K. Saha. Quel test for two linear restrictions in the nonlinear models. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/406.

Sawilowsky:2002:CQM

[Saw02a] Ellen F. Sawilowsky. Combining quantum mechanical calculations and a χ^{2} fit in a potential energy function for the CO_{2} + O^{+} reaction. Journal of Modern Applied Statistical Methods, 1(2): 405–410, November 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/46.

Sawilowsky:2002:FSE

[Saw02b] Shlomo S. Sawilowsky. Fermat, Schubert, Einstein, and Behrens–Fisher: the probable difference between two means when σ_{1}^{2} ≠ σ_{2}^{2}. Journal of Modern Applied Statistical Methods, 1(2):461–472, November 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/52.

Sawilowsky:2002:MRE

[Saw02c] Shlomo S. Sawilowsky. A measure of relative efficiency for location of a single sample. Journal of Modern Applied Statistical Methods, 1(1):52–60, May 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/8.

Sawilowsky:2003:DAC

[Saw03a] Shlomo S. Sawilowsky. Deconstructing arguments from the case against hypothesis testing. Journal of Modern Applied Statistical Methods, 2(2):467–474, November 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/103.

Sawilowsky:2003:DFS

[Saw03b] Shlomo S. Sawilowsky. A different future for social and behavioral science research. Journal of Modern Applied Statistical Methods, 2(1):128–132, May 2003. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/71.

Sawilowsky:2003:TBS

[Saw03c] Shlomo S. Sawilowsky. Trivials: the birth, sale, and final production of meta-analysis. Journal of Modern Applied Statistical Methods, 2(1):242–246, May 2003. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/64.

Sawilowsky:2003:YTY

[Saw03d] Shlomo S. Sawilowsky. You think you’ve got trivials? Journal of Modern Applied Statistical Methods, 2(1):218–225, May 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/60.

Sawilowsky:2004:CRC

[Saw04a] Shlomo S. Sawilowsky. A conversation with R. Clifford Blair on the occasion of his retirement. Journal of Modern Applied Statistical Methods, 3(2):518–566, November 2004. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/148.

Sawilowsky:2004:TRA

[Saw04b] Shlomo S. Sawilowsky. Teaching random assignment: Do you believe it works? Journal of Modern Applied Statistical Methods, 3(1):221–226, May 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/132.

Sawilowsky:2005:MLC

[Saw05] Shlomo S. Sawilowsky. Misconceptions leading to choosing the t test over the Wilcoxon Mann–Whitney test for shift in location parameter. Journal of Modern Applied Statistical Methods, 4(2): ??, ???? 2005. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/218.

Sawilowsky:2007:MVS

[Saw07] Shlomo S. Sawilowsky. Mathmatics in volume I of scripta universitatis. Journal of Modern Applied Statistical Methods, 6(1): ??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/309.

Sawilowsky:2009:NES

[Saw09] Shlomo S. Sawilowsky. New effect size rules of thumb. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/452.

Sawan:2013:ADS

[Saw13] Zakaria M. Sawan. An approach for dealing with statuses of nonstatistically significant interactions between treatments. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/655.

Sawilowsky:2016:RLT

[Saw16] Shlomo S. Sawilowsky. Rao–Lovric and the triwizard point null hypothesis tournament. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/837.

Shafiq:2017:JAC

[SAZ17] Mohamad Shafiq, Wan Muhamad Amir, and Nur Syabiha Zafakali. JMASM 46: Algorithm for comparison of robust regression methods in multiple linear regression by weighting least square regression (SAS). Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/953.

Smith:2009:ESM

[SB09] Lindsey J. Wolff Smith and S. Natasha Beretvas. Estimation of the standardized mean difference for repeated measures designs. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/453.

Stamey:2007:BSS

[SBY07] James D. Stamey, Thomas L. Bratcher, and Dean M. Young. Bayesian subset selection of binomial parameters using possibly misclassified data. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/328.

Sawilowsky:2006:JLM

[SC06] Shlomo S. Sawilowsky and John L. Cuzzocrea. Joseph Liouville’s ‘Mathematical Works of Évariste Galois’. Journal of Modern Applied Statistical Methods, 5(2):??, ???? 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/275.

Singh:2017:CCM

[SD17] J. R. Singh and Ab Latif Dar. Control charts for mean for non-normally correlated data. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/906.

Sigirli:2012:EGS

[SE12] Deniz Sigirli and Ilker Ercan. Examining growth with statistical shape analysis and comparison of growth models. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/627.

Sen:2002:SGM

[Sen02] Pranab K. Sen. Shifting goals and mounting challenges for statistical methodology. Journal of Modern Applied Statistical Methods, 1(1):2–12, May 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/2.

Seo:2019:JML

[Seo19] JangDong Seo. Joint models of longitudinal outcomes and informative time. Journal of Modern Applied Statistical Methods, 18 (1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/1039.

Serlin:2002:CC

[Ser02] Ronald C. Serlin. Constructive criticism. Journal of Modern Applied Statistical Methods, 1(2):??, ???? 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/28.

Serlin:2010:FWR

[Ser10] Ronald C. Serlin. Fisher was right. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/460.

Shirvani:2019:AAA

[SF19] Alireza Shirvani and Malek Fathizadeh. Assessing the accuracy of approximate confidence intervals proposed for the mean of Poisson distribution. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1020.

Sindhu:2014:BEP

[SFA14] Tahassum N. Sindhu, Navid Feroze, and Muhammad Aslam. Bayesian estimation of the parameters of two-component mixture of Rayleigh distribution under doubly censoring. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/731.

Sindhu:2016:DCD

[SFA16a] Tabassum Sindhu, Navid Feroze, and Muhammad Aslam. Doubly censored data from two-component mixture of inverse Weibull distributions: Theory and applications. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/854.

Sindhu:2016:SLC

[SFA16b] Tabassum Naz Sindhu, Navid Feroze, and Muhammad Aslam. Study of the left censored data from the Gumbel Type II distribution under a Bayesian approach. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/843.

Singh:2021:CTA

[SG21] Gurinder Pal Singh and Davinder Kumar Garg. Construction of three associate PBIB designs using K-maps. Journal of Modern Applied Statistical Methods, 20(1):??, February 2021. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1111.

Singh:2016:AUE

[SGM16] Rajesh Singh, S. B. Gupta, and Sachin Malik. Almost unbiased estimator using known value of population parameter(s) in sample surveys. Journal of Modern Applied Statistical Methods, 15 (1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/815.

Sheng:2007:JAI

[SH07] Yanyan Sheng and Todd C. Headrick. JMASM 27: an algorithm for implementing Gibbs sampling for 2PNO IRT models (Fortran). Journal of Modern Applied Statistical Methods, 6(1): ??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/308.

Sheng:2009:JGS

[SH09] Yanyan Sheng and Todd C. Headrick. JMASM 28: Gibbs sampling for 2PNO multi-unidimensional item response theory models (Fortran). Journal of Modern Applied Statistical Methods, 8 (2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/458.

Sadia:2014:CBC

[SH14] Farhana Sadia and Syed S. Hossain. Contrast of Bayesian and classical sample size determination. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/740.

Sharif:2015:RPU

[SH15] Nurul Sima Mohamad Sharif and Nor Aishah Hamzah. A robust panel unit root test in the presence of cross sectional dependence. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/779.

Shaffer:2006:CIS

[Sha06a] Juliet Popper Shaffer. Confidence intervals on subsets may be misleading. Journal of Modern Applied Statistical Methods, 5(1): 281, May 2006. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/244.

Shanmugam:2006:CBN

[SHA06b] Ramalingam Shanmugam, Anwar Hassan, and Peer Bilal Ahmad. Correlation between the number of epileptic and healthy children in family size that follows a size-biased modified power series distribution. Journal of Modern Applied Statistical Methods, 5(2):458–463, November 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/263.

Shanmugam:2008:CBS

[Sha08a] Ramalingam Shanmugam. Correlation between the sample mean and sample variance. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/375.

Sharp:2008:TIT

[Sha08b] Stephen Sharp. Tests for independence in two-way contingency tables with small samples. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/353.

Shanmugam:2011:NTM

[Sha11] Ramalingam Shanmugam. Is next twelve months period tumor recurrence free under restricted rate due to medication? A probabilistic warning. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/547.

Sinha:2011:LRM

[SIA11] Narayan Chanra Sinha, M. Ataharul Islam, and Kazi Saleh Ahamed. Logistic regression models for higher order transition probabilities of Markov chain for analyzing the occurrences of daily rainfall data. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/548.

Siev:2005:EIE

[Sie05] David Siev. An estimator of intervention effect on disease severity. Journal of Modern Applied Statistical Methods, 4(2):500–508, November 2005. CODEN ???? ISSN 1538-9472. URL https:/

/jmasm.com/index.php/jmasm/article/view/205.

Shabri:2007:LMS

[SJ07] Ani Shabri and Abdul Aziz Jemain. LQ-moments for statistical analysis of extreme events. Journal of Modern Applied Statistical Methods, 6(1):228–238, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/296.

Suganthi:2025:GTB

[SJ25] J. Suganthi and I. Paulraj Jayasimman. Graph theory-based multi-objective Archimedes search and rescue optimization algorithm paradigm: an explicit consideration to urban drainage networks layouts. Journal of Modern Applied Statistical Methods, 24(1):??, ???? 2025. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1395.

Smadi:2014:RSE

[SJAZ14] A. A. Smadi, J. J. Jaber, and A. G. Al-Zu’bi. Robustness of several estimators of the ACF of AR(1) process with non-Gaussian errors. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/696.

Shakil:2007:PMR

[SK07] M. Shakil and B. M. Golam Kibria. On the product of Maxwell and Rice random variables. Journal of Modern Applied Statistical Methods, 6(1):212–218, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/294.

Shah:2009:WBP

[SK09] Hiral A. Shah and Sema A. Kalaian. Which is the best parametric statistical method for analyzing Delphi data? Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/414.

Singh:2010:EME

[SK10] Housila P. Singh and Namrata Karpe. Effect of measurement errors on the separate and combined ratio and product estimators in stratified random sampling. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/497.

Suresh:2011:IOF

[SK11] S. Suresh and K. Senthamarai Kannan. Identifying outliers in fuzzy time series. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/580.

Subramani:2013:EVU

[SK13] J. Subramani and G. Kumarapandiyan. Estimation of variance using known coefficient of variation and median of an auxiliary variable. Journal of Modern Applied Statistical Methods, 12(1): ??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/642.

Sakate:2014:CEG

[SK14] D. M. Sakate and D. N. Kashid. Comparison of estimators in GLM with binary data. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/727.

Singh:2018:ECE

[SK18] G. N. Singh and Mohd Khalid. Efficient class of estimators for finite population mean using auxiliary information in twooccasion successive sampling. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/998.

Singh:2011:EPM

[SKB11] Housila P. Singh, Sunil Kumar, and Sandeep Bhougal. Estimation of population mean in successive sampling by sub-sampling non-respondents. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/522.

Sangaiah:2015:ESD

[SKB15] Ilangovan Sangaiah, A. Vincent Antony Kumar, and Appavu Balamurugan. An empirical study on different ranking methods for effective data classification. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/772.

Safiih:2009:SPS

[SKO09] L. Muhamad Safiih, A. A. Kamil, and M. T. Abu Osman. Semiparametric of sample selection model using fuzzy concepts. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/445.

Sujatha:2025:AIW

[SKSV25] S. Sujatha, A. Dinesh Kumar, R. Sivaraman, and M. Vasuki. Application of the inverse Weibull distribution to agricultural data using fuzzy parameter intervals. Journal of Modern Applied Statistical Methods, 24(1):??, ???? 2025. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1398.

Singh:2016:MBO

[SL16] D. Asir Antony Gnana Singh and E. Jebalamar Leavline. Modelbased outlier detection system with statistical preprocessing. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/825.

Sajobi:2017:RMV

[SL17] Tolulope T. Sajobi and Lisa M. Lix. Robust measures of variable importance for multivariate group designs. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/929.

Szczesniak:2016:SMM

[SLA16] Rhonda D. Szczesniak, Dan Li, and Raouf S. Amin. Semiparametric mixed models for nested repeated measures applied to ambulatory blood pressure monitoring data. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/799.

Santos:2017:MCP

[SLC17] Elisa Norberto Ferreira Santos, Gilberto Rodrigues Liska, and Marcelo Angelo Cirillo. Methodology for constructing perceptual maps incorporating measuring error in sensory acceptance tests. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/899.

Sen:2014:BDC

[SLD14] Sumen Sen, Rajan Lamichhane, and Norou Diawara. A bivariate distribution with conditional gamma and its multivariate form. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/726.

Sajobi:2011:DAR

[SLLL11] Tolulope T. Sajobi, Lisa M. Lix, Longhai Li, and William Laverty. Discriminant analysis for repeated measures data: Effects of mean and covariance misspecification on bias and error in discriminant function coefficients. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/565.

Sawilowsky:2002:UTU

[SM02] Shlomo S. Sawilowsky and Barry S. Markman. Using the t test with uncommon sample sizes. Journal of Modern Applied Statistical Methods, 1(1):143–146, May 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/20.

Shukur:2004:SPR

[SM04] Ghazi Shukur and Panagiotis Mantalos. Size and power of the RESET test as applied to systems of equations: a bootstrap approach. Journal of Modern Applied Statistical Methods, 3(2): 370–385, November 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/145.

Sawilowsky:2017:EWT

[SM17] Jack Sawilowsky and Barry Markman. Experiment-wise Type I error rates in nested (hierarchical) study designs. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/885.

Sen:2016:XDS

[SMC16] Subhradev Sen, Sudhansu S. Maiti, and N. Chandra. The xgamma distribution: Statistical properties and application. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/824.

Smith:2009:IVU

[Smi09] Julie M. Smith. Intermediate r values for use in the Fleishman power method. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/454.

Sabry:2020:PEB

[SMSN20] Mohamed Abd Elhamed Sabry, Hiba Zeyada Muhammed, Mostafa Shaaban, and Abd El Hady Nabih. Parameter estimation based on double ranked set samples with applications to Weibull distribution. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1098.

Subramani:2014:MBM

[SP14a] Jambulingam Subramani and G. Prabavathy. Median based modified ratio estimators with known quartiles of an auxiliary variable. Journal of Modern Applied Statistical Methods, 13(1): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/701.

Subramani:2014:TPM

[SP14b] Jambulingam Subramani and G. Prabavathy. Two parameter modified ratio estimators with two auxiliary variables for estimation of finite population mean with known skewness, kurtosis and correlation coefficient. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/699.

Sundaram:2016:DOT

[SP16] Sampath Sundaram and Deepa S. Parthasarathy. Determination of optimal tightened normal tightened plan using a genetic algorithm. Journal of Modern Applied Statistical Methods, 15 (1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/835.

Sofi:2019:NTP

[SPB19] Saima Manzoor Sofi, Safina Peerzada, and Mirza Abdul Khalique Baig. A new two-parametric ‘useful’ fuzzy information measure and its properties. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1054.

Spence:2004:RRC

[Spe04] Patric R. Spence. On the reporting of reliability in content analysis. Journal of Modern Applied Statistical Methods, 3(1):234– 238, May 2004. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/134.

Shukla:2012:CFT

[SPT12] Diwakar Shukla, Sharad Pathak, and Narendra Singh Thakur. Class(es) of factor-type estimator(s) in presence of measurement error. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/614.

Samuh:2015:EPE

[SQ15] Monjed H. Samuh and Areen Qtait. Estimation for the parameters of the exponentiated exponential distribution using a median ranked set sampling. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/764.

Stark:2007:SCA

[SR07] D. B. Stark and J. F. Reed III. Sensitivity curves for asymmetric trimming hinge estimators. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/319.

Supandi:2017:ECB

[SRA17] Epha Diana Supandi, Dedi Rosadi, and Abdurakhman. An empirical comparison between robust estimation and robust optimization to mean-variance portfolio. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/914.

Shang:2021:GFT

[SRW21] Justin Shang, Timothy J. Robinson, and Shaun S. Wulff. Goodness-of-fit tests in logistic regression with continuous covariates. Journal of Modern Applied Statistical Methods, 20(2): ??, July 2021. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1116.

Samawi:2004:SER

[SS04] Hani M. Samawi and Laith J. Saeid. Stratified extreme ranked set sample with application to ratio estimators. Journal of Modern Applied Statistical Methods, 3(1):117–133, May 2004. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/121.

Soloman:2009:IRB

[SS09] Shira R. Soloman and Shlomo S. Sawilowsky. Impact of rankbased normalizing transformations on the accuracy of test scores. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/435.

Sankle:2012:SSP

[SS12] R. Sankle and J. R. Singh. Single sampling plans for variables indexed by AQL and AOQL with measurement error. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/620.

Sharma:2013:GCE

[SS13] Prayas Sharma and Rajesh Singh. A generalized class of estimators for finite population variance in presence of measurement errors. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/672.

Singh:2014:SRR

[SS14] Sarjinder Singh and Stephen A. Sedory. Stochastic randomized response model for a quantitative sensitive random variable. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/711.

Singh:2015:MEP

[SS15] Rajesh Singh Singh and Prayas Sharma. Method of estimation in the presence of non-response and measurement errors simultaneously. Journal of Modern Applied Statistical Methods, 14(1): ??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/757.

Sharma:2016:EPM

[SS16] Anup Kumar Sharma and Garib Nath Singh. Estimation of population mean on recent occasion under non-response in hoccasion successive sampling. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/845.

Szulczyk:2018:UCC

[SS18] Kenneth R. Szulczyk and Shibley Sadique. Using cyclical components to improve the forecasts of the stock market and macroeconomic variables. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/977.

Shanker:2023:GGP

[SS23] Rama Shanker and Kamlesh Kumar Shukla. A generalization of generalized Poisson–Lindley distribution and its applications. Journal of Modern Applied Statistical Methods, 22(2):??, October 2023. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/1198.

Sharma:2024:EEE

[SS24] Anup Kumar Sharma and Alok Kumar Singh. Enhanced estimators and effective estimation procedures for population variance under missing at random data in successive sampling. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1179.

Sideridis:2018:ICR

[SSAH18] Georgios Sideridis, Abdullah Saddaawi, and Khaleel Al-Harbi. Internal consistency reliability in measurement: Aggregate and multilevel approaches. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/984.

Singh:2013:VSP

[SSK13] J. R. Singh, R. Sankle, and M. Ahmad Khanday. Variables sampling plan for correlated data. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/669.

Salamwade:2018:DLB

[SSM18] R. L. Salamwade, D. M. Sakate, and S. K. Mathur. φ-divergence loss-based artificial neural network. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/994.

Soh:2022:EAE

[SSMK22] Patrice Takam Soh, Florant Ulrich Simo, Ibrahim Moukouop, and Eugene Kouassi. An EM-algorithm for estimating the lifetime distribution with long terms survival data. Journal of Modern Applied Statistical Methods, 21(1):??, February 2022. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1119.

Shangodoyin:2010:ATT

[SSMS10] D. K. Shangodoyin, K. Setlhare, K. K. Moseki, and K. Sediakgotla. On the appropriate transformation technique and model selection in forecasting economic time series: an application to Botswana GDP data. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/486.

Singh:2016:IGE

[SSS16] G. N. Singh, Alok Kumar Singh, and Anup Kumar Sharma. An improved generalized estimation procedure of current population mean in two-occasion successive sampling. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/847.

Singh:2021:SPM

[SSS21] Priyanka Singh, Ajeet Kumar Singh, and V. K. Singh. A super population model approach to imputation for estimating population mean. Journal of Modern Applied Statistical Methods, 20 (1):??, February 2021. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1112.

Samawi:2002:DMR

[ST02] Hani M. Samawi and Eman M. Tawalbeh. Double median ranked set sample: Comparing to other double ranked samples for mean and ratio estimators. Journal of Modern Applied Statistical Methods, 1(2):428–442, November 2002. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/49.

Shih:2007:WMA

[ST07] Shou Hsing Shih and Chris P. Tsokos. A weighted moving average process for forcasting. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/334.

Shih:2008:WMA

[ST08] Shou Hsing Shih and Chris P. Tsokos. A weighted moving average process for forecasting. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/354.

Sharaf:2014:PST

[ST14] Taysseer Sharaf and Chris P. Tsokos. Predicting survival time of localized melanoma patients using discrete survival time method. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/695.

Stamatopoulos:2002:GCU

[Sta02] Constantine Stamatopoulos. Generation of combinations using Excel. Journal of Modern Applied Statistical Methods, 1(1):191– 194, May 2002. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/26.

Stephens:2016:LSE

[Ste16] Joshua Stephens. Longitudinal stability of effect sizes in education research. Journal of Modern Applied Statistical Methods, 15 (2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/840.

Singh:2011:MRP

[STJ11] Housila P. Singh, Rajesh Tailor, and Narendra Kumar Jatwa. Modified ratio and product estimators for population mean in systematic sampling. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/554.

Stoker:2004:THC

[Sto04] Howard Stoker. “Teaching” in honor of Cliff Blair. Journal of Modern Applied Statistical Methods, 3(2):578–580, November 2004. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/151.

Su:2005:DAF

[Su05] Steve Su. A discretized approach to flexibly fit generalized lambda distributions to data. Journal of Modern Applied Statistical Methods, 4(2):408–424, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/198.

Su:2011:MLL

[Su11] Steve Su. Maximum log likelihood estimation using EM algorithm and partition maximum log likelihood estimation for mixtures of generalized lambda distributions. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/567.

Su:2016:FFP

[Su16a] Steve Su. Fitting flexible parametric regression models with GL-Dreg in R. Journal of Modern Applied Statistical Methods, 15 (2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/877.

Suresh:2016:DBS

[SU16b] K. K. Suresh and S. Umamaheswari. Designing of Bayesian skip lot sampling plan under destructive testing. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/874.

Subramani:2013:GMR

[Sub13] Jambulingam Subramani. Generalized modified ratio estimator for estimation of finite population mean. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/666.

Scagliarini:2008:CMP

[SV08] Michele Scagliarini and Raffaele Vermiglio. Computing multivariate process capability indices (Excel). Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/369.

Schaffer:2012:GEV

[SV12] Jay R. Schaffer and Shawn VandenHul. A graphical examination of variable deletion within the MEWMA statistic. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/611.

Singh:2016:EPM

[SVS16] Rajesh Singh, Hemant K. Verma, and Prayas Sharma. Estimation of population mean using exponential type imputation technique for missing observations. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/804.

Safi:2006:EOP

[SW06] Samir Safi and Alexander White. The efficiency of OLS in the presence of auto-correlated disturbances in regression models. Journal of Modern Applied Statistical Methods, 5(1):107–117, May 2006. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/230.

Shehata:2008:RMC

[SW08] Yasser A. Shehata and Paul White. A randomization method to control the Type I error rates in best subset regression. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/374.

Smith:2019:ELF

[SWM19] Thomas J. Smith, David A. Walker, and Cornelius M. McKenna. An exploration of link functions used in ordinal regression. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/1030.

Sawilowsky:2002:TT

[SY02] Shlomo S. Sawilowsky and Jina S. Yoon. The trouble with trivials (p > 0.05). Journal of Modern Applied Statistical Methods, 1(1):143–144, May 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/19.

Singh:2019:NEA

[SY19] Housila P. Singh and Anita Yadav. A new exponential approach for reducing the mean squared errors of the estimators of population mean using conventional and non-conventional location parameters. Journal of Modern Applied Statistical Methods, 18 (1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/1035.

Srivastava:2018:REI

[SZH^{+}18] Deo Kumar Srivastava, Liang Zhu, Melissa M. Hudson, Jianmin Pan, and Shesh N. Rai. Robust estimation and inference on current status data with applications to phase IV cancer trial.

Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/971.

Takahashi:2017:JIM

[Tak17a] Masayoshi Takahashi. JMASM 44: Implementing multiple ratio imputation by the EMB algorithm (R). Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/918.

Takahashi:2017:MRI

[Tak17b] Masayoshi Takahashi. Multiple ratio imputation by the EMB algorithm: Theory and simulation. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/911.

Taylor:2011:JMP

[Tay11] Alan Taylor. JMASM 31: MANOVA procedure for power calculations (SPSS). Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/583.

Tailor:2012:RTE

[TC12] Rajesh Tailor and Sunil Chouhan. Ratio type estimator of ratio of two population means in stratified random sampling. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/608.

Thatcher:2003:TSC

[TD03] W. Gregory Thatcher and J. Wanzer Drane. On treating a survey of convenience sample as a simple random sample. Journal of Modern Applied Statistical Methods, 2(2):478–480, November 2003. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/105.

Thompson:2005:PTN

[TGhC^{+}05] Marilyn S. Thompson, Samuel B. Green, Yi hsin Chen, Shawn Stockford, and Wen juo Lo. Power of the t test for normal and mixed normal distributions. Journal of Modern Applied Statistical Methods, 4(2):591–597, November 2005. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/217.

Thompson:2009:ATM

[Tho09] Daniel R. Thompson. Assessing trends: Monte Carlo trials with four different regression methods. Journal of Modern Applied Statistical Methods, 8(2):??, ???? 2009. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/430.

Thomakos:2010:MUO

[Tho10] Dimitrios D. Thomakos. Median-unbiased optimal smoothing and trend extraction. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/473.

Thompson:2016:GEF

[Tho16] Christopher G. Thompson. Graphing effects as fuzzy numbers in meta-analysis. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/828.

Tighiouart:2003:MCT

[Tig03] Mourad Tighiouart. Modeling correlated time-varying covariate effects in a Cox-type regression model. Journal of Modern Applied Statistical Methods, 2(1):161–167, May 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/74.

Thangavel:2006:CSD

[TK06] Kuttiannan Thangavel and Durairaj Ashok Kumar. A combined standard deviation based data clustering algorithm. Journal of Modern Applied Statistical Methods, 5(1):258–265, May 2006. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/241.

Tailor:2014:SRT

[TL14] Rajesh Tailor and Hilal A. Lone. Separate ratio-type estimators of population mean in stratified random sampling. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/700.

Tikhomirova:2015:SMM

[TL15] Tatiana Tikhomirova and Yulia Lebedeva. Statistical modeling of migration attractiveness of the EU member states. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/784.

Troendle:2015:PFE

[TMB15] James F. Troendle, Keshia-Lee Martin, and Vance W. Berger. Per family error rates: a response. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/752.

Turner:2017:PEC

[TNH17] Heather J. Turner, Prathiba Natesan, and Robin K. Henson. Performance evaluation of confidence intervals for ordinal coefficient alpha. Journal of Modern Applied Statistical Methods, 16 (2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/932.

Tofallis:2008:LSP

[Tof08] Chris Tofallis. Least squares percentage regression. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/387.

Torkaman:2018:OMS

[Tor18] Parisa Torkaman. Optimal model selection for truncated data among non-nested competitive models. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/965.

Trivedi:2016:EPM

[TP16] Bhaktida S. Trivedi and M. N. Patel. Estimation of parameters of misclassified size biased Borel distribution. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/862.

Tailor:2011:IEP

[TS11] Rajesh Tailor and Balkishan Sharma. Improved estimation of the population mean using known parameters of an auxiliary variable. Journal of Modern Applied Statistical Methods, 10(1): ??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/523.

Tarray:2016:NPE

[TS16] Tanveer A. Tarray and Housila P. Singh. New procedures of estimating proportion and sensitivity using randomized response in a dichotomous finite population. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/817.

Tarray:2018:PTS

[TS18] Tanveer A. Tarray and Housila P. Singh. A proficient two-stage stratified randomized response strategy. Journal of Modern Applied Statistical Methods, 17(1):??, ???? 2018. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/981.

Tobias:2004:MAR

[TSK04] Aurelio Tobías, Marc Saez, and Manolis Kogevinas. Metaanalysis of results and individual patient data in epidemiologal studies. Journal of Modern Applied Statistical Methods, 3 (1):176–185, May 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/127.

Tripathi:2020:DTR

[TSS20] Ayush Tripathi, Umesh Singh, and Sanjay Kumar Singh. Does the type of records affect the estimates of the parameters? Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/1092.

Thomas:2006:DOB

[TTBG06] Seemon Thomas, Alex Thannippara, S. C. Bagui, and D. K. Ghosh. Determination of optimal block designs with preassigned variance for elementary contrasts. Journal of Modern Applied Statistical Methods, 5(2):390–394, November 2006. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/256.

Tayyab:2019:EOR

[TuAH19] Muhammad Tayyab, Muhammad Noor ul Amin, and Muhammad Hanif. Even order ranked set sampling with auxiliary variable. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1051.

Turner:2009:MSP

[Tur09] Paul Turner. Multiple search paths and the general-to-specific methodology. Journal of Modern Applied Statistical Methods, 8 (2):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/440.

Turpin:2017:ECD

[Tur17] Lonnie Turpin. Effectively comparing differences in proportions. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/933.

Tian:2007:CTR

[TW07] Tian Tian and Rand R. Wilcox. A comparison of two rank tests for repeated measure designs. Journal of Modern Applied Statistical Methods, 6(1):331–335, May 2007. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/306.

Tsokos:2011:NHP

[TX11] Chris P. Tsokos and Yong Xu. Non-homogeneous Poisson process for evaluating stage I & II ductal breast cancer treatment. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/573.

Talpur:2008:TDM

[TZA08] Mir Ghulam Hyder Talpur, Iffat Zamir, and M. Masoom Ali. Two dimension marginal distributions of crossing time and renewal numbers related to two-stage Erlang processes. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/377.

Thomas:2008:MRI

[TZZD08] D. Roland Thomas, PengCheng Zhu, Bruno D. Zumbo, and Shantanu Dutta. On measuring the relative importance of explanatory variables in a logistic regression. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/343.

Udomboso:2016:ANI

[UAD16] Christopher Godwin Udomboso, Godwin Nwazu Amahia, and Isaac Kwame Dontwi. An adjusted network information criterion for model selection in statistical neural network models. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/859.

Udomboso:2013:SPH

[Udo13] Christopher Godwin Udomboso. On some properties of a heterogeneous transfer function involving symmetric saturated linear (SATLINS) with hyperbolic tangent (TANH) transfer functions. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/685.

Udomboso:2020:LPH

[Udo20] Christopher Godwin Udomboso. On the level of precision of a heterogeneous transfer function in a statistical neural network model. Journal of Modern Applied Statistical Methods, 19(1): ??, ???? 2020. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1077.

Uraibi:2020:VRS

[Ura20] Hassan S. Uraibi. VIF-regression screening ultrahigh dimensional feature space. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1074.

Vallabados:2016:EPL

[VA16] Christopher A. Vallabados and Subbarayan A. Arumugam. An evaluation of Pareto, lognormal and PPS distributions: the size distribution of cities in Kerala, India. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/827.

Varghese:2011:CBD

[VJ11] Cini Varghese and Seema Jaggi. Control balanced designs involving sequences of treatments. Journal of Modern Applied Statistical Methods, 10(2):??, ???? 2011. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/571.

Vorburger:2009:APW

[VM09] Michael Vorburger and Breda Munoz. Approximations to power when comparing two small independent proportions. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/411.

VandenNoortgate:2003:PBV

[VO03] Wim Van den Noortgate and Patrick Onghena. A parametric bootstrap version of Hedges’ homogeneity test. Journal of Modern Applied Statistical Methods, 2(1):73–79, May 2003. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/67.

Vishwakarma:2008:OAF

[VS08] Gajendra K. Vishwakarma and Housila P. Singh. An optimum allocation with a family of estimators using auxiliary information in sample survey. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/382.

Vossoughi:2016:CSM

[VSS16] Mehrdad Vossoughi, Shila Shahvali, and Erfan Sadeghi. Comparison of some multivariate nonparametric tests in profile analysis to repeated measurements. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/852.

Vishwakarma:2020:GRP

[VZ20] Gajendra Kumar Vishwakarma and Sayed Mohammed Zeeshan. Generalized ratio-cum-product estimator for finite population mean under two-phase sampling scheme. Journal of Modern Applied Statistical Methods, 19(1):??, ???? 2020. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1076.

Wehry:2003:TER

[WA03] Stephanie Wehry and James Algina. Type I error rates of four methods for analyzing data collected in a groups vs individuals design. Journal of Modern Applied Statistical Methods, 2(2): 400–413, November 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/97.

Wehry:2007:AUG

[WA07] Stephanie Wehry and James Algina. Analyses of unbalanced groups-versus-individual research designs using three alternative approximate degrees of freedom tests: Test development and Type I error rates. Journal of Modern Applied Statistical Methods, 6(1):53–65, May 2007. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/282.

Walker:2003:JCK

[Wal03] David A. Walker. JMASM 9: Converting Kendall’s tau for correlational or meta-analytic analyses. Journal of Modern Applied Statistical Methods, 2(2):525–530, November 2003. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/110.

Walker:2004:VSM

[Wal04] David A. Walker. Validation studies: Matters of dimensionality, accuracy, and parsimony with predictive discriminant analysis and factor analysis. Journal of Modern Applied Statistical Methods, 3(1):186–199, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/128.

Walker:2005:BAT

[Wal05a] David A. Walker. Bias affiliated with two variants of Cohen’s d when determining U1 as a measure of the percent of nonoverlap. Journal of Modern Applied Statistical Methods, 4(1): 100–105, May 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/172.

Walker:2005:JSM

[Wal05b] David A. Walker. JMASM 19: a SPSS matrix for determining effect sizes from three categories: r and functions of r, differences between proportions, and standardized differences between means. Journal of Modern Applied Statistical Methods, 4 (1):333–342, May 2005. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/193.

Walker:2006:CSB

[Wal06] David A. Walker. A comparison of the Spearman–Brown and Flanagan–Rulon formulas for split half reliability under various variance parameter conditions. Journal of Modern Applied Statistical Methods, 5(2):443–451, November 2006. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/261.

Walker:2007:CES

[Wal07] David A. Walker. A comparison of eight shrinkage formulas under extreme conditions. Journal of Modern Applied Statistical Methods, 6(1):162–172, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/291.

Walker:2008:EHM

[Wal08] David A. Walker. Estimating how many observations are needed to obtain a required level of reliability. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/351.

Walker:2015:JTG

[Wal15] David A. Walker. JMASM 34: Two group program for Cohen’s d, Hedges’ g, η2, Radj2, ω2, ϵ2, confidence intervals, and power. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/786.

Walker:2016:JCI

[Wal16a] David A. Walker. JMASM 38: Confidence intervals for Kendall’s tau with small samples (SPSS). Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/833.

Walker:2016:GCO

[Wal16b] Robert W. Walker. On generalizing cumulative ordered regression models. Journal of Modern Applied Statistical Methods, 15 (2):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/861.

Walker:2017:RFE

[Wal17a] David A. Walker. In response to frane, “Errors in a program for approximating confidence intervals”. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/924.

Walker:2017:JPP

[Wal17b] David A. Walker. JMASM 48: the Pearson product-moment correlation coefficient and adjustment indices: the Fisher approximate unbiased estimator and the Olkin–Pratt adjustment (SPSS). Journal of Modern Applied Statistical Methods, 16(2): ??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/955.

Wang:2003:CEP

[Wan03] Daniel X. Wang. Comparison of estimates of proprietary and syndicated methods in auto industry surveys. Journal of Modern Applied Statistical Methods, 2(1):195–201, May 2003. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/78.

Watkins:2006:DPA

[Wat06] Marley W. Watkins. Determining parallel analysis criteria. Journal of Modern Applied Statistical Methods, 5(2):344–346, November 2006. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/251.

Wolf-Branigin:2002:ASR

[WB02] Michael Wolf-Branigin. Applying spatial randomness to community inclusion. Journal of Modern Applied Statistical Methods, 1(1):110–113, May 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/14.

Wolf-Branigin:2006:SPF

[WB06] Michael Wolf-Branigin. Supporting and preparing future decision-makers with the needed tools. Journal of Modern Applied Statistical Methods, 5(2):538–541, November 2006. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/270.

Wu:2007:EPA

[WBN07] Meng-Jia Wu, Betsy Jane Becker, and Yael Netz. Effects of physical activity on psychological change in advanced age: a multivariate meta-analysis. Journal of Modern Applied Statistical Methods, 6(1):2–7, May 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/277.

Wolf-Branigin:2009:ACD

[WBSMI09] Michael Wolf-Branigin, Hyon-Sook Suh, Star Muir, and Emily S. Ihara. Applying census data for small area estimation in community and social service planning. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/422.

Wendell:2005:CPO

[WC05] John P. Wendell and Sharon P. Cox. Coverage properties of optimized confidence intervals for proportions. Journal of Modern Applied Statistical Methods, 4(1):43–52, May 2005. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/165.

Wilcox:2009:QRI

[WC09] Rand R. Wilcox and Kathleen Costa. Quantile regression: on inferences about the slopes corresponding to one, two or three quantiles. Journal of Modern Applied Statistical Methods, 8(1): ??, ???? 2009. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/396.

Wilcox:2013:RRE

[WC13] Rand R. Wilcox and Florence Clark. Robust regression estimators when there are tied values. Journal of Modern Applied Statistical Methods, 12(2):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/662.

Wu:2008:WSF

[WCW08] Dongfeng Wu, Ricolindo L. Cariño, and Xiaoqin Wu. When sensitivity is a function of age and time spent in the preclinical state in periodic cancer screening. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/364.

Wilcox:2005:IAR

[WE05] Rand R. Wilcox and Mitchell Earleywine. Inferences about regression interactions via a robust smoother with an application to cannabis problems. Journal of Modern Applied Statistical Methods, 4(1):53–62, May 2005. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/167.

Weber:2007:EDD

[Web07] Michèle Weber. The effect of different degrees of freedom of the chi-square distribution on the statistical power of the t, permutation t, and Wilcoxon tests. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/317.

Wehry:2003:ALR

[Weh03] Stephanie Wehry. Alphabet letter recognition and emergent literacy abilities of rising kindergarten children living in low-income families. Journal of Modern Applied Statistical Methods, 2(2): 451–466, November 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/102.

Whittaker:2009:CMS

[WF09] Tiffany A. Whittaker and Carolyn F. Furlow. The comparison of model selection criteria when selecting among competing hierarchical linear models. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/409.

Wang:2015:CDV

[WFC15] Xuequn Wang, Brian F. French, and Paul F. Clay. Convergent and discriminant validity with formative measurement: a mediator perspective. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/756.

Whittaker:2002:DPI

[WFW02] Tiffany A. Whittaker, Rachel T. Fouladi, and Natasha J. Williams. Determining predictor importance in multiple regression under varied correlational and distributional conditions.

Journal of Modern Applied Statistical Methods, 1(2):354–366, November 2002. CODEN ???? ISSN 1538-9472. URL https:/

/jmasm.com/index.php/jmasm/article/view/41.

Wilcox:2016:WGA

[WH16] Rand Wilcox and Timothy Hayes. Within groups ANOVA when using a robust multivariate measure of location. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/839.

White:2007:RLK

[Whi07] Ian R. White. Regarding Lui K. J. (2006). Interval estimation of risk difference in simple compliance randomized trials. JMASM, 5, 395–407. Journal of Modern Applied Statistical Methods, 6(2): ??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/339. See [Lui06]

and reply [Lui07].

Wilson:2004:GQL

[Wil04] Jeffrey R. Wilson. A generalized quasi-likelihood model application to modeling poverty of Asian American women. Journal of Modern Applied Statistical Methods, 3(1):165–175, May 2004. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/126.

Wilcox:2005:WWA

[Wil05] Rand R. Wilcox. Within by within ANOVA based on medians. Journal of Modern Applied Statistical Methods, 4(1):2–10, May 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/162.

Wilcox:2006:ARO

[Wil06a] Rand R. Wilcox. ANCOVA: a robust omnibus test based on selected design points. Journal of Modern Applied Statistical Methods, 5(1):14–21, May 2006. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/223.

Wilcox:2006:IAC

[Wil06b] Rand R. Wilcox. Inferences about the components of a generalized additive model. Journal of Modern Applied Statistical Methods, 5(2):??, ???? 2006. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/246.

Wilcox:2007:OTW

[Wil07a] Rand R. Wilcox. An omnibus test when using a regression estimator with multiple predictors. Journal of Modern Applied Statistical Methods, 6(2):??, ???? 2007. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/311.

Wilcox:2007:FTI

[Wil07b] Rand R. Wilcox. On flexible tests of independence and homoscedasticity. Journal of Modern Applied Statistical Methods, 6(1):??, ???? 2007. CODEN ???? ISSN 1538-9472. URL https:

//jmasm.com/index.php/jmasm/article/view/280.

Wilcox:2008:EEP

[Wil08a] Rand R. Wilcox. Estimating explanatory power in a simple regression model via smoothers. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/371.

Wilcox:2008:TIQ

[Wil08b] Rand R. Wilcox. On a test of independence via quantiles that is sensitive to curvature. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/342.

Wilcox:2010:IAP

[Wil10] Rand R. Wilcox. Inferences about the population mean: Empirical likelihood versus bootstrap-t. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/461.

Wilcox:2011:CSA

[Wil11] Rand R. Wilcox. Comparing the strength of association of two predictors via smoothers or robust regression estimators. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/519.

Wilcox:2012:CTI

[Wil12] Rand R. Wilcox. Comparing two independent groups via a quantile generalization of the Wilcoxon–Mann–Whitney test. Journal of Modern Applied Statistical Methods, 11(2):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/610.

Wilcox:2015:ESA

[Wil15a] Rand Wilcox. Estimating the strength of an association based on a robust smoother. Journal of Modern Applied Statistical Methods, 14(1):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/749.

Wilcox:2015:IAS

[Wil15b] Rand Wilcox. Inferences about the skipped correlation coefficient: Dealing with heteroscedasticity and non-normality. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/769.

Wilcox:2016:AGT

[Wil16a] Rand Wilcox. ANCOVA: a global test based on a robust measure of location or quantiles when there is curvature. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/788.

Wilcox:2016:CTQ

[Wil16b] Rand Wilcox. Comparisons of two quantile regression smoothers. Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/790.

Wilcox:2017:RSL

[Wil17a] Rand Wilcox. The regression smoother LOWESS: a confidence band that allows heteroscedasticity and has some specified simultaneous probability coverage. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/926.

Wilcox:2017:RAC

[Wil17b] Rand Wilcox. Robust ANCOVA: Confidence intervals that have some specified simultaneous probability coverage when there is curvature and two covariates. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/882.

Wilcox:2018:IMD

[Wil18a] Rand Wilcox. An inferential method for determining which of two independent variables is most important when there is curvature. Journal of Modern Applied Statistical Methods, 17(1): ??, ???? 2018. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/958.

Wilcox:2018:LRI

[Wil18b] Rand Wilcox. Logistic regression: an inferential method for identifying the best predictors. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/989.

Wilcox:2018:RAC

[Wil18c] Rand Wilcox. Robust ANCOVA, curvature, and the curse of dimensionality. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/988.

Wilcox:2018:RNM

[Wil18d] Rand Wilcox. A robust nonparametric measure of effect size based on an analog of Cohen’s d, plus inferences about the median of the typical difference. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/987.

Wilcox:2019:BAW

[Wil19a] Rand Wilcox. Bivariate analogs of the Wilcoxon–Mann–Whitney test and the Patel–Hoel method for interactions. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1010.

Wilcox:2019:IWJ

[Wil19b] Rand Wilcox. Identifying which of J independent binomial distributions has the largest probability of success. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/1042.

Wilcox:2019:IAP

[Wil19c] Rand Wilcox. Inferences about the probability of success, given the value of a covariate, using a nonparametric smoother. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/1016.

Wilcox:2019:NIA

[Wil19d] Rand Wilcox. A note on inferences about the probability of success. Journal of Modern Applied Statistical Methods, 18(1): ??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1017.

Wilcox:2019:RDW

[Wil19e] Rand Wilcox. Regression: Determining which of p independent variables has the largest or smallest correlation with the dependent variable, plus results on ordering the correlations Winsorized. Journal of Modern Applied Statistical Methods, 18(2): ??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1043.

Wilcox:2019:RWT

[Wil19f] Rand Wilcox. Regression when there are two covariates: Some practical reasons for considering quantile grids. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/1011.

Wilcox:2002:PAW

[WK02a] Rand R. Wilcox and H. J. Keselman. Power analyses when comparing trimmed means. Journal of Modern Applied Statistical Methods, 1(1):24–31, May 2002. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/5.

Wilcox:2002:WGM

[WK02b] Rand R. Wilcox and H. J. Keselman. Within groups multiple comparisons based on robust measures of location. Journal of Modern Applied Statistical Methods, 1(2):281–287, November 2002. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/34.

Wilcox:2004:MLR

[WK04] Rand R. Wilcox and H. J. Keselman. Multivariate location: Robust estimators and inference. Journal of Modern Applied Statistical Methods, 3(1):2–12, May 2004. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/111.

Wang:2018:EEC

[WPN^{+}18] Yan Wang, Thanh Pham, Diep Nguyen, Eun Sook Kim, and Yi-Hsin Chen. Evaluating the efficacy of conditional analysis of variance under heterogeneity and non-normality. Journal of Modern Applied Statistical Methods, 17(2):??, ???? 2018. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1007.

Weber:2009:CPI

[WS09] Michèle Weber and Shlomo Sawilowsky. Comparative power of the independent t, permutation t, and Wilcoxon tests. Journal of Modern Applied Statistical Methods, 8(1):??, ???? 2009. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/397.

Walker:2016:JNP

[WS16a] David A. Walker and Thomas J. Smith. JMASM 36: Nine pseudo R^{2} indices for binary logistic regression models (SPSS). Journal of Modern Applied Statistical Methods, 15(1):??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/831.

Wang:2016:NTC

[WS16b] Ping Wang and Ping Sa. A new test for correlation on bivariate nonnormal distributions. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/851.

Walker:2019:LRU

[WS19] David A. Walker and Thomas J. Smith. Logistic regression under sparse data conditions. Journal of Modern Applied Statistical Methods, 18(2):??, ???? 2019. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/1058.

Wooluru:2015:ADU

[WSN15] Yerriswamy Wooluru, D. R. Swamy, and P. Nagesh. Approaches for detection of unstable processes: a comparative study. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/782.

Wu:2004:VAB

[Wu04] Dongfeng Wu. A visually adaptive Bayesian model in wavelet regression. Journal of Modern Applied Statistical Methods, 3 (1):200–212, May 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/129.

Wang:2010:GBS

[WW10] Li-Chih Wang and Chin-Lien Wang. A GA-based sales forecasting model incorporating promotion factors. Journal of Modern Applied Statistical Methods, 9(2):??, ???? 2010. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/509.

Wang:2012:IVF

[WWB12] Dandan Wang, Tiffany A. Whittaker, and S. Natasha Beretvas. The impact of violating factor scaling method assumptions on latent mean difference testing in structured means models. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/585.

Walker:2003:EIW

[WY03] David A. Walker and Denise Y. Young. Example of the impact of weights and design effects on contingency tables and chi-square analysis. Journal of Modern Applied Statistical Methods, 2(2): 425–432, November 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/99.

Wu:2017:UPI

[WZ17] Amrey D. Wu and Bruno D. Zumbo. Using Pratt’s importance measures in confirmatory factor analyses. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/928.

Wu:2011:GPG

[WZS11] Amery D. Wu, Bruno D. Zumbo, and Linda S. Siegel. General piecewise growth mixture model: Word recognition development for different learners in different phases. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/537.

Xiong:2008:MOH

[XGY^{+}08] Chengjie Xiong, Feng Gao, Yan Yan, Jingqin Luo, and YunJu Sung. Measuring overall heterogeneity in meta-analyses: Application to CSF biomarker studies in Alzheimer’s disease. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/363.

Xu:2005:EUD

[Xu05] Yonghong Jade Xu. An exploration of using data mining in educational research. Journal of Modern Applied Statistical Methods, 4(1):251–274, May 2005. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/185.

Xiong:2003:THU

[XYJ03] Chengjie Xiong, Yan Yan, and Ming Ji. Test of homogeneity for umbrella alternatives in dose–response relationship for Poisson variables. Journal of Modern Applied Statistical Methods, 2(2): 389–399, November 2003. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/96.

Xiong:2005:SME

[XZ05] Chengjie Xiong and Kejun Zhu. Statistical model and estimation of the optimum price for a chain of price setting firms. Journal of Modern Applied Statistical Methods, 4(2):553–566, November 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/212.

Yousef:2014:EGP

[YAS14] Omar M. Yousef and Sameer A. Al-Subh. Estimation of Gumbel parameters under ranked set sampling. Journal of Modern Applied Statistical Methods, 13(2):??, ???? 2014. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/741.

Ye:2017:IIM

[YD17] Feifei Ye and Laura Danie. The impact of inappropriate modeling of cross-classified data structures on random-slope models. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/948.

Yang:2017:CDM

[YHPR17] Si Yang, Lisa L. Harlow, Gavino Puggion, and Colleen A. Redding. A comparison of different methods of zero-inflated data analysis and an application in health surveys. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN

???? ISSN 1538-9472. URL https://jmasm.com/index.php/
jmasm/article/view/910.

Yap:2013:AAR

[YII13] Grace Lee Ching Yap, Wan Rosmanira Ismail, and Zaidi Isa. An alternative approach to reduce dimensionality in data envelopment analysis. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/648.

Yaya:2014:EFG

[YOO14] OlaOluwa S. Yaya, Olusanya E. Olubusoye, and Oluwadare O. Ojo. Estimates and forecasts of GARCH model under misspecified probability distributions: a Monte Carlo simulation approach. Journal of Modern Applied Statistical Methods, 13(2): ??, ???? 2014. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/745.

Yaya:2014:SAS

[YS14] OlaOluwa S. Yaya and Olanrewaju I. Shittu. Specifying asymmetric STAR models with linear and nonlinear GARCH innovations: Monte Carlo approach. Journal of Modern Applied Statistical Methods, 13(1):??, ???? 2014. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/713.

Yaya:2016:SVL

[YS16] OlaOluwa S. Yaya and Olanrewaju I. Shittu. Symmetric variants of logistic smooth transition autoregressive models: Monte Carlo evidences. Journal of Modern Applied Statistical Methods, 15(1): ??, ???? 2016. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/820.

Yu:2019:QLR

[YSVS19] Lili Yu, Varadan Sevilimedu, Robert Vogel, and Hani Samawi. Quasi-likelihood ratio tests for homoscedasticity in linear regression. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/1029.

Yamamoto:2010:SPQ

[YT10] Kouji Yamamoto and Sadao Tomizawa. Symmetry plus quasi uniform association model and its orthogonal decomposition for square contingency tables. Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/483.

Yadav:2024:DEM

[YTP24] Amita Yadav, Narendra Singh Thakur, and Sarla Pareek. Determining the effect of measurement errors and estimation of mean by factor-type estimator in simple random sampling. Journal of Modern Applied Statistical Methods, 23(1):??, ???? 2024. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1164.

Zhang:2008:CPN

[ZA08] Guili Zhang and James Algina. Coverage performance of the non-central F -based and percentile bootstrap confidence intervals for root mean square standardized effect size in one-way fixed-effects ANOVA. Journal of Modern Applied Statistical Methods, 7(1):??, ???? 2008. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/345.

Zhang:2011:RRM

[ZA11] Guili Zhang and James Algina. A robust root mean square standardized effect size in one-way fixed-effects ANOVA. Journal of Modern Applied Statistical Methods, 10(1):??, ???? 2011. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/525.

Zadkarami:2013:BIP

[ZC13] M. R. Zadkarami and O. Chatrabgoun. Bayesian inference of pair-copula constriction for multivariate dependency modeling of Iran’s macroeconomic variables. Journal of Modern Applied Statistical Methods, 12(1):??, ???? 2013. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/656.

Zhang:2017:UMI

[ZCB^{+}17] Yujia Zhang, Sara Crawford, Sheree Boulet, Michael Monsour, and Bruce Cohen. Using multiple imputation to address missing values of hierarchical data. Journal of Modern Applied Statistical Methods, 16(1):??, ???? 2017. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/915.

Zhang:2010:JPL

[ZD10] Dongquan Zhang and C. Mitchell Dayton. JMASM 30 PI-LCA: a SAS program computing the two-point mixture index of fit for two-class LCA models with dichotomous variables (SAS). Journal of Modern Applied Statistical Methods, 9(1):??, ???? 2010. CODEN ???? ISSN 1538-9472. URL https://jmasm.com/

index.php/jmasm/article/view/490.

Zumbo:2007:OVC

[ZGZ07] Bruno D. Zumbo, Anne M. Gadermann, and Cornelia Zeisser. Ordinal versions of coefficients alpha and theta for Likert rating scales. Journal of Modern Applied Statistical Methods, 6(1):21– 29, May 2007. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/279.

Zhang:2005:TGF

[Zha05] Biao Zhang. Testing the goodness of fit of multivariate multiplicative-intercept risk models based on case-control data. Journal of Modern Applied Statistical Methods, 4(1):11–34, May 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/163.

Zin:2008:NPQ

[ZJ08] Wan Zawiah Wan Zin and Abdul Aziz Jemain. Non-parametric quantile selection for extreme distributions. Journal of Modern Applied Statistical Methods, 7(2):??, ???? 2008. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/379.

Zumbo:2005:MDI

[ZK05] Bruno D. Zumbo and Kim H. Koh. Manifestation of differences in item-level characteristics in scale-level measurement invariance tests of multi-group confirmatory factor analyses. Journal of Modern Applied Statistical Methods, 4(1):275–282, May 2005. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/186.

Zumbo:2016:SRR

[ZK16] Bruno D. Zumbo and Edward Kroc. Some remarks on Rao and Lovric’s ‘Testing Point Null Hypothesis of a Normal Mean and the Truth: 21st Century Perspective’. Journal of Modern Applied Statistical Methods, 15(2):??, ???? 2016. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/838. See [RL16].

Zhou:2004:NES

[ZL04] Xinyue Zhou and Joel R. Levin. A note on extending Scheffé’s modified multiple-comparison procedure to other analysis situations. Journal of Modern Applied Statistical Methods, 3(2): 432–442, November 2004. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/155.

Zou:2019:USA

[ZLB19] Danjie Zou, Jennifer E. V. Lloyd, and Jennifer L. Baumbusch. Using SPSS to analyze complex survey data: a primer. Journal of Modern Applied Statistical Methods, 18(1):??, ???? 2019. CO-

DEN ???? ISSN 1538-9472. URL https://jmasm.com/index.
php/jmasm/article/view/1026.

Zhou:2023:WES

[ZRB23] Yi Zhou, Xinyue Ren, and Gordon Brooks. Which effect size calculation is the best to estimate the population effect size in the Welch T test? Journal of Modern Applied Statistical Methods, 22(1):??, February 2023. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/1132.

Zaninotto:2017:MDL

[ZS17] Paola Zaninotto and Amanda Sacker. Missing data in longitudinal surveys: a comparison of performance of modern techniques. Journal of Modern Applied Statistical Methods, 16(2):??, ???? 2017. CODEN ???? ISSN 1538-9472. URL https://jmasm.

com/index.php/jmasm/article/view/943.

Zhang:2006:VEC

[ZT06] Shenghai Zhang and Mary E. Thompson. Variance estimation and construction of confidence intervals for GEE estimator. Journal of Modern Applied Statistical Methods, 5(1):217– 229, May 2006. CODEN ???? ISSN 1538-9472. URL https://

jmasm.com/index.php/jmasm/article/view/238.

Zhang:2012:SUC

[ZT12] Lin Zhang and Linjuan Tang. A study on underwriting cycle of property insurance industry of China. Journal of Modern Applied Statistical Methods, 11(1):??, ???? 2012. CODEN ????

ISSN 1538-9472. URL https://jmasm.com/index.php/jmasm/
article/view/605.

Zumbo:2002:AIS

[Zum02] Bruno D. Zumbo. An adaptive inference strategy: the case of auditory data. Journal of Modern Applied Statistical Methods, 1(1):61–68, May 2002. CODEN ???? ISSN 1538-9472. URL

https://jmasm.com/index.php/jmasm/article/view/9.

Zimmerman:2015:RIH

[ZZ15] Donald W. Zimmerman and Bruno D. Zumbo. Resolving the issue of how reliability is related to statistical power: Adhering to mathematical definitions. Journal of Modern Applied Statistical Methods, 14(2):??, ???? 2015. CODEN ???? ISSN 1538-

9472. URL https://jmasm.com/index.php/jmasm/article/
view/770.